### v2

# Pinpointing Important Genetic Variants via A Feature-level Filter and Group Knockoffs

Abstract
Identifying variants that carry substantial information on the trait of interest remains a core topic in genetic studies. In analyzing the EADB-UKBB dataset to identify genetic variants associated with Alzheimer’s disease (AD), however, we recognize that both existing marginal association tests and conditional independence tests using knockoffs suffer either power loss or lack of informativeness, especially when strong correlations exist among variants. To address the limitations of existing knockoff filters, we propose a new feature-versus-group (FVG) filter that is more powerful and precise in identifying important features from a set of strongly correlated features using group knockoffs. In extensive simulation studies, the FVG filter controls the expected proportion of false discoveries and identifies important features with enhanced power and greater precision. Applying the proposed method to the EADB-UKBB dataset, we discover important variants from 84 loci (same as the most powerful group knockoff filter) with catching sets of substantially smaller size and higher purity.
Keywords: Alzheimer’s disease genetics, False discovery rate, Feature selection, Knockoffs.
1 Introduction
1.1 Conditional Independence Test to Discover Disease-associated Variants
Testing conditional independence between a set of features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the response of interest $Y$ is an important topic in various research areas, including causal inference (Peters,, 2015; Cai et al.,, 2022), genetic analysis (Khera and Kathiresan,, 2017; Zhu et al.,, 2018) and graphical model learning (Deka et al.,, 2016; Tugnait,, 2022). In the era of big data, as the number of features increases, the need for statistical approaches for simultaneous inference of conditional independence between hundreds of thousands of features and the response keeps emerging.
Specifically, developing statistical methods for conditional independence tests with type-I error rate control has been an important and popular research topic of large-scale genome-wide association studies (GWAS; Tang and He,, 2021; Hou et al.,, 2023; Morra et al.,, 2023) in recent years. With the ultimate goal of identifying novel targets for the development of genomic-driven medicine, the contribution of more genetic variants to complex phenotypes has been investigated in GWAS. However, our ability to detect causal genetic variants and to translate discoveries into insights of the underlying genetic mechanisms does not increase proportionally to the growth of the scale of GWAS. On one hand, with the stringent criterion to control the family-wise error rate (FWER), conventional GWAS are usually suboptimal in terms of the statistical power of detecting important variants. On the other hand, because conventional GWAS base their inference on marginal association models which regress $Y$ on one $X_{j}$ at a time, they can only provide blurred results with a lot of proxy variants correlated with the true causal variants (Schaid et al.,, 2018). Both issues become more severe when more variants are sequenced at a higher resolution because the $p$-value threshold of FWER control also becomes more stringent and there are more proxy variants strongly correlated with true causal variants. Although many fine-mapping methods have been proposed (Benner et al.,, 2017; Zou et al.,, 2022) to pinpoint the causal variants, they can only partially address the aforementioned issues without control of type-I error rate, calling for new statistical inference methods with better theoretical properties and empirical performance in multiple testing of conditional independence for real-world genetic data analysis.
In this article, we investigate the stage I meta-analysis of the European Alzheimer & Dementia Biobank (EADB) dataset and the UK Biobank (UKBB) dataset (Bellenguez et al.,, 2022) to identify genetic variants associated with AD. This meta-analyzed EADB–UKBB dataset is deposited in the European Bioinformatics Institute GWAS Catalog [^1] https://www.ebi.ac.uk/gwas/ (Accession number: GCST90027158), including $39,106$ clinically diagnosed AD cases, $46,828$ proxy cases, and $401,577$ control cases from $15$ European countries. We focus on $640,576$ directly genotyped variants with minor allele frequency (MAF) at least $0.01$, resulting in a processed dataset $\mathbb{G}$ of $487,511$ observations and $640,576$ features. With such a large dataset, we aim to discover as many important genetic variants that carry independent information of AD without making too many false discoveries. Mathematically, such a task can be transferred into a multiple testing problem of conditional independence between a large set of variants $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the binary response $Y$ which indicates whether a case is clinically diagnosed AD.
1.2 Existing Approaches and Their Limitations
Recent decades have witnessed fruitful works in the development of multiple testing procedures. To control the family-wise error rate (FWER) of making at least one false discovery, a lot of $p$-value-based methods have been developed as improvements of the classic Bonferroni correction, including the Šidák correction (Šidák,, 1967), Holm’s step-down procedure (Holm,, 1979) and Hochberg’s step-up procedure (Hochberg,, 1988). However, FWER has been criticized for its conservativeness in many settings, especially when most of the signals are weak. To perform powerful inference, Benjamini and Hochberg, (1995) proposed the false discovery rate (the expected proportion of false discoveries in all discoveries) as an alternative type-I error rate measure and developed a procedure to perform multiple testing with provable FDR control under the assumption that $p$-values are independent or positively dependent. Inspired by this pioneering work, Benjamini and Yekutieli, (2001) developed a generalized procedure without assumptions on dependencies among $p$-values, while Storey, (2002) and Whittemore, (2007) investigated the Bayesian counterpart of FDR. However, as most of the aforementioned methods require $p$-values of nulls for all features, their feasibility on multiple testing of conditional independence becomes questionable when the number of features $p$ greatly exceeds the sample size $n$. The main reason is that $p$-values of conditional independence under high-dimensional scenarios are nontrivial and generally not available except in a very few specific cases (Tibshirani et al.,, 2016; Lee et al.,, 2016).
Unlike all the above methods, the recently developed knockoff filter (Barber and Candès,, 2015; Candès et al.,, 2018) and its variants (Gimenez et al.,, 2019; Gimenez and Zou,, 2019; Barber and Candès,, 2019; Huang and Janson,, 2020; Ren and Candès,, 2023) are a family of approaches that can control finite-sample FDR in multiple testing of conditional independence hypotheses at the feature-versus-feature level
$H^{(\text{ff})}_{j}:X_{j}\perp Y|\textbf{X}_{-j},\text{ where }\textbf{X}_{-j}=(X_{1},\ldots,X_{j-1},X_{j+1},\ldots,X_{p})^{\mathsf{T}},\quad j=1,\ldots,p,$ (1)
without any assumption on the conditional distribution $Y|\textbf{X}$. Specifically, these methods first construct knockoffs $\widetilde{\textbf{X}}$ that mimic the dependency among features X and are not dependent to the response $Y$. By doing so, knockoffs play the role of synthetic controls: if $H^{(\text{ff})}_{j}$ is true, ${X}_{j}$ and $\widetilde{X}_{j}$ are indistinguishable even with information of $Y$; if $H^{(\text{ff})}_{j}$ is false, ${X}_{j}$ and $\widetilde{X}_{j}$ possess different dependencies with $Y$. By comparing dependencies between $Y$ and X with dependencies between $Y$ and $\widetilde{\textbf{X}}$, these methods have both controlled FDR and superior power in detecting true causal features as validated in extensive simulation studies.
However, in the analysis of real-world genetic data, methods that infer $H^{(\text{ff})}_{j}$’s usually suffer severe power loss. The main reason is that linkage disequilibrium could introduce high correlations among genetic variants, making it difficult to distinguish variants that are truly associated with the response (whose $H^{(\text{ff})}_{j}$’s are false) from their strongly correlated null variants (whose $H^{(\text{ff})}_{j}$’s are true). This would significantly decrease the number of discoveries as shown by the simulation studies in Barber and Candès, (2015) and Candès et al., (2018). Such a power loss issue would deteriorate when more genetic variants with strong correlations are recorded under a growing resolution of sequencing. For example, when we apply the model-X knockoff filter (Candès et al.,, 2018) to the EADB-UKBB dataset (Bellenguez et al.,, 2022), only 138 AD-associated genetic variants from 49 loci (here, we define two loci as different if they are at least 1Mb away from each other) are identified under the target FDR level $0.10$ as shown in Table 1. Compared with the 51 loci discovered by marginal association test ($p$-value threshold: $5\times 10^{-8}$) in Bellenguez et al., (2022), although model-X knockoff filter identified 6 new loci, it misses several important AD loci with strong signals, including the “FAM193B” locus, the “FAM157C” locus, the “NSF” locus, the “SIGLEC11” locus and the “APP” locus.
Table 1: Summary of results by applying different methods to the EADB-UKBB dataset.
Method Number of Average number of Average Average
identified variants per size of purity of
loci identified locus catching sets catching sets
Marginal association test 51 18.078 18.078 0.477
Model-X knockoff filter 49 2.816 1.000 1.000
Group knockoff filter 84 23.714 13.105 0.610
: Catching sets refer to
(a) sets of variants within different loci identified by marginal association test;
(b) variants identified by the model-X knockoff filter;
(c) variant groups identified by the group knockoff filter.
: Purity refers to the minimum absolute correlation within the catching set.
To address this power loss issue, Dai and Barber, (2016) proposed the group knockoff filter. By defining groups $B_{1},\ldots,B_{k}$ such that features with strong correlations are in the same group and changing the inference target as
$H^{\text{(gg)}}_{k}:\textbf{X}_{B_{k}}\perp Y|\textbf{X}_{-B_{k}},\quad k=1,\ldots,K,$ (2)
at the group-versus-group level, important features and their strongly correlated null features can be rejected as a group. Because the grouping is done to ensure features from different groups are weakly correlated, the group knockoff filter can circumvent the obstacle of high correlation in inferring $H^{(\text{ff})}_{j}$ with significantly higher power. To apply the group knockoff filter to the EADB-UKBB dataset (Bellenguez et al.,, 2022), we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ and construct $321,535$ variant groups by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$. By doing so, variants in the EADB-UKBB dataset (Bellenguez et al.,, 2022) are grouped such that the average absolute correlation of variants from different groups is lower than 0.5. Under such a group structure, the group knockoff filter identifies 152 AD-associated variant groups from 84 loci under the target FDR level $0.10$. Specifically, among these 152 identified groups, 126 are located outside the APOE/APOC region with the strongest association to AD (chromosome 19, positions 44909011 $\sim$ 45912650), including groups in loci “FAM193B”, “FAM157C”, “NSF”, “SIGLEC11” and “APP” that are missed by the model-X knockoff filter. In addition, the group knockoff filter also identifies 39 new loci that are not discovered by Bellenguez et al., (2022).
However, doing so can only provide a series of AD-associated variant groups without information on which variants contribute more to AD variation within each group, leaving the target of identifying AD-associated variants not fully addressed. This can be observed from Table 1. With an average size $13.105$ of the identified groups given by the group knockoff filter, 35 identified groups contain at least 10 genetic variants and the 5 largest groups contain 314, 232, 136, 121 and 117 variants respectively. With such large sizes, which genetic variants in these groups contribute more to AD variation remains unanswered.
1.3 Our Contribution
In this article, to address the limitations of existing knockoff filters in discovering important variants of AD in the analysis of the EADB-UKBB dataset, we propose a new knockoff filter that is more powerful in inferring conditional independence hypotheses and identifying important features from a set of strongly correlated features. Our contribution in this article is two-folded.
  • Presenting a family of conditional independence hypotheses that allow identifying important features using group knockoffs. Compared to existing conditional independence hypotheses, the new hypotheses allow the use of group knockoffs construction mechanism (Dai and Barber,, 2016) to provide informative discoveries at the level of features with higher power.
  • Developing the new knockoff filter with FDR control, high empirical power and informativeness [^2] In this article, informativeness refers to the precision of catching sets in identifying conditionally independent signals and is quantified by the size and purity of catching sets.. Specifically, we thoroughly investigate the theoretical properties of the developed knockoff filter and empirically evaluate their performance via simulated experiments.
The rest of this article is organized as follows. In Section 2, we develop our knockoff filter and perform theoretical analysis. Section 3 illustrates the empirical performance of our method and its advantage over existing alternatives in power and informativeness via simulated experiments. Finally, we apply our method to the AD analysis of the EADB-UKBB dataset described in Section 4 and demonstrate that the method exhibits high power and precision. Section 5 concludes with discussions.
2 Methodology
2.1 Problem Statement
Consider independent and identically distributed (i.i.d.) samples $\{(\textbf{x}_{i},y_{i})|i=1,\ldots,n\}$ from a joint distribution $f(\mathbf{X},Y)$, where features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ are partitioned into $K$ disjoint groups $B_{1},\ldots,B_{K}$. Let $\mathbb{X}=[\textbf{x}_{1},\ldots,\textbf{x}_{n}]^{\mathsf{T}}$ denote the data matrix and $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ be the vector of responses. Our target is to perform multiple testing on conditional independence hypotheses,
$H_{j}^{(\text{fg})}:X_{j}\perp Y|\textbf{X}_{-B_{k}},\quad k=1,\ldots,K,\quad j\in B_{k},$ (3)
where $\textbf{X}_{-B_{k}}=(\textbf{X}^{\mathsf{T}}_{B_{1}},\ldots,\textbf{X}^{\mathsf{T}}_{B_{k-1}},\textbf{X}^{\mathsf{T}}_{B_{k+1}},\ldots,\textbf{X}^{\mathsf{T}}_{B_{K}})^{\mathsf{T}}$ is the subvector of X created by excluding all features in the same group $B_{k}$ of $X_{j}$. Here, we refer $H_{j}^{(\text{fg})}$ as the conditional independence hypothesis at the feature-versus-group level as it depicts the independence between a feature and the response conditional on other feature groups which omits information of other features in the same group. Under the belief that the response $Y$ only depends on a relatively small number of features, we would like to find as many important features without making too many false discoveries with respect to $H_{j}^{(\text{fg})}$’s. In other words, our target is to obtain a rejection set $\mathcal{R}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is rejected}\}$ such that the false discovery rate (FDR, the expectation of false discovery proportion),
$\text{FDR}^{(\text{fg})}=\mathbf{E}\{\text{FDP}^{(\text{fg})}\},\quad\text{where }\text{FDP}^{(\text{fg})}=\frac{\#(\mathcal{R}^{(\text{fg})}\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}},\quad\mathcal{H}_{0}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is true}\},$
is controlled under the target level $\alpha$ ($10<\alpha<$ 1).
Under the same setup of Candès et al., (2018) that the distribution of features X is completely known while no information is provided regarding the conditional distribution $Y|\textbf{X}$, our method differs in the hypotheses to test (3) from existing approaches, which test either conditional independence hypotheses (1) at the feature-versus-feature level or conditional independence hypotheses (2) at the group-versus-group level. Such a difference is the origin of our method’s advantages over the existing knockoff filters. On one hand, unlike the group knockoff filter (Dai and Barber,, 2016), the conclusion provided by our method is at the feature level and is thus more informative on which features contribute more to the variation of $Y$. On the other hand, compared to those that infer $H^{(\text{ff})}_{j}$’s, our method can also utilize group knockoffs that are more liberal in construction to gain power in selecting important features. However, such advantages do not come free as new challenge is posted that no existing filters can provide control on $\text{FDR}^{(\text{fg})}$, which is elaborated and addressed in the following sections.
2.2 Feature Importance Scores Under Group Knockoffs
To infer conditional independence hypotheses (3) with group knockoffs, our inference procedure, similar to the existing ones, include four main steps as follows.
  • $\diamond$
    (Knockoffs Construction): Construct the knockoff copy $\widetilde{\textbf{X}}$ of X such that $\perp Y|\textbf{X}$ X and the joint distribution of $(\textbf{X},\widetilde{\textbf{X}})$ is exchangeable at the group level (Dai and Barber,, 2016),
    $(\textbf{X},\widetilde{\textbf{X}})_{\text{swap}(\mathcal{S})}{\displaystyle\mathop{=\joinrel=}^{D}}(\textbf{X},\widetilde{\textbf{X}}),\quad\forall\mathcal{S}\subset\{1,\ldots,K\},$ (4)
    where $(\textbf{X},\widetilde{\textbf{X}})_{\text{swap}(\mathcal{S})}$ is obtained by swapping $\textbf{X}_{B_{k}}$ and $\widetilde{\textbf{X}}_{B_{k}}$ in $(\textbf{X},\widetilde{\textbf{X}})$ for all $k\in\mathcal{S}$.
  • $\diamond$
    (Feature Importance Scores Calculation): For each $H^{(\text{fg})}_{j}$, compute feature importance scores $T_{j}$ for feature $X_{j}$ and $\widetilde{T}_{j}$ for its knockoff copy $\widetilde{X}_{j}$.
  • $\diamond$
    (Feature Statistics Calculation): For each $H^{(\text{fg})}_{j}$, summarize importance scores $T_{j}$ and $\widetilde{T}_{j}$ as a feature statistic $W_{j}=w_{j}(T_{j},\widetilde{T}_{j})$ using an anti-symmetric function $w_{j}$ such that $w_{j}(T_{j},\widetilde{T}_{j})=-w_{j}(\widetilde{T}_{j},T_{j})$.
  • $\diamond$
    (Feature Filtering): Reject $H^{(\text{fg})}_{j}$ if $W_{j}$ is positive and large enough (i.e., $T_{j}$ dominates $\widetilde{T}_{j}$).
However, directly inputting $W_{1},\ldots,W_{p}$ to the model-X knockoff filter at feature level (Candès et al.,, 2018) cannot provide control on $\text{FDR}^{(\text{fg})}$ of the rejection set. The main reason is that under the group knockoffs model $f(\textbf{X},\widetilde{\textbf{X}})$, $W_{1},\ldots,W_{p}$ do not possess the coin-flipping property (i.e., conditional on $|W_{j}|$’s and $\text{sign}(W_{j})$’s for those false $H^{(\text{fg})}_{j}$’s, $\text{sign}(W_{j})$’s for those true $H^{(\text{fg})}_{j}$’s are independently and uniformly distributed on {±$\{\pm 1\}$), the key property that guarantees FDR control in Candès et al., (2018). Specifically, for most of the widely-used feature importance scores (e.g., the marginal correlations, the Lasso coefficient-difference statistics and the Lasso signed max statistics; Barber and Candès,, 2015, Candès et al.,, 2018), $W_{j}$’s for those true $H^{(\text{fg})}_{j}$’s are dependent coin flips. This comes from (4) where exchangeability only exists among groups, making signs of $W_{j}$’s from the same group not independent.
Thus, we relax the unachievable coin-flipping property to the between-group coin-flipping property which can be satisfied by feature importance scores $W_{j}$’s in the form of
$\begin{cases}T_{j}=t([\mathbb{X}_{j},\mathbb{X}_{B_{k}\setminus\{j\}},\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{k}\setminus\{j\}},\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}],\textbf{y}),\\ \widetilde{T}_{j}=t([\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{k}\setminus\{j\}},{\mathbb{X}}_{j},{\mathbb{X}}_{B_{k}\setminus\{j\}},\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}],\textbf{y}),\end{cases}\quad k=1,\ldots,K\text{, }j\in B_{k},$ (5)
where $\text{where }\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}\text{ denotes the set of unordered pairs }\{\mathbb{X}_{j^{\prime}},\widetilde{\mathbb{X}}_{j^{\prime}}\}\text{ for all }j^{\prime}\notin B_{k},$
as shown in Theorem 1.
Theorem 1.
For any feature importance scores $T_{j}$ and $\widetilde{T}_{j}$ in the form of (5), feature statistics $W_{j}=w_{j}(T_{j},\widetilde{T}_{j})$ ($j=1,\ldots,p$) with antisymmetric functions $w_{j}$’s satisfy the between-group coin-flipping property that
  1. $\star$
    conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$’s for those false $H^{(\text{fg})}_{j}$’s,
    1. $\diamond$
      (Uniformity) $\text{\rm sign}(W_{j})$ is uniformly distributed on {±$\{\pm 1\}$ for all $j\in\mathcal{H}_{0}^{(\text{fg})}$;
    2. $\diamond$
      (Between-group independence) for any $k\neq k^{\dagger}$, $\text{\rm sign}(W_{j})$ and $\text{\rm sign}(W_{j^{\dagger}})$ are independent for any $j\in B_{k}\cap\mathcal{H}_{0}^{(\text{fg})}$ and $j^{\dagger}\in B_{k^{\dagger}}\cap\mathcal{H}_{0}^{(\text{fg})}$,
if $H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ for all $j\in B_{k}$ and $k=1,\ldots,K$.
The proof of Theorem 1 is provided in Section A of the Appendix. Specifically, Theorem 1 relies on a condition “$H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ for all $j\in B_{k}$ and $k=1,\ldots,K$”. Such a condition is generally valid in genetic data analysis, analogous to the usual expectation in fine-mapping that the association between an important variant $X_{j}$ and the response $Y$ is in the same direction as the direct effect of $X_{j}$ on $Y$.
Examples of feature importance scores in the form of (5) include:
  • $\diamond$
    marginal correlation with response: absolute values of the empirical marginal correlations $T_{j}=\Big{|}\widehat{\rho}(X_{j},Y)\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\rho}(\widetilde{X}_{j},Y)\Big{|}$;
  • $\diamond$
    joint lasso: absolute values of lasso estimators $T_{j}=\Big{|}\widehat{\beta}_{j}\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\widetilde{\beta}}_{j}\Big{|}$ of the linear model
    $Y=\sum_{j=1}^{p}\left({\beta}_{j}{{X}}_{j}+\widetilde{\beta}_{j}\widetilde{{X}}_{j}\right)+e;$ (6)
  • $\diamond$
    marginal correlation with lasso residual: absolute values of estimated marginal correlations $T_{j}=\Big{|}\widehat{\rho}(X_{j},\hat{e})\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\rho}(\widetilde{X}_{j},\hat{e})\Big{|}$ between $(X_{j},\widetilde{X}_{j})^{\mathsf{T}}$ and the lasso residual of the linear model
    $Y=\sum_{j^{\dagger}\notin B_{k}}\left({\beta}_{j^{\dagger}}{{X}}_{j^{\dagger}}+\widetilde{\beta}_{j^{\dagger}}\widetilde{{X}}_{j^{\dagger}}\right)+e,\quad\text{where }j\in B_{k};$
  • $\diamond$
    separate lasso: absolute values of lasso estimators $T_{j}=\Big{|}\widehat{\beta}_{j}\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\widetilde{\beta}}_{j}\Big{|}$ of the separate linear model
    $Y={\beta}_{j}{{X}}_{j}+\widetilde{\beta}_{j}\widetilde{{X}}_{j}+\sum_{j^{\dagger}\notin B_{k}}\left({\beta}_{j^{\dagger}}{{X}}_{j^{\dagger}}+\widetilde{\beta}_{j^{\dagger}}\widetilde{{X}}_{j^{\dagger}}\right)+e,\quad\text{where }j\in B_{k}.$
Although the above feature importance scores can be used to compute feature statistics $W_{1},\ldots,W_{j}$ with the between-group coin-flipping property, they suffer several problems in practice. On one hand, with not much information of $H_{j}^{(\text{fg})}$’s, marginal correlation with response usually lacks power. So do the joint lasso feature importance scores, which do not directly depict evidence against $H_{j}^{(\text{fg})}$’s. On the other hand, with the need to perform $p$ and $K$ lasso linear model fittings, separate lasso and marginal correlation with lasso residual are both computationally intensive.
Faced with such a tradeoff between computational efficiency and power, we propose the combined feature importance scores as computationally efficient proxies of the separate lasso importance scores
$\begin{cases}T_{j}=|g_{j}|\cdot|\widehat{\rho}(X_{j},Y)|,\\ \widetilde{T}_{j}=|g_{j}|\cdot|\widehat{\rho}(\widetilde{{X}}_{j},Y)|,\\ \end{cases}\quad k=1,\ldots,K\text{, }j\in B_{k},$ (7)
where $g_{j}$ is computed as follows.
  1. 1.
    Fit the joint linear model (6) and compute $\widehat{\gamma}_{j}=(\widehat{\beta}_{j}+\widehat{\widetilde{\beta}}_{j})/2$ and $\widehat{\delta}_{j}=(\widehat{\beta}_{j}-\widehat{\widetilde{\beta}}_{j})/2$ as the approximate estimator of the reparametrized model
    $Y=\sum_{j=1}^{p}{\gamma}_{j}\left({{X}}_{j}+\widetilde{{X}}_{j}\right)+\sum_{j=1}^{p}{\delta}_{j}\left({{X}}_{j}-\widetilde{{X}}_{j}\right)+e.$
  2. 2.
    For $k=1,\ldots,K$ and $j\in B_{k}$, compute
    $g_{j}$ $=\widehat{\gamma}_{j}+\sum_{j^{\dagger}\in B_{k}\setminus\{j\}}\frac{\psi_{j,j^{\dagger}}-\boldsymbol{\Psi}_{j,-B_{k}}\boldsymbol{\Psi}^{-1}_{-B_{k},-B_{k}}\boldsymbol{\Psi}_{-B_{k},j^{\dagger}}}{\psi_{j,j}-\boldsymbol{\Psi}_{j,-B_{k}}\boldsymbol{\Psi}^{-1}_{-B_{k},-B_{k}}\boldsymbol{\Psi}_{-B_{k},j}}\widehat{\gamma}_{j^{\dagger}},$
    where $\boldsymbol{\Psi}=(\psi_{j,j^{\dagger}})_{p\times p}=\boldsymbol{\Sigma}-\textbf{D}/2,\quad\boldsymbol{\Sigma}=\text{Var}(\textbf{X}),\quad\textbf{D}=\boldsymbol{\Sigma}-\text{Cov}(\textbf{X},\widetilde{\textbf{X}}),$
    as the approximate estimator of the model
    $Y={g}_{j}\left({{X}}_{j}+\widetilde{{X}}_{j}\right)+\sum_{j^{\prime}\notin B_{k}}{\gamma}_{j^{\prime}}\left({{X}}_{j^{\prime}}+\widetilde{{X}}_{j^{\prime}}\right)+{h}_{j}\left({{X}}_{j}-\widetilde{{X}}_{j}\right)+\sum_{j^{\prime}\notin B_{k}}{\delta}_{j^{\prime}}\left({{X}}_{j^{\prime}}-\widetilde{{X}}_{j^{\prime}}\right)+e.$
    In other words, ${g}_{j}$ measures the overall contribution of ${X}_{j}$ and $\widetilde{{X}}_{j}$ to the variation of $Y$ beyond features from other groups.
By doing so, feature statistics $W_{j}=w_{j}(T_{j},\widetilde{T}_{j})$ ($j=1,\ldots,p$) with antisymmetric functions $w_{j}$’s satisfy the between-group coin-flipping property because
$\text{sign}(W_{j})=\text{sign}(w_{j}(|\widehat{\rho}({{X}}_{j},Y)|,|\widehat{\rho}(\widetilde{{X}}_{j},Y)|)),$
whose between-group coin-flipping property is validated in Theorem 1. In addition, as $g_{j}$ is derived to measure the total dependency between $(X_{j},\widetilde{{X}}_{j})$ and the response $Y$ conditional on $(\textbf{X}_{-B_{k}},\widetilde{\textbf{X}}_{-B_{k}})$, $|W_{j}|$’s computed from the combined feature importance scores provide proxies of $|W_{j}|$’s computed from the separate lasso feature importance scores.
2.3 Naive Feature Filters under Group Knockoffs
By Theorem 1, feature statistics $W_{j}$’s only satisfy the between-group coin-flipping property and thus directly applying the model-X knockoff filter (Candès et al.,, 2018) would violate the FDR control. However, we can still utilize the between-group independence among $W_{j}$’s for inference with FDR control as follows.
Recognizing that independence remains true among $W_{j}$’s for those true $H^{(\text{fg})}_{j}$’s from different groups, we consider aligning $W_{j}$’s in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$. Let $\mathcal{C}_{l}$ be the set of features whose feature statistics are aligned in the $l$-th row and consider the rejection set in the form of $\mathcal{R}^{(\text{fg})}(t)=\{j|W_{j}\geq t\}$.
Table 2: Alignment of feature statistics $W_{j}$’s such that different groups correspond to different columns. Here, $W_{(kl)}$ corresponds to the feature from the $k$-th group that is aligned in the $l$-th row.
Row $B_{1}$ $B_{2}$ $B_{3}$cdots
1 $W_{(11)}$ $W_{(21)}$ $W_{(31)}$cdots
[…]
  • On one hand, we find that within each row, there exists at most one $W_{j}$ from each group. As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within the same row satisfy the coin-flipping property introduced in Candès et al., (2018). Thus, we can estimate the false discovery proportion (FDP) of the rejection subset $\mathcal{R}^{(l)}(t)=\{j\in\mathcal{C}_{l}|W_{j}\geq t\}$ as
    $\widehat{\text{FDP}}^{(l)}(t)=\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t\}}{1\vee\#\{j\in\mathcal{C}_{l}|W_{j}\geq t\}},$ (8)
    for any $t>0$ and $l=1,2\ldots$ in the analogous way of Candès et al., (2018).
  • On the other hand, we have $\text{FDP}^{(\text{fg})}(t)$ of $\mathcal{R}^{(\text{fg})}(t)$ is a weighted mean of $\text{FDP}^{(l)}(t)$’s of subsets $\mathcal{R}^{(l)}(t)=\mathcal{R}^{(\text{fg})}(t)\cap\mathcal{C}_{l}$ ($l=1,2,\ldots$) that
    $\text{FDP}^{(\text{fg})}(t)=$$\frac{\#(\mathcal{R}^{(\text{fg})}(t)\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}$
    $=$ $\sum_{l=1}^{\phi(t)}\frac{\#(\mathcal{R}^{(l)}(t)\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}$
    $=$ $\sum_{l=1}^{\phi(t)}\frac{1\vee\#\mathcal{R}^{(l)}(t)}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}\times\frac{\#(\mathcal{R}^{(l)}(t)\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(l)}(t)}$
    $=$ $\sum_{l=1}^{\phi(t)}\frac{1\vee\#\mathcal{R}^{(l)}(t)}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}\times\text{FDP}^{(l)}(t),$
    where the second equality comes from the fact that rejection subsets $\mathcal{R}^{(l)}(t)$ can be nonempty only for the first $\phi(t)=\max_{k}\#\{j\in B_{k}||W_{j}|\geq t\}$ rows because $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$ with each group $B_{k}$.
As a result, using $\widehat{\text{FDP}}^{(l)}(t)$’s in (8), we can estimate $\text{FDP}^{(\text{fg})}(t)$ of $\mathcal{R}^{(\text{fg})}(t)$ by
$\widehat{\text{FDP}}^{(\text{fg})}(t)$ $=\sum_{l=1}^{\phi(t)}\frac{1\vee\#\mathcal{R}^{(l)}(t)}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}\times\widehat{\text{FDP}}^{(l)}(t)$
$=\sum_{l=1}^{\phi(t)}\frac{1\vee\#\{j\in\mathcal{C}_{l}|W_{j}\geq t\}}{1\vee\#\{j|W_{j}\geq t\}}\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t\}}{1\vee\#\{j\in\mathcal{C}_{l}|W_{j}\geq t\}}$
$=\frac{{\phi(t)}+\#\{j|W_{j}\leq-t\}}{1\vee\#\{j|W_{j}\geq t\}},$
leading to the naive filter (Algorithm 1).
Algorithm 1 Naive feature filter with group knockoffs.
1:  Input: Groups $B_{1},\ldots,B_{K}$, feature statistics $W_{1},\ldots,W_{p}$ and the target level $\alpha>0$.
2:  Compute
$t_{\alpha}=\min\Biggl{\{}t>0\Bigg{|}\widehat{\text{FDP}}^{(\text{fg})}(t)=\frac{{\phi(t)}+\#\{j|W_{j}\leq-t\}}{1\vee\#\{j|W_{j}\geq t\}}\leq\alpha\Biggr{\}},$ (9)
where $\phi(t)=\max_{k}\#\{j\in B_{k}||W_{j}|\geq t\}$.
3:  Output: The rejection set $\mathcal{R}^{(\text{fg})}=\{j|W_{j}\geq t_{\alpha}\}$.
2.4 Rigorous Feature Filters under Group Knockoffs
Although the naive filter seems intuitive in FDR control, such a control can not be proven. To fill the gap, we provide an alternative interpretation of Algorithm 1 and consider a variation that provides provable FDR control as follows.
With the target to reject as many false $H_{j}^{(\text{fg})}$’s as possible while $\text{FDR}^{(\text{fg})}$ is under control, multiple testing of $H_{j}^{(\text{fg})}$’s can be transferred into an optimization problem. Based on the construction of feature statistics that $W_{j}$’s for those false $H_{j}^{(\text{fg})}$’s tend to be positive and large, Algorithm 1 solves the following optimization problem
[…]
$\mathcal{R}^{(\text{fg})}=\cup_{l}\mathcal{R}^{(l)}=\cup_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\},\quad t^{(1)}=t^{(2)}=\cdots.$
Based on such an observation, we utilize theoretical results of Katsevich and Sabatti, (2019) that for any $t>0$,
$\textbf{E}\left[\sup_{t>0}\frac{\#\{j\in\mathcal{H}_{0}^{(\text{fg})}\cap\mathcal{C}_{l}|W_{j}\geq t\}}{1+\#\{j\in\mathcal{H}_{0}^{(\text{fg})}\cap\mathcal{C}_{l}|W_{j}\leq-t\}}\right]\leq 1.93,$
to modify the above optimization problem as
[…]
$\mathcal{R}^{(\text{fg})}=\cup_{l}\mathcal{R}^{(l)}=\cup_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\},$
where $v_{l}$’s are pre-calculated budgets for contribution of $\mathcal{R}^{(l)}$’s to $\widehat{\text{FDP}}^{(\text{fg})}$. By doing so, the rejection set $\mathcal{R}^{(\text{fg})}$ obtained by solving (10) has controlled $\text{FDR}^{(\text{fg})}$ as shown in Theorem 2.
Theorem 2.
The rejection set $\mathcal{R}^{(\text{fg})}$ obtained by solving (10) controls $\text{FDR}^{(\text{fg})}$ at the target level $\alpha>0$.
The proof of Theorem 2 is provided in Section B of the Appendix. Similar to the discussion in Sesia et al., (2020), the correction factor $1.93$ is required in the proof for technical reasons, while empirical evidence suggests that it may be practically unnecessary. To solve the optimization problem (10), we develop the feature-versus-group (FVG) filter (Algorithm 2) in an analogous manner to the GGM knockoff filter (Li and Maathuis,, 2021).
We also provide other variations for FDR control with details in Section C of the Appendix.
Algorithm 2 Feature-versus-group (FVG) filter with group knockoffs.
1:  Input: Groups $B_{1},\ldots,B_{K}$, feature statistics $W_{1},\ldots,W_{p}$ and the target level $\alpha>0$.
2:  Align $W_{j}$’s as shown in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$.
3:  Compute budget $v_{l}$ for the $l$-th row using $|W_{1}|,\ldots,|W_{p}|$ for $l=1,2,\ldots$ such that $\sum_{l=1}^{\infty}v_{l}=1$.
4:  Compute grids $\mathcal{G}_{l}=\{1/v_{l},2/v_{l},\ldots,({neg}_{l}+1)/v_{l}\}$ for the $l$-th row where $neg_{l}=\#\{j\in\mathcal{C}_{l}|W_{j}<0\}$.
5:  Combine $\mathcal{G}_{comb}=(\cup_{l}\mathcal{G}_{l})\cup\{0\}$ and sort values ${grid}_{(1)}>{grid}_{(2)}>\ldots$ in $\mathcal{G}_{comb}$.
6:  Initialize $b=0$
7:  repeat
8:     Update $b=b+1$.
9:     Compute
$t^{(l)}=\min\Biggl{\{}t>0\Bigg{|}\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t\}}{v_{l}}\leq{grid}_{(b)}\Biggr{\}},$ (11)
for $l=1,2,\ldots$.
10:  until $\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\leq\frac{v_{l}\alpha}{1.93}$ for all $l$.
11:  Output: The rejection set $\mathcal{R}^{(\text{fg})}=\cup_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}$.
Remark 1.
In the FVG filter (Algorithm 2), the budgets $v_{l}$’s can be computed adaptively to the prior belief of the number of important features within each groups. Here, we provide two strategies of specifying $v_{l}$’s.
  • If it is believed that all important groups have similar effect size on the response $Y$, we suggest letting
    $v_{l}=\frac{\sum_{k}|W_{(kl)}|}{\sum_{l}\sum_{k}|W_{(kl)}|},\quad l=1,2,\ldots,$
    such that contribution of $\mathcal{R}^{(l)}$’s to $\widehat{\text{FDP}}^{(\text{fg})}$ are constrained by the sum of signal magnitudes $\sum_{k}|W_{(kl)}|$ in the $l$-th row of Table 2.
  • If it is believed that most of important groups have small effect size while a small number of important groups have large effect size, we suggest letting
    $v_{l}=\frac{\sum_{k}|W_{(kl)}|/l}{\sum_{l}\sum_{k}|W_{(kl)}|/l},\quad l=1,2,\ldots,$
    such that more budgets are allocated to the first several rows of Table 2 and more important features in important groups with small effect sizes are identified.
3 Simulated Experiments
To evaluate the proposed FVG filter in both FDR control and identifying important features, we conduct extensive experiments under various scenarios. Specifically, throughout simulated experiments, we generate group knockoffs under the maximum entropy (ME) construction of Chu et al., (2023), compute feature statistics as $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ with the joint lasso importance scores.
3.1 Simulated Experiments with Synthetic Data
To examine how the FVG filter performs under different sample sizes, we simulate 1000 datasets with a fixed number of features $p=250$ for sample sizes $n=500,1000$ and $2000$. Specifically, we partition these $250$ features into $K=50$ groups of equal size $B_{1}=\{X_{1},\ldots,X_{5}\},\ldots,B_{50}=\{X_{246},\ldots,X_{250}\}$. Based on such a partition, we generate for each simulated dataset $n$ feature vectors $\textbf{x}_{1},\ldots,\textbf{x}_{n}$ from the multivariate normal distribution,
$\text{MVN}\left\{\mathbf{0},\begin{pmatrix}0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ 0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ \vdots&\vdots&\ddots&\vdots\\ 0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ \end{pmatrix}\right\}.$ (12)
Given features $\textbf{x}_{1},\ldots,\textbf{x}_{n}$, responses $y_{1},\ldots,y_{n}$ are simulated from the linear model
$Y=\sum_{k=1}^{K}\textbf{X}_{B_{k}}^{\mathsf{T}}\boldsymbol{\beta}_{B_{k}}+e,\quad e\sim\text{N}(0,4^{2}),$ (13)
where
$\boldsymbol{\beta}_{B_{k}}=\begin{cases}(\textbf{1}_{2}^{\mathsf{T}},-0.2\cdot\textbf{1}_{3}^{\mathsf{T}})^{\mathsf{T}},&k=1,2,\\ (2.5,-1.2\cdot\textbf{1}_{2}^{\mathsf{T}},\textbf{0}_{2}^{\mathsf{T}})^{\mathsf{T}},&k=3,4,\\ (0.3\cdot\textbf{1}_{2}^{\mathsf{T}},\textbf{0}_{3}^{\mathsf{T}})^{\mathsf{T}},&k=5,\ldots,10\\ \textbf{0}_{5},&k=11,\ldots,50,\\ \end{cases}$ (14)
Under (12)-(14), we have $H_{j}^{(\text{fg})}$’s are false for $j=51,\ldots,250$ and there are 28 important features. Empirical FDR and power of the FVG filter in identifying important features under different sample sizes are visualized in Figure 1. For comparison, we also evaluate the performance of model-X knockoff filter (Candès et al.,, 2018), group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019) in FDR control and identifying important features.
(a) $n=500$.
(b) $n=1000$.
(c) $n=2000$.
Figure 1: Empirical FDR and power of FVG filter with respect to the target FDR level ($\alpha$) over 1000 simulated datasets of different sample sizes in comparison to model-X knockoff filter (Candès et al.,, 2018), group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019).
It is clear that the FVG filter manages to control $\text{FDR}^{(\text{fg})}$ under any target level $\alpha$. For a fixed target level $\alpha$, we find the empirical power increases to $1$ as $n$ grows, suggesting that the FVG filter is consistent in identifying important features. Although existing methods can also control FDR and have growing power, they suffer different types of performance deterioration in identifying important features. On one hand, group knockoff filter suffers power loss when the target FDR level is lower than $0.10$ due to the threshold issue mentioned in Gimenez and Zou, (2019). Even with power improvement as the target FDR level grows from $0.10$ and $0.20$, the informativeness of rejection sets obtained by group knockoff filter is poor as all important groups ($B_{1},\ldots,B_{10}$) are rejected as a whole without distinguishing which features are more important. As a comparison, the average size of nonempty catching sets $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ ($k=1,\ldots,50$) obtained by FVG filter is less than $2$ under the target level $\alpha=0.1$. On the other hand, although the multilayer knockoff filter has power improvement over model-X knockoff filter by focusing important feature selection on those identified feature groups, there remains a significant power gap between FVG filter and the multilayer knockoff filter when the target FDR is large.
3.2 Simulated Experiments with Real-world Genetic Data
To mimic the dependency structure among features in real-world genetic analysis and investigate how the proposed method performs in real applications, we simulate 500 datasets based on $6095$ variants in the APOE/APOC region (chromosome 19, positions 44909011 $\sim$ 45912650) from the whole-genome sequencing data (NG00067.v5) of the Alzheimer’s Disease Sequencing Project (Leung et al.,, 2019). Specifically, we restrict our study to $6952$ individuals with estimated European ancestry rate [^3] Estimated by SNPWeights v2.1 (Chen et al.,, 2013) using reference populations from the 1000 Genomes Consortium (The 1000 Genomes Project Consortium,, 2015). greater than 80% and $p=1000$ variants $X_{1},\ldots,X_{p}$ whose minor allele frequency (MAF) is larger than $1\%$ and pairwise correlations are in $[-0.95,0.95]$ to avoid the existence of statistically indistinguishable variants with nearly-perfect correlations. Among these variants, we compute the correlation $\text{cor}(X_{i},X_{j})$ between any pair of variants $(X_{i},X_{j})$ over all 6952 individuals and construct variant groups ${B}_{1},\ldots,{B}_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(X_{i},X_{j})|)_{p\times p}$. By randomly sampling $n=5000$ individuals without replacement, each simulated dataset is obtained by collecting the sampled individuals’ variants and generating responses $y_{1},\ldots,y_{n}$ from the linear model
$Y=\sum_{j=1}^{p}X_{j}{\beta}_{j}+e,\quad e\sim\text{N}(0,4),$ (15)
where only $k=10$ randomly selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.225)$. As a result, important features refer to variants $X_{j}$’s whose corresponding coefficients $\beta_{j}$’s are nonzero and important groups are those containing at least one important feature. Specifically, most important groups only contain one important feature.
Figure 2: Empirical performance of the model-X knockoff filter (Candès et al.,, 2018), the group knockoff filter (Dai and Barber,, 2016), the multilayer knockoff filter (Katsevich and Sabatti,, 2019) and our FVG filter over 500 simulated data: (a) empirical FDR and power with respect to the target FDR level; (b) average size of catching sets; (c) average purity of catching sets.
To illustrate the empirical benefit of the FVG filter, we apply it to find important features in these 500 simulated datasets with comparison to the model-X knockoff filter (Candès et al.,, 2018), the group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019). Empirical FDR and power of all methods are presented in Figure 2 (a). For fair comparison, we use the joint lasso feature importance score when implementing both existing methods and our algorithm. It is found that our FVG filter can empirically control FDR. Compared to the model-X knockoff filter which can only distinguish around $50\%$ of important features when the target FDR level is relaxed to $20\%$, our FVG filter, the multilayer knockoff filter and the group knockoff filter are able to identify about $70\%\sim 80\%$ of important features (or feature groups) when the target FDR level is $10\%$. However, the inference results given by the multilayer knockoff filter and the group knockoff filter are far from being informative and precise. To illustrate this, we compute catching set $\mathcal{R}_{k}$ of each variant group from the rejection set $\mathcal{R}$, where
  • $\mathcal{R}_{k}=B_{k}$ if $H^{\text{(gg)}}_{k}$ is rejected by the group knockoff filter;
  • $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ if the rejection set $\mathcal{R}$ is obtained from the model-X knockoff filter or our FVG filter.
  • For each group $B_{k}$, we let
    • $\mathcal{R}_{k}=B_{k}$ if $H^{\text{(gg)}}_{k}$ but all $H^{\text{(ff)}}_{j}$’s ($j\in B_{k}$) is rejected by the multilayer knockoff filter;
    • $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ if $H^{\text{(gg)}}_{k}$ and some $H^{\text{(ff)}}_{j}$’s ($j\in B_{k}$) are rejected by the multilayer knockoff filter.
We then calculate the average size and purity (minimum absolute correlation) of catching sets to measure the informativeness and precision of $\mathcal{R}$ given by different filters in Figure 2 (b)-(c). Here, it is desired that sizes of catching sets are smaller as more unimportant variants are excluded. Even when there remain multiple features in catching sets, high purity is desirable to exclude unimportant variants weakly correlated to the important ones (or poor-quality proxy variants). In contrast to the group knockoff filter whose average size and purity of catching sets are almost $6$ and smaller than $20\%$ respectively (which means that catching sets obtained by the group knockoff filter contain a lot of poor-quality proxy variants), the precision of our FVG filter in pinpointing true important variants is close to the model-X knockoff filter and is higher than the multilayer knockoff filter. With an average size close to $1$ and average purity greater than $80\%$, our FVG filter can directly obtain informative catching sets containing only one or two features in most cases. In summary, our FVG filter manage to simultaneously achieve high power and high informativeness in identifying important features.
4 AD Analysis of EADB–UKBB Dataset
Given the promising performance of the proposed filter in simulated experiments, we apply our method to the EADB-UKBB dataset where model-X knockoff filter exhibits power loss and the group knockoff filter returns not informative enough results in Section 1.2.
To infer which variants are associated with AD, we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ and construct variant groups $B_{1},\ldots,B_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$. By doing so, we obtain the same $321,535$ groups $B_{1},\ldots,B_{k}$ as the group knockoff filter in Section 1.2 where the average absolute correlation of variants from different groups are lower than 0.5. Based on the correlation matrix $(\text{cor}(G_{i},G_{j}))_{p\times p}$ and variant groups $B_{1},\ldots,B_{K}$, we generate group knockoffs $\widetilde{\mathbb{G}}$ under the maximum entropy (ME) construction of Chu et al., (2023). Given the response of interest $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ where $y_{i}=1$ if the $i$-th observation corresponds to a clinically diagnosed AD case and $y_{i}=0$ otherwise, we use absolute values of lasso estimators of the logistic regression model,
$\text{logit}\{\text{Pr}(Y=1|{\textbf{G}},\widetilde{\textbf{G}})\}=\beta_{0}+\sum_{j=1}^{p}\left\{{G}_{j}{\beta}_{j}+\widetilde{G}_{j}\widetilde{\beta}_{j}\right\}$
as feature importance scores. In other words, we compute $T_{j}=\left|\widehat{\beta}_{j}\right|$ and $\widetilde{T}_{j}=\left|\widehat{\widetilde{\beta}}_{j}\right|$, which are also used in other knockoff-based methods and provide more informative rejection sets without large power loss given the large sample size. We then calculate feature statistics $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ and implement the FVG filter with target FDR level $\alpha=0.10$. Variants in the rejection set $\mathcal{R}^{(\text{fg})}$ under target FDR level $\alpha=0.10$ are presented in Figure 3 with their post-selection importance scores $W_{j}\cdot\text{I}(W_{j}>0)$ and names of the closest genes.
Figure 3: Manhattan plot of genetic variants under our FVG filter. Identified variants under target FDR level $\alpha=0.10$ are highlighted with names of their closest genes.
Details of variants identified by the FVG filter under the target FDR level $0.10$ are provided in Table 3. In total, we identify 205 variants that contribute to AD variation substantially, which comes from 154 groups in 84 loci. Similar to the literature, multiple variants are identified in the APOE/APOC region with the strongest association to AD (chromosome 19, positions 44909011 $\sim$ 45912650). In addition, we also manage to identify variants rs744373 and rs2118506 close to gene “BIN1”, variant rs1532278 close to gene “CLU”, variant rs10792832 close to gene “PICALM”, variants rs4277405, rs4309 and rs3730025 close to gene “ACE”, all of which are also reported in He et al., (2022) and Bellenguez et al., (2022). Compared with the model-X knockoff filter, the FVG filter only misses 2 loci identified by the model-X knockoff filter but discover 37 more loci, which suggests that our method would not suffer from the same power loss issue of the model-X knockoff filter.
Table 3: Details of variants identified by the FVG filter under the EADB–UKBB dataset and target FDR level $\alpha=0.10$.
Identified Position Closest Identified Position Closest Identified Position Closest Identified Position Closest
Variant Gene Variant Gene Variant Gene Variant Gene
Chromosome 1 rs3997700 41251889 TREML5P rs9516168 93085490 LINC00363 rs3852860 44879709 NECTIN2
rs11579514 161173522 B4GALT3 rs9296558 47484147 CD2AP Chromosome 14 rs365653 44858389 NECTIN2
rs782791 203029691 PPFIA4 rs10948363 47520026 CD2AP rs74093831 105733666 ELK2AP rs73052307 44881148 NECTIN2
rs4562624 207512620 CR1 rs12525381 696327 PTPRK rs873533 106667442 LINC00221 rs11668327 44895376 TOMM40
rs3818361 207611623 CR1 rs117591349 78478451 MIR10524 rs17125924 52924962 FERMT2 rs449647 44905307 APOE
rs6701713 207612944 CR1 rs313161 85332752 LINC02535 rs4904929 92470949 SLC24A4 rs10420036 44882772 NECTIN2
rs4844610 207629207 CR1 Chromosome 7 Chromosome 15 rs283813 44885917 NECTIN2
Chromosome 2 rs221834 100745552 ZAN rs59685680 50709337 SPPL2A rs7254892 44886339 NECTIN2
rs116038905 105805908 NCK2 rs144867634 111940111 DOCK4 rs28470808 50735410 SPPL2A rs7412 44908822 APOE
rs74851408 117201103 DDX18 rs10952168 1554792 TMEM184A rs2043085 58388755 LIPC rs445925 44912383 APOC1
rs6733839 127135234 BIN1 rs56402156 143406388 TCAF2 rs593742 58753575 ADAM10 rs390082 44913574 APOC1
rs730482 127136908 BIN1 rs12703526 143410495 TCAF2 rs117618017 63277703 APH1B rs190712692 44921921 APOC1
rs744373 127137039 BIN1 rs9640386 143435090 TCAF2 rs3816385 64126456 SNX1 rs41290108 44874585 NECTIN2
rs745717 127136321 BIN1 rs2189965 28132395 JAZF1 rs56117790 79075361 RASGRF1 rs73033507 44928146 APOC1P1
rs2118506 127139927 BIN1 rs2122623 37842715 NME8 Chromosome 16 rs79701229 44881674 NECTIN2
rs10496731 134840058 ACMSD rs7809394 55077689 EGFR rs12325539 30022312 C16orf92 rs157580 44892009 TOMM40
rs4954187 134842848 ACMSD rs12669393 7815748 UMAD1 rs11865499 31120929 KAT8 rs2075649 44892073 TOMM40
rs3819121 134865290 ACMSD rs4725076 8204779 ICA1 rs8058370 81738205 CMIP rs8106922 44898409 TOMM40
rs6725887 202881162 WDR12 rs1476679 100406823 ZCWPW1 rs12446759 81739398 CMIP rs405697 44901434 TOMM40
rs10933431 233117202 INPP5D rs41280968 100105324 AP4M1 rs11548656 81883307 PLCG2 rs7259620 44904531 TOMM40
rs7421448 233117495 INPP5D rs13228694 100342684 STAG3L5P-PVRIG2P-PILRB rs1071644 81937798 PLCG2 rs440446 44905910 APOE
rs35349669 233159830 INPP5D Chromosome 8 rs56407236 90103687 FAM157C rs6857 44888997 NECTIN2
rs3771371 71400409 ZNF638 rs2436889 102564430 ODF1 Chromosome 17 rs2075650 44892362 TOMM40
rs113068192 76786226 LRRTM4 rs4731 11808828 FDFT1 rs79759095 44451394 GPATCH8 rs157581 44892457 TOMM40
Chromosome 3 rs79832570 144042819 SPATC1 rs5910 44372421 ITGA2B rs34404554 44892652 TOMM40
rs7634492 136105288 PPP2R3A rs61732533 144053248 OPLAH rs199439 46716137 NSF rs75627662 44910319 APOE
rs16824536 155069722 MME rs34674752 144099319 SHARPIN rs199451 46724418 NSF rs438811 44913484 APOC1
rs61762319 155084189 MME rs73223431 27362470 PTK2B rs2671659 49397167 PHB rs283814 44885967 NECTIN2
rs4974180 56200107 ERC2 rs1532278 27608798 CLU rs616338 49219935 ABI3 rs439401 44911194 APOE
Chromosome 4 rs9297949 94957217 NDUFAF6 rs2072153 49312652 ZNF652 rs59325138 44913034 APOC1
rs6448453 11024404 CLNK Chromosome 9 rs2526378 58326988 TSPOAP1-AS1 rs584007 44913221 APOC1
rs13116567 12239152 LINC02270 rs1131773 93077974 SUSD3 rs2632516 58331728 TSPOAP1-AS1 rs157595 44922203 APOC1
rs502746 14019593 LINC01182 Chromosome 10 rs9896864 63458947 CYB561 rs4803770 44924096 APOC1P1
rs116791081 143428212 GAB1 rs7920721 11678309 ECHDC3 rs4277405 63471557 ACE rs71352239 44926286 APOC1P1
Chromosome 5 rs7068614 11679478 ECHDC3 rs4309 63482562 ACE rs118060185 44928417 APOC1P1
rs25992 14707491 ANKH rs10764258 20634336 MIR4675 rs3730025 63480412 ACE rs144311893 44920687 APOC1
rs871269 151052827 TNIP1 rs117980033 29966853 JCAD rs9901675 7581494 CD68 rs114533385 44933496 APOC1P1
rs335473 177559423 FAM193B rs1227759 69831131 COL13A1 Chromosome 19 rs79429216 44942260 APOC4-APOC2
rs4700716 180138792 RASGEF1C rs6586028 80494228 TSPAN14 rs445752 43956053 ZNF221 rs12721109 44943964 APOC4-APOC2
rs10068419 86923485 LINC02059 Chromosome 11 rs2571174 44430314 ZNF229 rs16979595 44974124 CLPTM1
Chromosome 6 rs74685827 121482368 SORL1 rs2293164 44475997 ZNF180 rs10411314 45224801 EXOC3L2
rs976271 114361563 HDAC2-AS2 rs2298813 121522975 SORL1 rs1625502 44540845 CEACAM22P rs75987583 45745299 MEIOSIN
rs2130357 27919052 OR2B2 rs3781837 121578263 SORL1 rs10405071 44732632 BCL3 rs875121 49948251 SIGLEC11
rs13408 28244970 ZKSCAN4 rs3740890 130385218 ZBTB44-DT rs2965169 44747899 BCL3 rs1710398 51223655 CD33
rs3130931 31167111 POU5F1 rs17439299 15599040 LINC02751 rs116480598 44753996 BCL3 rs3865444 51224706 CD33
rs3132451 31614248 AIF1 rs1685404 47222114 DDB2 rs28399654 44813331 BCAM rs12974390 54264111 LILRB5
rs532965 32610196 HLA-DRB1 rs2293579 47419207 PSMC3 rs10419669 44794812 CBLC rs10404195 5908958 VMAC
rs9469112 32447376 HLA-DRA rs1582763 60254475 MS4A4E rs10405693 44823407 BCAM rs4147918 1058177 RDH13
rs9271162 32609938 HLA-DRB1 rs11824773 60309467 MS4A4A rs73572039 44839373 NECTIN2 19:1051137_CTG_C 1051138 RDH13
rs9273349 32658092 HLA-DQB1 rs17854357 65834089 SNX32 rs77301115 44893716 TOMM40 rs3752246 1056493 RDH13
rs9275184 32686937 HLA-DQB1 rs536841 86076782 PICALM rs112849259 44894050 TOMM40 rs35917007 1849148 REXO1
rs9275614 32716480 HLA-DQA2 rs10792832 86156833 PICALM rs79398853 44895528 TOMM40 Chromosome 20
rs114812713 41066261 OARD1 rs3844143 86139201 PICALM rs75687619 44896087 TOMM40 rs6064392 56409712 CASS4
rs7748513 41160234 TREM2 Chromosome 12 rs76366838 44896639 TOMM40 rs6069749 56446519 CASS4
rs2093395 41187288 TREML2 rs17860282 51346566 CELA1 rs115881343 44899959 TOMM40 Chromosome 21
rs143332484 41161469 TREM2 Chromosome 13 rs3865427 44877704 NECTIN2 rs62222965 26225443 APP
rs9394778 41247320 TREML5P
Figure 4: Comparison of our FVG filter and the group knockoff filter in analyzing the EADB–UKBB dataset: (a) Venn diagram of variant groups identified by different filters; (b) bar plot of size of catching sets identified by different filters (columns); and (c) histogram of purity of catching sets identified by different filters.
Compared to the group knockoff filter which identifies 152 AD-associated variant groups from 84 loci under the same target FDR level $0.10$, our FVG filter does not suffer power loss. As shown in Figure 4 (a), among all variant groups identified by either the group knockoff filter or our FVG filter [^4] Here, we say a variant group is identified by our FVG filter if it contains at least one variant identified by our FVG filter ., 143 groups are identified by both methods while 11 groups and 9 groups are only identified by our filter and the group knockoff filter respectively. This suggests the consistency of inference results between our FVG filter and the group knockoff filter. However, our filter manages to detect signals at variant rs1532278 close to gene “CLU”, while none of variant groups in the “CLU” locus are identified by the group knockoff filter. Without significant power loss, our method is shown to succeed in pinpointing AD-associated variants with higher precision. As shown in Figure 4 (b)-(c), all catching sets obtained by our method are of size no larger than $6$ while most of catching sets (129/154) are of size 1 and purity greater than $90\%$. In other words, most of catching sets only contain the important variants or strongly correlated proxy variants. In contrast, 37 catching sets obtained by the group knockoff filter (Dai and Barber,, 2016) contain more than $8$ variants and are far from being informative with strong impurity. Specifically, for the five largest catching sets obtained by the group knockoff filter shown in Table 4, our FVG filter manages to distinguish one or two variants from more than 100 proxy variants in each. This suggests the ability of the proposed method in pinpointing AD-associated variants. Comprehensive comparisons of our method with all existing methods are presented in Table 5.
Table 4: List of variants identified by the FVG filter in the five largest catching sets obtained by the group knockoff filter.
Identified Chromosome Position Catching sets obtained by Closest
Variant the group knockoff filter (size) Gene
rs13408 6 28244970 6:28017819:28917607:9 (121) ZKSCAN4
rs3130931 6 31167111 6:30798168:31571217:20 (136) POU5F1
rs3132451 6 31614248 6:31571218:32682663:7 (232) AIF1
rs9469112 6 32447376 6:31571218:32682663:76 (171) HLA-DRA
rs199439 17 46716137 17:43056905:45876021:36 (314) NSF
rs199451 17 46724418
Table 5: Summary of results by applying different methods on the EADB-UKBB dataset.
Method Number of Average number of Average Average
identified variants per size of purity of
loci identified locus catching sets catching sets
Marginal association test 51 18.078 18.078 0.477
Model-X knockoff filter 49 2.816 1.000 1.000
Group knockoff filter 84 23.714 13.105 0.610
FVG filter 84 2.440 1.257 0.958
5 Discussions
As more genetic variants are sequenced using the rapidly developing whole-genome sequencing technology, it is now possible to explain more diseases’ heritability via large-scale genetic studies. Noting that existing knockoff methods either lose power or lack informativeness because of strong correlations among variants in Alzheimer’s disease analysis of the EADB-UKBB dataset, we developed a new filter to select important features using group knockoffs. Specifically, we first define a new family of conditional independence hypotheses that allow inference at feature level under group knockoffs construction. Based on theoretical properties of test statistics when group knockoffs are used, we develop the FVG knockoff filter with FDR control. Extensive simulated experiments (with both synthetic data and real-world genetic data) empirically validate the FDR control of our proposed filter. Compared with the existing knockoff filters, our FVG filter is shown to be powerful in identifying important features with higher precision. When applying the proposed method to AD analysis of the real-world EADB–UKBB dataset, we identify as many signals as the group knockoff filter with as high precision as the model-$X$ knockoff filter at the same target FDR level $0.10$. Within several variant groups identified by the group knockoff filter, our method precisely distinguishes the variants that contribute more to AD variation.
Simultaneous inference of conditional independence at different layers of groups is of great necessity as we need to balance the power and the resolution (informativeness) of groups (Katsevich and Sabatti,, 2019; Sesia et al.,, 2020; Gablenz and Sabatti,, 2023). For example, in genetic analysis, variant groups with higher resolution are typically smaller and have higher correlations, making it challenging to identify important signals, but easier to interpret if they are identified. Thus, it is interesting to incorporate the proposed filter into the multilayer testing framework (Barber and Ramdas,, 2017; Katsevich and Sabatti,, 2019; Sesia et al.,, 2020; Gablenz and Sabatti,, 2023) to perform simultaneous inference of $H^{(\text{fg})}$’s at multiple layers of group structure. In addition, it is also of great interest to adopt the proposed filter in causal inference. By elaborately designing group knockoffs that can eliminate confounding effects, the proposed filter can perform causal features selection with provable FDR control when unmeasured confounders exist. We leave these tasks as future works.
Appendix A Proof of Theorem 1
Without loss of generality, we let $k=1$, $k^{\dagger}=2$, $j\in B_{1}$ and $j^{\dagger}\in B_{2}$.
  • $\star$
    (Uniformity):
    Since $H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ where $j\in B_{k}$, if $j\in\mathcal{H}_{0}^{(\text{fg})}$, we have $H_{1}^{\text{(gg)}}$ is true (and $H_{j^{\prime}}^{(\text{fg})}$ is true for all $j^{\prime}\in B_{k}$) and $(X_{j},\textbf{X}_{B_{1}\setminus\{j\}})\perp Y|\textbf{X}_{-B_{1}}$ by (2). Because $(\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}},\widetilde{\textbf{X}}_{-B_{1}})\perp Y|(X_{j},\textbf{X}_{B_{1}\setminus\{j\}},\textbf{X}_{-B_{1}})$, we have
    $(X_{j},\textbf{X}_{B_{1}\setminus\{j\}},\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}},\widetilde{\textbf{X}}_{-B_{1}})\perp Y|\textbf{X}_{-B_{1}},$
    and thus
    $(X_{j},\textbf{X}_{B_{1}\setminus\{j\}},\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}})\perp Y|(\textbf{X}_{-B_{1}},\widetilde{\textbf{X}}_{-B_{1}}).$ (16)
    By (4), we have conditional on $(\textbf{X}_{-B_{1}},\widetilde{\textbf{X}}_{-B_{1}})$,
    $(X_{j},\textbf{X}_{B_{1}\setminus\{j\}},\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}}){\displaystyle\mathop{=\joinrel=}^{D}}(\widetilde{X}_{j},\widetilde{\textbf{X}}_{B_{1}\setminus\{j\}},X_{j},\textbf{X}_{B_{1}\setminus\{j\}}).$ (17)
    Thus, (16)-(17) lead to
    $([\mathbb{X}_{j},\mathbb{X}_{B_{1}\setminus\{j\}},\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{1}\setminus\{j\}},\{\mathbb{X}_{-B_{1}},\widetilde{\mathbb{X}}_{-B_{1}}\}],\textbf{y})$ (18)
[…]
    Note that swapping $\mathbb{X}_{j},\mathbb{X}_{B_{1}\setminus\{j\}}$ and $\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{1}\setminus\{j\}}$ does not change $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s, (5) and (18) implies that conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s,
    $(T_{j},\widetilde{T}_{j}){\displaystyle\mathop{=\joinrel=}^{D}}(\widetilde{T}_{j},T_{j}),$ (19)
[…]
  • $\star$
    (Between-group Independence):
    By (4), we have
    $({\textbf{X}}_{B_{1}},\widetilde{\textbf{X}}_{B_{1}},{\textbf{X}}_{B_{2}},\widetilde{\textbf{X}}_{B_{2}})|(\textbf{X}_{-(B_{1}\cup B_{2})},\widetilde{\textbf{X}}_{-(B_{1}\cup B_{2})})$
[…]
    If $j,j^{\dagger}\in\mathcal{H}_{0}^{(\text{fg})}$, by (5) and (19), we have conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s,
    $({T}_{j},\widetilde{T}_{j},{T}_{j^{\dagger}},\widetilde{T}_{j^{\dagger}}){\displaystyle\mathop{=\joinrel=}^{D}}({T}_{j},\widetilde{T}_{j},\widetilde{T}_{j^{\dagger}},{T}_{j^{\dagger}}){\displaystyle\mathop{=\joinrel=}^{D}}(\widetilde{T}_{j},{T}_{j},{T}_{j^{\dagger}},\widetilde{T}_{j^{\dagger}}){\displaystyle\mathop{=\joinrel=}^{D}}(\widetilde{T}_{j},{T}_{j},\widetilde{T}_{j^{\dagger}},{T}_{j^{\dagger}}).$
    and thus $\text{\rm sign}(W_{j})$ and $\text{\rm sign}(W_{j^{\dagger}})$ are independent conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s.
Appendix B Proof of Theorem 2
As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within each row of Table 2 satisfy the coin-flipping property introduced in Candès et al., (2018). In other words, for each $l$, $\text{sign}(W_{j})$’s for those $j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}$ are i.i.d. coin flips conditional on $\{W_{j}|j\in\mathcal{C}_{l}\}$ and $\{\text{\rm sign}(W_{j})|j\in\mathcal{C}_{l}\setminus\mathcal{H}_{0}^{(\text{fg})}\}$. Thus, by Lemma 3 of Katsevich and Sabatti, (2019), we have
$\textbf{E}\left\{\sup_{t>0}\frac{V^{(l)}_{\mathcal{H}_{0},+}(t)}{1+V^{(l)}_{\mathcal{H}_{0},-}(t)}\right\}\leq 1.93,\quad\text{where }\begin{cases}V^{(l)}_{\mathcal{H}_{0},+}(t)=\#\{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}|W_{j}\geq t\},\\ V^{(l)}_{\mathcal{H}_{0},-}(t)=\#\{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}|W_{j}\leq-t\}.\\ \end{cases}$ (20)
Thus, the FDR of the rejection set $\mathcal{R}^{(\text{fg})}$ obtained by solving (10) satisfies
${\text{FDR}}^{(\text{fg})}=$ $\mathbf{E}\Biggl{\{}\frac{\#(\mathcal{R}^{(\text{fg})}\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}}\Biggr{\}}$
$=$ $\mathbf{E}\Biggl{\{}\frac{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\Biggr{\}}$
$=$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot\frac{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
$\leq$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
$\leq$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\frac{v_{l}\alpha}{1.93}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
$\leq$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\frac{v_{l}\alpha}{1.93}\cdot\sup_{t>0}\frac{V^{(l)}_{\mathcal{H}_{0},+}(t)}{1+V^{(l)}_{\mathcal{H}_{0},-}(t)}\Biggr{\}}$
$=$ $\sum_{l=1}^{\infty}\frac{v_{l}\alpha}{1.93}\cdot\mathbf{E}\Biggl{\{}\sup_{t>0}\frac{V^{(l)}_{\mathcal{H}_{0},+}(t)}{1+V^{(l)}_{\mathcal{H}_{0},-}(t)}\Biggr{\}}$
$=$ $\sum_{l=1}^{\infty}{v_{l}\alpha}$
$=$ $\alpha,$ (21)
by (20).
Appendix C Extensions
C.1 Multiple Knockoffs
Although Algorithm 2 with group knockoffs can obtain the rejection set with FDR control, randomness in group knockoffs generation could produce greatly different $\mathcal{R}^{(\text{fg})}$ in different runs, especially in the case that the number of false $H_{j}^{(\text{fg})}$’s is small and Algorithm 2 could return an empty rejection set (Gimenez and Zou,, 2019). Here we extend the proposed Algorithm 2 to multiple group knockoffs $\widetilde{\textbf{X}}^{(1)},\ldots,\widetilde{\textbf{X}}^{(M)}$ where original and knockoff features are simultaneously exchangeable at the level of feature groups. That is to say, with the convention that $\textbf{X}=\widetilde{\textbf{X}}^{(0)}$, $\widetilde{\textbf{X}}^{(0)},\widetilde{\textbf{X}}^{(1)},\ldots,\widetilde{\textbf{X}}^{(M)}$ satisfy
$(\widetilde{\textbf{X}}^{(\sigma_{1}(0))}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{K}(0))}_{B_{K}},\widetilde{\textbf{X}}^{(\sigma_{1}(1))}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{K}(1))}_{B_{K}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{1}(M))}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{K}(M))}_{B_{K}})$ (22)
[…]
for any permutations $\sigma_{1},\ldots,\sigma_{K}$ of $\{0,1,\ldots,M\}$. With importance scores $\{T_{j}^{(m)}|j=1,\ldots,p;m=0,\ldots,M\}$ obtained in an analogous way of (5), we follow He et al., (2021) to compute feature statistics
$\kappa_{j}=\arg\max_{m}T_{j}^{(m)},\quad\tau_{j}=\max_{m}T_{j}^{(m)}-\text{\rm median}\{T_{j}^{(m)}|m\neq\kappa_{j}\},\quad j=1,\ldots,p.$ (23)
Specifically, $\kappa_{j}$ and $\tau_{j}$ are the multiple group knockoffs counterparts of $\text{sign}(W_{j})$ and $|W_{j}|$ respectively (Gimenez and Zou,, 2019) that
  1. $\star$
    conditional on $\tau_{1},\ldots,\tau_{p}$ and $\kappa_{j}$’s for those false $H^{(\text{fg})}_{j}$’s,
    1. $\diamond$
      (Uniformity) $\kappa_{j}$ uniformly distributes on $\{0,1,\ldots,M\}$ for all $j\in\mathcal{H}_{0}^{(\text{fg})}$;
    2. $\diamond$
      (Between-Group Independence) for any $k\neq k^{\dagger}$, $\kappa_{j}$ and $\kappa_{j^{\dagger}}$ are independent for any $j\in B_{k}\cap\mathcal{H}_{0}^{(\text{fg})}$ and $j^{\dagger}\in B_{k^{\dagger}}\cap\mathcal{H}_{0}^{(\text{fg})}$.
Based on $\kappa_{j}$’s and $\tau_{j}$’s, we provide Algorithm 3 as the multiple knockoffs counterpart of Algorithm 2.
Algorithm 3 Feature filter with multiple group knockoffs (naive version).
1:  Input: Groups $B_{1},\ldots,B_{K}$, feature statistics $\{(\kappa_{j},\tau_{j})|j=1,\ldots,p\}$ and the target level $\alpha>0$.
2:  Align $(\kappa_{j},\tau_{j})$’s as shown in Table 6 such that $(\kappa_{j},\tau_{j})$’s for all features in the group $B_{k}$ are in the $k$-th column and $\tau_{(k1)}\geq\tau_{(k2)}\geq\cdots$.
3:  Compute budget $v_{l}$ for the $l$-th row using $\tau_{1},\ldots,\tau_{p}$ for $l=1,2,\ldots$ such that $\sum_{l=1}^{\infty}v_{l}=1$.
4:  Compute grids $\mathcal{G}_{l}=\{1/v_{l},2/v_{l},\ldots,({neg}_{l}+1)/v_{l}\}$ for the $l$-th row where $neg_{l}=\#\{j\in\mathcal{C}_{l}|\kappa_{j}\neq 0\}$.
5:  Combine $\mathcal{G}_{comb}=(\cup_{l}\mathcal{G}_{l})\cup\{0\}$ and sort values ${grid}_{(1)}>{grid}_{(2)}>\ldots$ in $\mathcal{G}_{comb}$.
6:  Initialize $b=0$
7:  repeat
8:     Update $b=b+1$.
9:     Compute
$t^{(l)}=\min\Biggl{\{}t>0\Bigg{|}\frac{1+\#\{j\in\mathcal{C}_{l}|\kappa_{j}\neq 0,\tau_{j}\geq t\}}{v_{l}}\leq{grid}_{(b)}\Biggr{\}},$ (24)
for $l=1,2,\ldots$.
10:  until $\frac{\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)}{M}\times\frac{1+\#\{j\in\mathcal{C}_{l}|\kappa_{j}\neq 0,\tau_{j}\geq t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|\kappa_{j}=0,\tau_{j}\geq t^{(l)}\}]}\leq\frac{v_{l}\alpha}{1.93}$ for all $l$.
11:  Output: The rejection set $\mathcal{R}^{(\text{fg})}=\cup_{l}\{j\in\mathcal{C}_{l}|\kappa_{j}=0,\tau_{j}\geq t^{(l)}\}$.
Table 6: Alignment of feature statistics $(\kappa_{j},\tau_{j})$’s such that different groups correspond to different columns. Here, $(\kappa_{(kl)},\tau_{(kl)})$ corresponds to the feature from the $k$-th group that is aligned in the $l$-th row.
Row $B_{1}$ $B_{2}$ $B_{3}$cdots
1 $(\kappa_{(11)},\tau_{(11)})$ $(\kappa_{(21)},\tau_{(21)})$ $(\kappa_{(31)},\tau_{(31)})$cdots
[…]
C.2 Filters based on E-values
Another variation relies on the connection between knockoff filter and $e$-values discussed in Ren and Barber, (2024). Following Ren and Barber, (2024)’s way of computing $e$-values of a set of hypotheses whose $W_{j}$’s possess the coin-flipping property, we can analogously compute $e$-values of hypotheses whose $W_{j}$’s are aligned in the same row of Table 2 as shown in steps 4-7 of Algorithm 4. By doing so, an e-value $e_{j}$ is calculated for each hypothesis $H_{j}^{(\text{fg})}$ and we obtain the rejection set $\mathcal{R}^{(\text{fg})}$ by applying the e-BH procedure (Wang and Ramdas,, 2022) to $e$-values $e_{1},\ldots,e_{p}$ (as shown in steps 8-10 of Algorithm 4).
Algorithm 4 Rigorous feature filter with group knockoffs using e-values.
1:  Input: Groups $B_{1},\ldots,B_{K}$, feature statistics $W_{1},\ldots,W_{p}$ and the target level $\alpha>0$.
2:  Align $W_{j}$’s as shown in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$.
3:  Compute the number of feature $p_{l}$ in the $l$-th row ($l=1,2,\ldots$).
4:  for $l=1,2,\ldots$ do
5:     Compute
$t_{l,\alpha}=\min\Biggl{\{}t_{l}>0\Bigg{|}\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t_{l}\}}{1\vee\#\{j\in\mathcal{C}_{l}|W_{j}\geq t_{l}\}}\leq\alpha/2\Biggr{\}}.$
6:     Compute e-value $e_{j}$ for each $j\in\mathcal{C}_{l}$ as
$e_{j}=\frac{p_{l}\times I(W_{j}\geq t_{l,\alpha})}{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t_{l,\alpha}\}}.$
7:  end for
8:  Compute $e_{(1)}\geq\ldots\geq e_{(p)}$ as the order statistics of $e_{1},\ldots,e_{p}$.
9:  Compute $\widehat{J}=\max\{J|e_{(J)}\geq p/(\alpha J)\}$ or $\widehat{J}=0$ if $e_{(J)}<p/(\alpha J)$ for $J=1,\ldots,p$.
10:  Output: The rejection set $\mathcal{R}^{(\text{fg})}=\{j|e_{j}\geq e_{(\widehat{J})}\}$.
Theorem 3.
The rejection set $\mathcal{R}^{(\text{fg})}$ obtained by Algorithm 4 controls $\text{FDR}^{(\text{fg})}$ at the target level $\alpha>0$.
Proof of Theorem 3.
Because $W_{j}$’s that are in the same row of Table 2 possess the coin-flipping property, by (7) of Ren and Barber, (2024), we have
$\textbf{E}\left\{\frac{\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\geq t_{l,\alpha})}{1+\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\leq-t_{l,\alpha})}\right\}\leq 1.$
As a result, we have for $l=1,2,\ldots,$
$\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}=$ $p_{l}\times\textbf{E}\left\{\frac{\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\geq t_{l,\alpha})}{1+\sum_{j\in\mathcal{C}_{l}}I(W_{j}\leq-t_{l,\alpha})}\right\}$
$\leq$ $p_{l}\times\textbf{E}\left\{\frac{\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\geq t_{l,\alpha})}{1+\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\leq-t_{l,\alpha})}\right\}$
$\leq$ $p_{l},$
and thus
$\sum_{j\in\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}=\sum_{l=1}^{\infty}\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}\leq\sum_{l=1}^{\infty}p_{l}=p.$
Finally, by Theorem 2 of of Ren and Barber, (2024), as $\sum_{j\in\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}\leq p$, the e-BH procedure in steps 8-10 of Algorithm 4 returns the rejection set $\mathcal{R}^{(\text{fg})}$ with FDR control. ∎

### v3

@@
-Pinpointing Important Genetic Variants via A Feature-level Filter and Group Knockoffs
+Refining Genetic Discoveries of Group Knockoffs via A Feature-level Filter
 Abstract
 Identifying variants that carry substantial information on the trait of interest remains a core topic in genetic studies.
-In analyzing the EADB-UKBB dataset to identify genetic variants associated with Alzheimer’s disease (AD), however, we recognize that both existing marginal association tests and conditional independence tests using knockoffs suffer either power loss or lack of informativeness, especially when strong correlations exist among variants.
-To address the limitations of existing knockoff filters, we propose a new feature-versus-group (FVG) filter that is more powerful and precise in identifying important features from a set of strongly correlated features using group knockoffs.
-In extensive simulation studies, the FVG filter controls the expected proportion of false discoveries and identifies important features with enhanced power and greater precision.
-Applying the proposed method to the EADB-UKBB dataset, we discover important variants from 84 loci (same as the most powerful group knockoff filter) with catching sets of substantially smaller size and higher purity.
-Keywords: Alzheimer’s disease genetics, False discovery rate, Feature selection, Knockoffs.
+In analyzing the EADB-UKBB dataset to identify genetic variants associated with Alzheimer’s disease (AD), however, we recognize that both existing marginal association tests and conditional independence tests using existing knockoff filters suffer either power loss or lack of informativeness, especially when strong correlations exist among variants.
+To address these limitations, we propose a new feature-versus-group (FVG) filter that achieves balance between the power and precision in identifying important features from a set of strongly correlated features using group knockoffs.
+In extensive simulation studies, the FVG filter controls the expected proportion of false discoveries and identifies important features in smaller catching sets without large power loss.
+Applying the proposed method to the EADB-UKBB dataset, we discover important variants from 89 loci (similar to the most powerful group knockoff filter) with catching sets of substantially smaller size and higher purity and verify the biological informativeness of our discoveries.
+Keywords: Alzheimer’s disease genetics, False discovery rate, Genetic variant selection, Knockoffs.
 1 Introduction
 1.1 Conditional Independence Test to Discover Disease-associated Variants
-Testing conditional independence between a set of features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the response of interest $Y$ is an important topic in various research areas, including causal inference (Peters,, 2015; Cai et al.,, 2022), genetic analysis (Khera and Kathiresan,, 2017; Zhu et al.,, 2018) and graphical model learning (Deka et al.,, 2016; Tugnait,, 2022).
+Testing conditional independence between a set of features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the response of interest $Y$ is an important topic in various research areas, including causal inference (Peters2015; Cai2022), genetic analysis (Khera2017; Zhu2018) and graphical model learning (Deka2016; Tugnait2022).
 In the era of big data, as the number of features increases, the need for statistical approaches for simultaneous inference of conditional independence between hundreds of thousands of features and the response keeps emerging.
-Specifically, developing statistical methods for conditional independence tests with type-I error rate control has been an important and popular research topic of large-scale genome-wide association studies (GWAS; Tang and He,, 2021; Hou et al.,, 2023; Morra et al.,, 2023) in recent years.
+Specifically, developing statistical methods for conditional independence tests with type-I error rate control has been an important and popular research topic of large-scale genome-wide association studies (GWAS; Tang2021; Hou2023; Morra2023) in recent years.
 With the ultimate goal of identifying novel targets for the development of genomic-driven medicine, the contribution of more genetic variants to complex phenotypes has been investigated in GWAS.
@@
 On one hand, with the stringent criterion to control the family-wise error rate (FWER), conventional GWAS are usually suboptimal in terms of the statistical power of detecting important variants.
-On the other hand, because conventional GWAS base their inference on marginal association models which regress $Y$ on one $X_{j}$ at a time, they can only provide blurred results with a lot of proxy variants correlated with the true causal variants (Schaid et al.,, 2018).
+On the other hand, because conventional GWAS base their inference on marginal association models which regress $Y$ on one $X_{j}$ at a time, they can only provide blurred results with a lot of proxy variants correlated with the true causal variants (Schaid2018).
 Both issues become more severe when more variants are sequenced at a higher resolution because the $p$-value threshold of FWER control also becomes more stringent and there are more proxy variants strongly correlated with true causal variants.
-Although many fine-mapping methods have been proposed (Benner et al.,, 2017; Zou et al.,, 2022) to pinpoint the causal variants, they can only partially address the aforementioned issues without control of type-I error rate, calling for new statistical inference methods with better theoretical properties and empirical performance in multiple testing of conditional independence for real-world genetic data analysis.
-In this article, we investigate the stage I meta-analysis of the European Alzheimer & Dementia Biobank (EADB) dataset and the UK Biobank (UKBB) dataset (Bellenguez et al.,, 2022) to identify genetic variants associated with AD.
-This meta-analyzed EADB–UKBB dataset is deposited in the European Bioinformatics Institute GWAS Catalog [^1] https://www.ebi.ac.uk/gwas/ (Accession number: GCST90027158), including $39,106$ clinically diagnosed AD cases, $46,828$ proxy cases, and $401,577$ control cases from $15$ European countries.
-We focus on $640,576$ directly genotyped variants with minor allele frequency (MAF) at least $0.01$, resulting in a processed dataset $\mathbb{G}$ of $487,511$ observations and $640,576$ features.
-With such a large dataset, we aim to discover as many important genetic variants that carry independent information of AD without making too many false discoveries.
+This calls for new statistical inference methods with better theoretical properties and empirical performance in multiple testing of conditional independence for real-world genetic data analysis.
+In this article, we investigate the stage I meta-analysis of the European Alzheimer & Dementia Biobank (EADB) dataset and the UK Biobank (UKBB) dataset (Bellenguez2022) to identify genetic variants associated with AD.
+This meta-analyzed EADB–UKBB dataset is deposited in the European Bioinformatics Institute GWAS Catalog (https://www.ebi.ac.uk/gwas/; accession number: GCST90027158), including $39,106$ clinically diagnosed AD cases, $46,828$ proxy cases, and $401,577$ control cases from $15$ European countries.
+We focus on $650,706$ directly genotyped variants with minor allele frequency (MAF) at least $0.01$, resulting in a processed dataset $\mathbb{G}$ of $487,511$ observations and $650,706$ features.
+With such a large dataset, we aim to discover as many as possible important genetic variants that carry independent information of AD without making too many false discoveries.
 Mathematically, such a task can be transferred into a multiple testing problem of conditional independence between a large set of variants $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ and the binary response $Y$ which indicates whether a case is clinically diagnosed AD.
@@
 Recent decades have witnessed fruitful works in the development of multiple testing procedures.
-To control the family-wise error rate (FWER) of making at least one false discovery, a lot of $p$-value-based methods have been developed as improvements of the classic Bonferroni correction, including the Šidák correction (Šidák,, 1967), Holm’s step-down procedure (Holm,, 1979) and Hochberg’s step-up procedure (Hochberg,, 1988).
+To control the family-wise error rate (FWER) of making at least one false discovery, a lot of $p$-value based methods have been developed as improvements of the classic Bonferroni correction, including the Šidák correction (Sidak1967), Holm’s step-down procedure (Holm1979) and Hochberg’s step-up procedure (Hochberg1988).
 However, FWER has been criticized for its conservativeness in many settings, especially when most of the signals are weak.
-To perform powerful inference, Benjamini and Hochberg, (1995) proposed the false discovery rate (the expected proportion of false discoveries in all discoveries) as an alternative type-I error rate measure and developed a procedure to perform multiple testing with provable FDR control under the assumption that $p$-values are independent or positively dependent.
-Inspired by this pioneering work, Benjamini and Yekutieli, (2001) developed a generalized procedure without assumptions on dependencies among $p$-values, while Storey, (2002) and Whittemore, (2007) investigated the Bayesian counterpart of FDR.
+To perform powerful inference, Benjamini1995 proposed the false discovery rate (the expected proportion of false discoveries in all discoveries) as an alternative type-I error rate measure and developed a procedure to perform multiple testing with provable FDR control under the assumption that $p$-values are independent or positively dependent (Yekutieli2001).
+Inspired by this pioneering work, Yekutieli2001 developed a generalized procedure without assumptions on dependencies among $p$-values, while Storey2002 and Whittemore2007 investigated the Bayesian counterpart of FDR.
 However, as most of the aforementioned methods require $p$-values of nulls for all features, their feasibility on multiple testing of conditional independence becomes questionable when the number of features $p$ greatly exceeds the sample size $n$.
-The main reason is that $p$-values of conditional independence under high-dimensional scenarios are nontrivial and generally not available except in a very few specific cases (Tibshirani et al.,, 2016; Lee et al.,, 2016).
-Unlike all the above methods, the recently developed knockoff filter (Barber and Candès,, 2015; Candès et al.,, 2018) and its variants (Gimenez et al.,, 2019; Gimenez and Zou,, 2019; Barber and Candès,, 2019; Huang and Janson,, 2020; Ren and Candès,, 2023) are a family of approaches that can control finite-sample FDR in multiple testing of conditional independence hypotheses at the feature-versus-feature level
+The main reason is that $p$-values of conditional independence under high-dimensional scenarios are nontrivial and generally not available except in a very few specific cases (Tibshirani2016; Lee2016).
+Unlike all the above methods, the recently developed model-X knockoff filter (Candes2018) and its variations (Gimenez2019; Gimenez2019b; Barber2019; Bates2020; Huang2020; Ren2023; Li2024) are a family of approaches that can control finite-sample FDR in multiple testing of conditional independence hypotheses at the feature-versus-feature level
 $H^{(\text{ff})}_{j}:X_{j}\perp Y|\textbf{X}_{-j},\text{ where }\textbf{X}_{-j}=(X_{1},\ldots,X_{j-1},X_{j+1},\ldots,X_{p})^{\mathsf{T}},\quad j=1,\ldots,p,$ (1)
 without any assumption on the conditional distribution $Y|\textbf{X}$.
-Specifically, these methods first construct knockoffs $\widetilde{\textbf{X}}$ that mimic the dependency among features X and are not dependent to the response $Y$.
+Specifically, these methods first construct knockoffs $\widetilde{\textbf{X}}$ that mimic the dependency among features X and are not dependent on the response $Y$.
 By doing so, knockoffs play the role of synthetic controls: if $H^{(\text{ff})}_{j}$ is true, ${X}_{j}$ and $\widetilde{X}_{j}$ are indistinguishable even with information of $Y$; if $H^{(\text{ff})}_{j}$ is false, ${X}_{j}$ and $\widetilde{X}_{j}$ possess different dependencies with $Y$.
-By comparing dependencies between $Y$ and X with dependencies between $Y$ and $\widetilde{\textbf{X}}$, these methods have both controlled FDR and superior power in detecting true causal features as validated in extensive simulation studies.
-However, in the analysis of real-world genetic data, methods that infer $H^{(\text{ff})}_{j}$’s usually suffer severe power loss.
-The main reason is that linkage disequilibrium could introduce high correlations among genetic variants, making it difficult to distinguish variants that are truly associated with the response (whose $H^{(\text{ff})}_{j}$’s are false) from their strongly correlated null variants (whose $H^{(\text{ff})}_{j}$’s are true).
-This would significantly decrease the number of discoveries as shown by the simulation studies in Barber and Candès, (2015) and Candès et al., (2018).
+By comparing the dependencies between $Y$ and X with those between $Y$ and $\widetilde{\textbf{X}}$, these methods have both controlled FDR and superior power in detecting true causal features as validated in extensive simulation studies.
+However, in the analysis of real-world genetic data where genetic variants are the features to be selected, methods that infer $H^{(\text{ff})}_{j}$’s usually suffer severe power loss.
+The main reason is that linkage disequilibrium (LD) could introduce high correlations among genetic variants, making it difficult to distinguish variants that are truly associated with the response (whose $H^{(\text{ff})}_{j}$’s are false) from their strongly correlated null variants (whose $H^{(\text{ff})}_{j}$’s are true).
+This would significantly decrease the number of discoveries as shown by the simulation studies in Barber2015 and Candes2018.
 Such a power loss issue would deteriorate when more genetic variants with strong correlations are recorded under a growing resolution of sequencing.
-For example, when we apply the model-X knockoff filter (Candès et al.,, 2018) to the EADB-UKBB dataset (Bellenguez et al.,, 2022), only 138 AD-associated genetic variants from 49 loci (here, we define two loci as different if they are at least 1Mb away from each other) are identified under the target FDR level $0.10$ as shown in Table 1.
-Compared with the 51 loci discovered by marginal association test ($p$-value threshold: $5\times 10^{-8}$) in Bellenguez et al., (2022), although model-X knockoff filter identified 6 new loci, it misses several important AD loci with strong signals, including the “FAM193B” locus, the “FAM157C” locus, the “NSF” locus, the “SIGLEC11” locus and the “APP” locus.
-Table 1: Summary of results by applying different methods to the EADB-UKBB dataset.
+For example, when we apply the model-X knockoff filter (Candes2018) to the EADB-UKBB dataset (Bellenguez2022) , only 162 AD-associated genetic variants from 38 loci (here, we define two loci as different if they are at least 1Mb away from each other) are identified under the target FDR level $0.10$ as shown in Table 1.
+Compared with the 54 loci discovered by marginal association test ($p$-value threshold: $5\times 10^{-8}$) in Bellenguez2022, although model-X knockoff filter identified 7 new loci, it misses several important AD loci with strong signals, including the “ZCWPW1” locus on chromosome 7, the “PTK2B” locus on chromosome 8, the “ECHDC3” locus on chromosome 10, the ”MS4A4E” locus on chromosome 11, the “FERMT2” locus on chromosome 14, the “FAM157C” locus on chromosome 16 and the “SIGLEC11” locus on chromosome 19.
+This can seen in Figures 1 (a)-(b) where we displays the difference of inference results between marginal association test and model-X knockoff filter within a high LD region between positions 31571218 $\sim$ 32682663, chromosome 6.
+As correlation among genetic variants is high in this region, most of genetic variants possess a small $p$-values, which not only implies the existence of contributing variants but also obstacles our distinguishing of the contributing ones from all proxy ones.
+Such great correlations also result in no discoveries (under FDR $0.10$) of contributing variants of the model-X knockoff filter as the importance score of the most contributing variant does not exceed the data-driven threshold.
+Table 1: Summary of results by applying different existing methods to the EADB-UKBB dataset, where bold values indicate the best performance under different evaluation metrics.
 Method Number of Average number of Average Average
@@
 loci identified locus catching sets catching sets
-Marginal association test 51 18.078 18.078 0.477
-Model-X knockoff filter 49 2.816 1.000 1.000
-Group knockoff filter 84 23.714 13.105 0.610
+Marginal association test 54 17.093 17.093 0.489
+Model-X knockoff filter 38 4.263 1.000 1.000
+Group knockoff filter 91 19.923 9.157 0.651
+KnockoffZoom (2 layers) 91 15.152 5.621 0.851
+KnockoffZoom (3 layers) 91 14.681 5.409 0.858
+KeLP (2 layers) 45 10.422 2.535 0.918
+KeLP (3 layers) 38 10.605 2.488 0.932
 : Catching sets refer to
@@
 (c) variant groups identified by the group knockoff filter.
+(d) resolution-adaptive discoveries made by KnockoffZoom and KeLP that can be either variants identified
+   at the variant level or variant groups identified at the group level containing no identified variants.
 : Purity refers to the minimum absolute correlation within the catching set.
-To address this power loss issue, Dai and Barber, (2016) proposed the group knockoff filter.
-By defining groups $B_{1},\ldots,B_{k}$ such that features with strong correlations are in the same group and changing the inference target as
+: Size and purity of catching sets provided by model-X knockoff filter are trivially 1 and thus not included
+   in comparison.
+Refer to captionRefer to caption
+(a) Manhattan plot of marginal association test and the correlation plot within the high-LD region.
+(b) Manhattan plot of feature importance statistics of
+different genetic variants under model-X knockoff filter.
+(c) Manhattan plot of the contribution of different variants to the importance statistics of the group under the group knockoff filter.
+Figure 1: Comparison of inference results provided by marginal association test (Bellenguez2022), model-X knockoff filter (Candes2018) and the group knockoff filter (Dai2016) in a high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6.
+(Displaying order of genetic variants is reorganized according to the hierarchical clustering of $(1-|\text{cor}(G_{i},G_{j})|)$ for better visualization.)
+To address this power loss issue in real-world genetic data analysis, people turn their inference target from the feature-versus-feature level to the group-versus-group level and use group knockoff filter to detect signals lie in high LD region.
+First introduced by Dai2016 and developed by multiple works (Katsevich2019; Sesia2020; Sesia2021; Spector2022; Chu2024; Gablenz2024), the group knockoff filter defines groups $B_{1},\ldots,B_{k}$ such that features with strong correlations are empirically allocated in the same group and perform multiple testing on
 $H^{\text{(gg)}}_{k}:\textbf{X}_{B_{k}}\perp Y|\textbf{X}_{-B_{k}},\quad k=1,\ldots,K,$ (2)
-at the group-versus-group level, important features and their strongly correlated null features can be rejected as a group.
+at the group-versus-group level, where important features and their strongly correlated null features can be rejected as a group.
 Because the grouping is done to ensure features from different groups are weakly correlated, the group knockoff filter can circumvent the obstacle of high correlation in inferring $H^{(\text{ff})}_{j}$ with significantly higher power.
-To apply the group knockoff filter to the EADB-UKBB dataset (Bellenguez et al.,, 2022), we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ and construct $321,535$ variant groups by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$.
-By doing so, variants in the EADB-UKBB dataset (Bellenguez et al.,, 2022) are grouped such that the average absolute correlation of variants from different groups is lower than 0.5.
-Under such a group structure, the group knockoff filter identifies 152 AD-associated variant groups from 84 loci under the target FDR level $0.10$.
-Specifically, among these 152 identified groups, 126 are located outside the APOE/APOC region with the strongest association to AD (chromosome 19, positions 44909011 $\sim$ 45912650), including groups in loci “FAM193B”, “FAM157C”, “NSF”, “SIGLEC11” and “APP” that are missed by the model-X knockoff filter.
-In addition, the group knockoff filter also identifies 39 new loci that are not discovered by Bellenguez et al., (2022).
+To apply the group knockoff filter to the EADB-UKBB dataset (Bellenguez2022), we generate the second-order group knockoffs as follows by assuming that the distribution of X can be well approximated by a multivariate Gaussian distribution.
+To empirically construct the genetic variant groups according to the correlation, we first use the UK Biobank directly genotyped data (which also records genetic information of Europeans) as the reference panel and compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ via the Pan-UKB consortium (https://pan.ukbb.broadinstitute.org).
+We then perform the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$ and construct $321,569$ variant groups.
+By doing so, the average absolute correlation of variants (features) from different groups is lower than 0.5.
+With second-order group knockoffs generated under such a group structure and the maximum-entropy construction using Algorithm 2 of Chu2024, the group knockoff filter identifies 198 AD-associated variant groups from 91 loci under the target FDR level $0.10$.
+Specifically, among these 198 identified groups, 121 are located outside the strongest APOE/APOC region (chromosome 19, positions 44909011 $\sim$ 45912650), including groups in loci “ZCWPW1”, “PTK2B”, “ECHDC3”, ”MS4A4E”, “FERMT2”, “FAM157C” and“SIGLEC11” that are missed by the model-X knockoff filter.
+In addition, the group knockoff filter also identifies 43 new loci that are not discovered by Bellenguez2022.
 However, doing so can only provide a series of AD-associated variant groups without information on which variants contribute more to AD variation within each group, leaving the target of identifying AD-associated variants not fully addressed.
-This can be observed from Table 1.
-With an average size $13.105$ of the identified groups given by the group knockoff filter, 35 identified groups contain at least 10 genetic variants and the 5 largest groups contain 314, 232, 136, 121 and 117 variants respectively.
-With such large sizes, which genetic variants in these groups contribute more to AD variation remains unanswered.
+This can be observed from Table 1, where each identified groups given by the group knockoff filter are treated as a catching set of genetic signals.
+With an average size $19.923$ of catching sets, 30 catching sets contain at least 10 genetic variants and the 5 largest sets contain 314, 232, 117, 84 and 57 variants respectively.
+With such large sizes, which genetic variants in these catching sets contribute more to AD variation remains unanswered.
+This can also be seen in Figure 1 (c).
+Within the high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6, all genetic variants are clustered in the same group and such a group is identified as a sole catching set.
+However, such an inference result loses the information that all of the importance of this variant group is contributed by the variant rs9469112 as shown in Figure 1 (c).
+To improve the precision of catching sets of genetic signals, several approaches have been developed to return catching sets of multiple resolutions or in a resolution-adaptive manner.
+One of the pioneering method is the multilayer knockoff filter (Katsevich2019) which simultaneously infers $H^{(\text{ff})}_{j}$’s and $H^{(\text{gg})}_{k}$’s, while the later developed KnockoffZoom [^1] In this article, we adopt Algorithm 5 in the supplementary note 4 of Sesia2020.
+(Sesia2020) and knockoff e-value linear program (KeLP; Gablenz2024) are implemented to identify genetic signals in resolution-adaptive manners by using catching sets at different resolutions.
+Specifically, KnockoffZoom gradually refines catching sets by sequentially performing group knockoff inference from lower resolutions to higher resolutions and finally using model-X knockoff filter at the feature-versus-feature level.
+KeLP, in contrast, utilizes the method of Ren2024 to compute $e$-values of all variant groups at different resolutions and all variants and identifies signals at the finest possible resolution by solving an optimization problem with nesting constraints.
+However, when we apply KnockoffZoom and KeLP to identify contributing genetic variants under the target FDR level $0.10$ at both the feature-versus-feature level as in the model-X knockoff filter and group-versus-group level as in the group knockoff filter (which we label as “(2 layers)”), existing issues are not fully addressed.
+On one hand, as the inference of KnockoffZoom (2 layers) at the group-versus-group level is the same as the group knockoff filter, KnockoffZoom (2 layers) identifies the same loci as the group knockoff filter and does not suffer from any power loss.
+However, as signal-to-noise ratios of $H^{(\text{ff})}_{j}$’s are low due to high LD, half of groups (99/198) contain no identified variants at the feature-versus-feature level.
+Thus, KnockoffZoom (2 layers) only records moderate improvement of average size (decreases from 9.157 to 5.621) and purity (increases from 0.651 to 0.851) of catching sets.
+This can be seen from Figure 7 (a) in Appendix A, where importance score of the most contributing variant in the high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6, remains lower than the data-driven threshold.
+Even when we add an intermediate layer (obtained by the hierarchical clustering with average linkage and cutoff value $0.25$) between the current two levels in applying KnockoffZoom (we refer to as “KnockoffZoom (3 layers)”), such a phenomenon retains as shown in Figures 7 (b)-(c).
+KeLP, on the other hand, suffer from greater power loss even compared to the model-X knockoff filter, no matter with 2 layers or 3 layers.
+The main reason is that computing e-values for resolution-adaptive inference under the target FDR level $0.10$ requires inferences under a target FDR level lower than $0.10$ throughout all resolutions.
+As only variants and groups with nonzero e-values can be selected by KeLP, lower target FDR levels would decrease the number of variants and groups with nonzero e-values and thus the number of catching sets.
+Another category of commonly used approaches are various Bayesian fine-mapping approaches (Wilson2010; Guan2011; Hormozdiari2014; Chen2015; Benner2016; Wang2020; Zou2022; Li2024a).
+For comprehensive review, please refer to Schaid2018 and Li2025.
+Generally, existing Bayesian fine-mapping approaches fit a parametric model between the response and features (genetic variants) in identified important regions, compute the posterior inclusion probability (PIP) of each feature to be contributive via Bayesian approaches and construct credible sets of features to contain at least one signal with large posterior probability.
+Potentially, one can perform fine-mapping within variant groups identified by the group knockoff filter and obtain credible sets of smaller size and higher precision.
+However, using such a two-step heuristic approach still faces the tradeoff between power and precision via the selection of coverage probability threshold.
 1.3 Our Contribution
-In this article, to address the limitations of existing knockoff filters in discovering important variants of AD in the analysis of the EADB-UKBB dataset, we propose a new knockoff filter that is more powerful in inferring conditional independence hypotheses and identifying important features from a set of strongly correlated features.
-Our contribution in this article is two-folded.
-  • Presenting a family of conditional independence hypotheses that allow identifying important features using group knockoffs.
-Compared to existing conditional independence hypotheses, the new hypotheses allow the use of group knockoffs construction mechanism (Dai and Barber,, 2016) to provide informative discoveries at the level of features with higher power.
-  • Developing the new knockoff filter with FDR control, high empirical power and informativeness [^2] In this article, informativeness refers to the precision of catching sets in identifying conditionally independent signals and is quantified by the size and purity of catching sets..
-Specifically, we thoroughly investigate the theoretical properties of the developed knockoff filter and empirically evaluate their performance via simulated experiments.
+In this article, to address the limitations of existing knockoff methods in discovering important variants of AD in the analysis of the EADB-UKBB dataset, we propose a new knockoff filter that achieves balance between the statistical power and precision in identifying important features from a set of strongly correlated features.
+Built upon group knockoffs, our filter simultaneously selects groups containing important features and performs fine-mapping to identify the most promising features within each selected group.
+Thus, our method can return refined catching sets that are small in size and high in purity without large power loss.
+Specifically, our approach of refinement is motivated by penalized regressions that can efficiently learn the sparse nature of causality via the sparse feature importance, which is leveraged in searching of catching sets under alternative variant-level hypotheses.
 The rest of this article is organized as follows.
-In Section 2, we develop our knockoff filter and perform theoretical analysis.
-Section 3 illustrates the empirical performance of our method and its advantage over existing alternatives in power and informativeness via simulated experiments.
-Finally, we apply our method to the AD analysis of the EADB-UKBB dataset described in Section 4 and demonstrate that the method exhibits high power and precision.
+In Section 2, we develop our new knockoff filter and perform theoretical analysis.
+Section 3 illustrates the empirical performance of our method and its advantage over existing alternatives in power and precision via simulated experiments.
+Finally, we apply our method to the AD analysis of the EADB-UKBB dataset in Section 4 and demonstrate that the method exhibits balance between power and precision.
 Section 5 concludes with discussions.
@@
 2.1 Problem Statement
-Consider independent and identically distributed (i.i.d.) samples $\{(\textbf{x}_{i},y_{i})|i=1,\ldots,n\}$ from a joint distribution $f(\mathbf{X},Y)$, where features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ are partitioned into $K$ disjoint groups $B_{1},\ldots,B_{K}$.
+Consider independent and identically distributed (i.i.d.) samples $\{(\textbf{x}_{i},y_{i})|i=1,\ldots,n\}$ from a joint distribution $f(\mathbf{X},Y)$, where features $\textbf{X}=(X_{1},\ldots,X_{p})^{\mathsf{T}}$ are partitioned into $K$ disjoint groups $B_{1},\ldots,B_{K}$ for inference.
+For example, in analysis of genetic data, features (variants) with high correlation are usually allocated to the same group such that features from different groups are weakly or moderately correlated.
 Let $\mathbb{X}=[\textbf{x}_{1},\ldots,\textbf{x}_{n}]^{\mathsf{T}}$ denote the data matrix and $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ be the vector of responses.
@@
 Here, we refer $H_{j}^{(\text{fg})}$ as the conditional independence hypothesis at the feature-versus-group level as it depicts the independence between a feature and the response conditional on other feature groups which omits information of other features in the same group.
-Under the belief that the response $Y$ only depends on a relatively small number of features, we would like to find as many important features without making too many false discoveries with respect to $H_{j}^{(\text{fg})}$’s.
+Under the belief that the response $Y$ only depends on a relatively small number of features as validated in many GWAS studies, we would like to find as many important features as possible without making too many false discoveries with respect to $H_{j}^{(\text{fg})}$’s.
 In other words, our target is to obtain a rejection set $\mathcal{R}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is rejected}\}$ such that the false discovery rate (FDR, the expectation of false discovery proportion),
-$\text{FDR}^{(\text{fg})}=\mathbf{E}\{\text{FDP}^{(\text{fg})}\},\quad\text{where }\text{FDP}^{(\text{fg})}=\frac{\#(\mathcal{R}^{(\text{fg})}\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}},\quad\mathcal{H}_{0}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is true}\},$
-is controlled under the target level $\alpha$ ($10<\alpha<$ 1).
-Under the same setup of Candès et al., (2018) that the distribution of features X is completely known while no information is provided regarding the conditional distribution $Y|\textbf{X}$, our method differs in the hypotheses to test (3) from existing approaches, which test either conditional independence hypotheses (1) at the feature-versus-feature level or conditional independence hypotheses (2) at the group-versus-group level.
-Such a difference is the origin of our method’s advantages over the existing knockoff filters.
-On one hand, unlike the group knockoff filter (Dai and Barber,, 2016), the conclusion provided by our method is at the feature level and is thus more informative on which features contribute more to the variation of $Y$.
-On the other hand, compared to those that infer $H^{(\text{ff})}_{j}$’s, our method can also utilize group knockoffs that are more liberal in construction to gain power in selecting important features.
-However, such advantages do not come free as new challenge is posted that no existing filters can provide control on $\text{FDR}^{(\text{fg})}$, which is elaborated and addressed in the following sections.
+$\text{FDR}^{(\text{fg})}=\mathbf{E}\{\text{FDP}^{(\text{fg})}\},\quad\text{where }\text{FDP}^{(\text{fg})}=\frac{\#(\mathcal{R}^{(\text{fg})}\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}},\quad\mathcal{H}_{0}^{(\text{fg})}=\{j|H_{j}^{(\text{fg})}\text{ is true}\},$ (4)
+is controlled under the target level $\alpha$ ($0<\alpha<1$).
+Our method is under the same setup of Candes2018 that the distribution of features X is known completely or well approximated up to the first two moments, while no information is provided regarding the conditional distribution $Y|\textbf{X}$.
+That is to say, our framework does not assume that the contribution of $X_{j}$’s to $Y$ are additive (independent) and the FDR control remains valid even the underlying conditional distribution $Y|\textbf{X}$ is deviated from the linear model.
+Our method differs from existing approaches, which test either (1) at the feature-versus-feature level or (2) at the group-versus-group level, as follows.
+  • Compared to those that infer $H^{(\text{ff})}_{j}$’s, our method can utilize group knockoffs that are more liberal in construction to achieve higher power in identifying important features.
+  • It is true that in most of GWAS studies where highly correlated features are allocated in the same groups, our hypotheses $H_{j}^{(\text{fg})}$’s are equivalent to $H_{k}^{(\text{gg})}$’s at the group-versus-group level for $j\in B_{k}$.
+That is to say, including feature $X_{j}$ in $\mathcal{R}^{(\text{fg})}$ implies that the group $B_{k}$ contains important features.
+The difference is that our method simultaneously selects groups containing important features and performs fine-mapping to identify the most promising features within each selected group.
+Thus, instead of returning variant groups with many variants as large and impure catching sets, our method can further refine the selected variant groups.
+In addition, our method can achieve better balance between power and precision compared to the two-step heuristic alternative – applying post-selection fine-mapping to groups selected by the group knockoff filter.
+However, such a gain does not come free as new challenge is posted that no existing filters can provide control on $\text{FDR}^{(\text{fg})}$, which is elaborated and addressed in the following sections.
 2.2 Feature Importance Scores Under Group Knockoffs
@@
   • $\diamond$
-    (Knockoffs Construction): Construct the knockoff copy $\widetilde{\textbf{X}}$ of X such that $\perp Y|\textbf{X}$ X and the joint distribution of $(\textbf{X},\widetilde{\textbf{X}})$ is exchangeable at the group level (Dai and Barber,, 2016),
+[…]
     $(\textbf{X},\widetilde{\textbf{X}})_{\text{swap}(\mathcal{S})}{\displaystyle\mathop{=\joinrel=}^{D}}(\textbf{X},\widetilde{\textbf{X}}),\quad\forall\mathcal{S}\subset\{1,\ldots,K\},$ (5)
-    where $(\textbf{X},\widetilde{\textbf{X}})_{\text{swap}(\mathcal{S})}$ is obtained by swapping $\textbf{X}_{B_{k}}$ and $\widetilde{\textbf{X}}_{B_{k}}$ in $(\textbf{X},\widetilde{\textbf{X}})$ for all $k\in\mathcal{S}$.
+[…]
   • $\diamond$
@@
     (Feature Filtering): Reject $H^{(\text{fg})}_{j}$ if $W_{j}$ is positive and large enough (i.e., $T_{j}$ dominates $\widetilde{T}_{j}$).
-However, directly inputting $W_{1},\ldots,W_{p}$ to the model-X knockoff filter at feature level (Candès et al.,, 2018) cannot provide control on $\text{FDR}^{(\text{fg})}$ of the rejection set.
-The main reason is that under the group knockoffs model $f(\textbf{X},\widetilde{\textbf{X}})$, $W_{1},\ldots,W_{p}$ do not possess the coin-flipping property (i.e., conditional on $|W_{j}|$’s and $\text{sign}(W_{j})$’s for those false $H^{(\text{fg})}_{j}$’s, $\text{sign}(W_{j})$’s for those true $H^{(\text{fg})}_{j}$’s are independently and uniformly distributed on {±$\{\pm 1\}$), the key property that guarantees FDR control in Candès et al., (2018).
-Specifically, for most of the widely-used feature importance scores (e.g., the marginal correlations, the Lasso coefficient-difference statistics and the Lasso signed max statistics; Barber and Candès,, 2015, Candès et al.,, 2018), $W_{j}$’s for those true $H^{(\text{fg})}_{j}$’s are dependent coin flips.
-This comes from (4) where exchangeability only exists among groups, making signs of $W_{j}$’s from the same group not independent.
-Thus, we relax the unachievable coin-flipping property to the between-group coin-flipping property which can be satisfied by feature importance scores $W_{j}$’s in the form of
-$\begin{cases}T_{j}=t([\mathbb{X}_{j},\mathbb{X}_{B_{k}\setminus\{j\}},\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{k}\setminus\{j\}},\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}],\textbf{y}),\\ \widetilde{T}_{j}=t([\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{k}\setminus\{j\}},{\mathbb{X}}_{j},{\mathbb{X}}_{B_{k}\setminus\{j\}},\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}],\textbf{y}),\end{cases}\quad k=1,\ldots,K\text{, }j\in B_{k},$ (5)
-where $\text{where }\{\mathbb{X}_{-B_{k}},\widetilde{\mathbb{X}}_{-B_{k}}\}\text{ denotes the set of unordered pairs }\{\mathbb{X}_{j^{\prime}},\widetilde{\mathbb{X}}_{j^{\prime}}\}\text{ for all }j^{\prime}\notin B_{k},$
-as shown in Theorem 1.
+[…]
+Thus, we relax the unachievable coin-flipping property to the between-group coin-flipping property and propose feature statistics $W_{j}$’s using anti-symmetric functions $w_{j}$’s and
+[…]
+[…]
+Here, the swapping between $T_{j}$ and $\widetilde{T}_{j}$ requires the swapping between $\mathbb{X}_{B_{k}}$ and $\widetilde{\mathbb{X}}_{B_{k}}$, while it only requires the swapping between $\mathbb{X}_{j}$ and $\widetilde{\mathbb{X}}_{j}$ in the model-X knockoff filter.
+However, most of commonly used feature importance scores in the model-X knockoff filter also satisfy (6).
+Examples include:
+  • $\diamond$
+    marginal correlation with response: absolute values of the empirical marginal correlations $T_{j}=\Big|\widehat{\rho}(X_{j},Y)\Big|$ and $\widetilde{T}_{j}=\Big|\widehat{\rho}(\widetilde{X}_{j},Y)\Big|$;
+  • $\diamond$
+    joint lasso: absolute values of lasso estimators $T_{j}=\Big|\widehat{\beta}_{j}\Big|$ and $\widetilde{T}_{j}=\Big|\widehat{\widetilde{\beta}}_{j}\Big|$ of the linear model (Tibshirani1996)
+    $Y=\sum_{j=1}^{p}\left({\beta}_{j}{{X}}_{j}+\widetilde{\beta}_{j}\widetilde{{X}}_{j}\right)+e;$ (7)
+  • $\diamond$
+[…]
+    $Y=\sum_{j^{\dagger}\notin B_{k}}\left({\beta}_{j^{\dagger}}{{X}}_{j^{\dagger}}+\widetilde{\beta}_{j^{\dagger}}\widetilde{{X}}_{j^{\dagger}}\right)+e,\quad\text{where }j\in B_{k};$
+  • $\diamond$
+    separate lasso: absolute values of lasso estimators $T_{j}=\Big|\widehat{\beta}_{j}\Big|$ and $\widetilde{T}_{j}=\Big|\widehat{\widetilde{\beta}}_{j}\Big|$ of the separate linear model
+    $Y={\beta}_{j}{{X}}_{j}+\widetilde{\beta}_{j}\widetilde{{X}}_{j}+\sum_{j^{\dagger}\notin B_{k}}\left({\beta}_{j^{\dagger}}{{X}}_{j^{\dagger}}+\widetilde{\beta}_{j^{\dagger}}\widetilde{{X}}_{j^{\dagger}}\right)+e,\quad\text{where }j\in B_{k}.$
+In our practice, we suggest using joint lasso importance score because it is computationally efficient and induces sparsity for better refinement of catching sets as stated in Section 2.1.
+As shown in Theorem 1, feature importance scores in the form of (6) satisfy the between-group coin-flipping property.
 Theorem 1.
@@
     1. $\diamond$
-      (Uniformity) $\text{\rm sign}(W_{j})$ is uniformly distributed on {±$\{\pm 1\}$ for all $j\in\mathcal{H}_{0}^{(\text{fg})}$;
+      (Uniformity) $\text{\rm sign}(W_{j})$ is uniformly distributed on $\{\pm 1\}$ for all $j\in\mathcal{H}_{0}^{(\text{fg})}$;
     2. $\diamond$
@@
 if $H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ for all $j\in B_{k}$ and $k=1,\ldots,K$.
-The proof of Theorem 1 is provided in Section A of the Appendix.
-Specifically, Theorem 1 relies on a condition “$H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ for all $j\in B_{k}$ and $k=1,\ldots,K$”.
+The proof of Theorem 1 is provided in Appendix B.
+Specifically, Theorem 1 relies on a condition $H_{j}^{(\text{fg})}$ implies $H_{k}^{\text{(gg)}}$ for all $j\in B_{k}$ and $k=1,\ldots,K$”.
+That is to say, if $H_{k}^{\text{(gg)}}$ is false and group $B_{k}$ is important, all $H_{j}^{(\text{fg})}$’s ($j\in B_{k}$) are false and any $j\in B_{k}$ included in the selection sets would not be counted as false discovery.
 Such a condition is generally valid in genetic data analysis, analogous to the usual expectation in fine-mapping that the association between an important variant $X_{j}$ and the response $Y$ is in the same direction as the direct effect of $X_{j}$ on $Y$.
-Examples of feature importance scores in the form of (5) include:
-  • $\diamond$
-    marginal correlation with response: absolute values of the empirical marginal correlations $T_{j}=\Big{|}\widehat{\rho}(X_{j},Y)\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\rho}(\widetilde{X}_{j},Y)\Big{|}$;
-  • $\diamond$
-    joint lasso: absolute values of lasso estimators $T_{j}=\Big{|}\widehat{\beta}_{j}\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\widetilde{\beta}}_{j}\Big{|}$ of the linear model
-    $Y=\sum_{j=1}^{p}\left({\beta}_{j}{{X}}_{j}+\widetilde{\beta}_{j}\widetilde{{X}}_{j}\right)+e;$ (6)
-  • $\diamond$
-    marginal correlation with lasso residual: absolute values of estimated marginal correlations $T_{j}=\Big{|}\widehat{\rho}(X_{j},\hat{e})\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\rho}(\widetilde{X}_{j},\hat{e})\Big{|}$ between $(X_{j},\widetilde{X}_{j})^{\mathsf{T}}$ and the lasso residual of the linear model
-    $Y=\sum_{j^{\dagger}\notin B_{k}}\left({\beta}_{j^{\dagger}}{{X}}_{j^{\dagger}}+\widetilde{\beta}_{j^{\dagger}}\widetilde{{X}}_{j^{\dagger}}\right)+e,\quad\text{where }j\in B_{k};$
-  • $\diamond$
-    separate lasso: absolute values of lasso estimators $T_{j}=\Big{|}\widehat{\beta}_{j}\Big{|}$ and $\widetilde{T}_{j}=\Big{|}\widehat{\widetilde{\beta}}_{j}\Big{|}$ of the separate linear model
-    $Y={\beta}_{j}{{X}}_{j}+\widetilde{\beta}_{j}\widetilde{{X}}_{j}+\sum_{j^{\dagger}\notin B_{k}}\left({\beta}_{j^{\dagger}}{{X}}_{j^{\dagger}}+\widetilde{\beta}_{j^{\dagger}}\widetilde{{X}}_{j^{\dagger}}\right)+e,\quad\text{where }j\in B_{k}.$
-Although the above feature importance scores can be used to compute feature statistics $W_{1},\ldots,W_{j}$ with the between-group coin-flipping property, they suffer several problems in practice.
-On one hand, with not much information of $H_{j}^{(\text{fg})}$’s, marginal correlation with response usually lacks power.
-So do the joint lasso feature importance scores, which do not directly depict evidence against $H_{j}^{(\text{fg})}$’s.
-On the other hand, with the need to perform $p$ and $K$ lasso linear model fittings, separate lasso and marginal correlation with lasso residual are both computationally intensive.
-Faced with such a tradeoff between computational efficiency and power, we propose the combined feature importance scores as computationally efficient proxies of the separate lasso importance scores
-$\begin{cases}T_{j}=|g_{j}|\cdot|\widehat{\rho}(X_{j},Y)|,\\ \widetilde{T}_{j}=|g_{j}|\cdot|\widehat{\rho}(\widetilde{{X}}_{j},Y)|,\\ \end{cases}\quad k=1,\ldots,K\text{, }j\in B_{k},$ (7)
-where $g_{j}$ is computed as follows.
-  1. 1.
-    Fit the joint linear model (6) and compute $\widehat{\gamma}_{j}=(\widehat{\beta}_{j}+\widehat{\widetilde{\beta}}_{j})/2$ and $\widehat{\delta}_{j}=(\widehat{\beta}_{j}-\widehat{\widetilde{\beta}}_{j})/2$ as the approximate estimator of the reparametrized model
-    $Y=\sum_{j=1}^{p}{\gamma}_{j}\left({{X}}_{j}+\widetilde{{X}}_{j}\right)+\sum_{j=1}^{p}{\delta}_{j}\left({{X}}_{j}-\widetilde{{X}}_{j}\right)+e.$
-  2. 2.
-    For $k=1,\ldots,K$ and $j\in B_{k}$, compute
-    $g_{j}$ $=\widehat{\gamma}_{j}+\sum_{j^{\dagger}\in B_{k}\setminus\{j\}}\frac{\psi_{j,j^{\dagger}}-\boldsymbol{\Psi}_{j,-B_{k}}\boldsymbol{\Psi}^{-1}_{-B_{k},-B_{k}}\boldsymbol{\Psi}_{-B_{k},j^{\dagger}}}{\psi_{j,j}-\boldsymbol{\Psi}_{j,-B_{k}}\boldsymbol{\Psi}^{-1}_{-B_{k},-B_{k}}\boldsymbol{\Psi}_{-B_{k},j}}\widehat{\gamma}_{j^{\dagger}},$
-    where $\boldsymbol{\Psi}=(\psi_{j,j^{\dagger}})_{p\times p}=\boldsymbol{\Sigma}-\textbf{D}/2,\quad\boldsymbol{\Sigma}=\text{Var}(\textbf{X}),\quad\textbf{D}=\boldsymbol{\Sigma}-\text{Cov}(\textbf{X},\widetilde{\textbf{X}}),$
-    as the approximate estimator of the model
-    $Y={g}_{j}\left({{X}}_{j}+\widetilde{{X}}_{j}\right)+\sum_{j^{\prime}\notin B_{k}}{\gamma}_{j^{\prime}}\left({{X}}_{j^{\prime}}+\widetilde{{X}}_{j^{\prime}}\right)+{h}_{j}\left({{X}}_{j}-\widetilde{{X}}_{j}\right)+\sum_{j^{\prime}\notin B_{k}}{\delta}_{j^{\prime}}\left({{X}}_{j^{\prime}}-\widetilde{{X}}_{j^{\prime}}\right)+e.$
-    In other words, ${g}_{j}$ measures the overall contribution of ${X}_{j}$ and $\widetilde{{X}}_{j}$ to the variation of $Y$ beyond features from other groups.
-By doing so, feature statistics $W_{j}=w_{j}(T_{j},\widetilde{T}_{j})$ ($j=1,\ldots,p$) with antisymmetric functions $w_{j}$’s satisfy the between-group coin-flipping property because
-$\text{sign}(W_{j})=\text{sign}(w_{j}(|\widehat{\rho}({{X}}_{j},Y)|,|\widehat{\rho}(\widetilde{{X}}_{j},Y)|)),$
-whose between-group coin-flipping property is validated in Theorem 1.
-In addition, as $g_{j}$ is derived to measure the total dependency between $(X_{j},\widetilde{{X}}_{j})$ and the response $Y$ conditional on $(\textbf{X}_{-B_{k}},\widetilde{\textbf{X}}_{-B_{k}})$, $|W_{j}|$’s computed from the combined feature importance scores provide proxies of $|W_{j}|$’s computed from the separate lasso feature importance scores.
-2.3 Naive Feature Filters under Group Knockoffs
-By Theorem 1, feature statistics $W_{j}$’s only satisfy the between-group coin-flipping property and thus directly applying the model-X knockoff filter (Candès et al.,, 2018) would violate the FDR control.
+2.3 Naive Feature Filter under Group Knockoffs
+By Theorem 1, feature statistics $W_{j}$’s only satisfy the between-group coin-flipping property and thus directly applying the model-X knockoff filter (Candes2018) would violate the FDR control.
 However, we can still utilize the between-group independence among $W_{j}$’s for inference with FDR control as follows.
-Recognizing that independence remains true among $W_{j}$’s for those true $H^{(\text{fg})}_{j}$’s from different groups, we consider aligning $W_{j}$’s in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$.
+Recognizing that independence remains true among $\text{sign}(W_{j})$’s for those true $H^{(\text{fg})}_{j}$’s from different groups, we consider aligning $W_{j}$’s in Table 2 such that $W_{j}$’s for all features in the group $B_{k}$ are in the $k$-th column and $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$.
 Let $\mathcal{C}_{l}$ be the set of features whose feature statistics are aligned in the $l$-th row and consider the rejection set in the form of $\mathcal{R}^{(\text{fg})}(t)=\{j|W_{j}\geq t\}$.
@@
 Here, $W_{(kl)}$ corresponds to the feature from the $k$-th group that is aligned in the $l$-th row.
-Row $B_{1}$ $B_{2}$ $B_{3}$cdots
-1 $W_{(11)}$ $W_{(21)}$ $W_{(31)}$cdots
-[…]
-[…]
-[…]
-[…]
+Row $B_{1}$ $B_{2}$ $B_{3}$ $\cdots$
+1 $W_{(11)}$ $W_{(21)}$ $W_{(31)}$ $\cdots$
+2 $W_{(12)}$ $W_{(22)}$ $W_{(32)}$ $\cdots$
+3 $W_{(13)}$ $W_{(23)}$ $W_{(33)}$ $\cdots$
+4 $W_{(14)}$ $W_{(24)}$ $W_{(34)}$ $\cdots$
+$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\ddots$
   • On one hand, we find that within each row, there exists at most one $W_{j}$ from each group.
-As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within the same row satisfy the coin-flipping property introduced in Candès et al., (2018).
+As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within the same row satisfy the coin-flipping property introduced in Candes2018.
 Thus, we can estimate the false discovery proportion (FDP) of the rejection subset $\mathcal{R}^{(l)}(t)=\{j\in\mathcal{C}_{l}|W_{j}\geq t\}$ as
     $\widehat{\text{FDP}}^{(l)}(t)=\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t\}}{1\vee\#\{j\in\mathcal{C}_{l}|W_{j}\geq t\}},$ (8)
-    for any $t>0$ and $l=1,2\ldots$ in the analogous way of Candès et al., (2018).
+    for any $t>0$ and $l=1,2\ldots$ in the analogous way of Candes2018.
   • On the other hand, we have $\text{FDP}^{(\text{fg})}(t)$ of $\mathcal{R}^{(\text{fg})}(t)$ is a weighted mean of $\text{FDP}^{(l)}(t)$’s of subsets $\mathcal{R}^{(l)}(t)=\mathcal{R}^{(\text{fg})}(t)\cap\mathcal{C}_{l}$ ($l=1,2,\ldots$) that
@@
     $=$ $\sum_{l=1}^{\phi(t)}\frac{1\vee\#\mathcal{R}^{(l)}(t)}{1\vee\#\mathcal{R}^{(\text{fg})}(t)}\times\text{FDP}^{(l)}(t),$
-    where the second equality comes from the fact that rejection subsets $\mathcal{R}^{(l)}(t)$ can be nonempty only for the first $\phi(t)=\max_{k}\#\{j\in B_{k}||W_{j}|\geq t\}$ rows because $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$ with each group $B_{k}$.
+    where the second equality comes from the fact that rejection subsets $\mathcal{R}^{(l)}(t)$ can be nonempty only for the first $\phi(t)=\max_{k}\#\{j\in B_{k}||W_{j}|\geq t\}$ rows because $|W_{(k1)}|\geq|W_{(k2)}|\geq\cdots$ in each group $B_{k}$.
 As a result, using $\widehat{\text{FDP}}^{(l)}(t)$’s in (8), we can estimate $\text{FDP}^{(\text{fg})}(t)$ of $\mathcal{R}^{(\text{fg})}(t)$ by
@@
 3:Output: The rejection set $\mathcal{R}^{(\text{fg})}=\{j|W_{j}\geq t_{\alpha}\}$.
-2.4 Rigorous Feature Filters under Group Knockoffs
-Although the naive filter seems intuitive in FDR control, such a control can not be proven.
+2.4 FVG Filter
+Although the naive filter seems intuitive in FDR control, such a control can not be strictly proven.
 To fill the gap, we provide an alternative interpretation of Algorithm 1 and consider a variation that provides provable FDR control as follows.
@@
 Based on the construction of feature statistics that $W_{j}$’s for those false $H_{j}^{(\text{fg})}$’s tend to be positive and large, Algorithm 1 solves the following optimization problem
-[…]
+$\max|\mathcal{R}^{(\text{fg})}|,\quad\text{s.t.
+}$ $\widehat{\text{FDP}}^{(\text{fg})}=\sum_{l}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\leq\alpha,$
 $\mathcal{R}^{(\text{fg})}=\cup_{l}\mathcal{R}^{(l)}=\cup_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\},\quad t^{(1)}=t^{(2)}=\cdots.$
-Based on such an observation, we utilize theoretical results of Katsevich and Sabatti, (2019) that for any $t>0$,
+Based on such an observation, we utilize theoretical results of Katsevich2019 that for any $t>0$,
 $\textbf{E}\left[\sup_{t>0}\frac{\#\{j\in\mathcal{H}_{0}^{(\text{fg})}\cap\mathcal{C}_{l}|W_{j}\geq t\}}{1+\#\{j\in\mathcal{H}_{0}^{(\text{fg})}\cap\mathcal{C}_{l}|W_{j}\leq-t\}}\right]\leq 1.93,$
 to modify the above optimization problem as
-[…]
+$\max|\mathcal{R}^{(\text{fg})}|,\quad\text{s.t.
+}$ $\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\leq\frac{v_{l}\alpha}{1.93},\quad l=1,2,\ldots,$ (10)
 $\mathcal{R}^{(\text{fg})}=\cup_{l}\mathcal{R}^{(l)}=\cup_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\},$
@@
 The rejection set $\mathcal{R}^{(\text{fg})}$ obtained by solving (10) controls $\text{FDR}^{(\text{fg})}$ at the target level $\alpha>0$.
-The proof of Theorem 2 is provided in Section B of the Appendix.
-Similar to the discussion in Sesia et al., (2020), the correction factor $1.93$ is required in the proof for technical reasons, while empirical evidence suggests that it may be practically unnecessary.
-To solve the optimization problem (10), we develop the feature-versus-group (FVG) filter (Algorithm 2) in an analogous manner to the GGM knockoff filter (Li and Maathuis,, 2021).
-We also provide other variations for FDR control with details in Section C of the Appendix.
+The proof of Theorem 2 is provided in Appendix C. Observing that the optimization problem (10) is analogous to the ones in (19) and (21) of Li2021 that
+  • $\star$
+    feature statistics $W_{j}$’s within the each row of Table 2 satisfy the coin-flipping property while $W_{j}$’s for $j=1,\ldots,p$ do not (which is analogous to the $p\times p$ feature statistics matrix in Section 3.1 of Li2021) and;
+  • $\star$
+    the contribution of each row of Table 2 to $\widehat{\text{FDP}}^{(\text{fg})}$, measured by
+    $\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\times\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]},$
+    is required to be constrained under the row-specific level $\frac{v_{l}\alpha}{1.93}$ with $\sum_{l}\frac{v_{l}\alpha}{1.93}=\frac{\alpha}{1.93}$,
+we develop the feature-versus-group (FVG) filter (Algorithm 2), whose steps 4-10 adopt the threshold calculation strategy in Algorithm 2 of Li2021 to compute the optimal thresholds $t^{(l)}$’s.
+Similar to the discussion in supplementary note 4 of Sesia2020, the correction factor $1.93$ is required in the proof for technical reasons, while empirical evidence suggests that it may be practically unnecessary.
+Thus, for all experiments in Sections 34, we implement the FVG filter without the correction factor $1.93$.
+For empirical reference, we compare the naive filter (Algorithm 1), the FVG filter with and without the $1.93$ factor via experiments in Appendix D.
 Algorithm 2 Feature-versus-group (FVG) filter with group knockoffs.
@@
 11:Output: The rejection set $\mathcal{R}^{(\text{fg})}=\cup_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}$.
+We also provide other variations with details in Appendix E.
 Remark 1.
-In the FVG filter (Algorithm 2), the budgets $v_{l}$’s can be computed adaptively to the prior belief of the number of important features within each groups.
+In the FVG filter (Algorithm 2), the budgets $v_{l}$’s can be computed adaptively to the prior belief of the number of important features within each group.
 Here, we provide two strategies of specifying $v_{l}$’s.
@@
     such that more budgets are allocated to the first several rows of Table 2 and more important features in important groups with small effect sizes are identified.
-3 Simulated Experiments
-To evaluate the proposed FVG filter in both FDR control and identifying important features, we conduct extensive experiments under various scenarios.
-Specifically, throughout simulated experiments, we generate group knockoffs under the maximum entropy (ME) construction of Chu et al., (2023), compute feature statistics as $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ with the joint lasso importance scores.
-3.1 Simulated Experiments with Synthetic Data
-To examine how the FVG filter performs under different sample sizes, we simulate 1000 datasets with a fixed number of features $p=250$ for sample sizes $n=500,1000$ and $2000$.
-Specifically, we partition these $250$ features into $K=50$ groups of equal size $B_{1}=\{X_{1},\ldots,X_{5}\},\ldots,B_{50}=\{X_{246},\ldots,X_{250}\}$.
-Based on such a partition, we generate for each simulated dataset $n$ feature vectors $\textbf{x}_{1},\ldots,\textbf{x}_{n}$ from the multivariate normal distribution,
-$\text{MVN}\left\{\mathbf{0},\begin{pmatrix}0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ 0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ \vdots&\vdots&\ddots&\vdots\\ 0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&0.3\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}&\cdots&0.3\cdot\textbf{I}+0.7\cdot\mathbf{1}\mathbf{1}^{\mathsf{T}}\\ \end{pmatrix}\right\}.$ (12)
-Given features $\textbf{x}_{1},\ldots,\textbf{x}_{n}$, responses $y_{1},\ldots,y_{n}$ are simulated from the linear model
-$Y=\sum_{k=1}^{K}\textbf{X}_{B_{k}}^{\mathsf{T}}\boldsymbol{\beta}_{B_{k}}+e,\quad e\sim\text{N}(0,4^{2}),$ (13)
-where
-$\boldsymbol{\beta}_{B_{k}}=\begin{cases}(\textbf{1}_{2}^{\mathsf{T}},-0.2\cdot\textbf{1}_{3}^{\mathsf{T}})^{\mathsf{T}},&k=1,2,\\ (2.5,-1.2\cdot\textbf{1}_{2}^{\mathsf{T}},\textbf{0}_{2}^{\mathsf{T}})^{\mathsf{T}},&k=3,4,\\ (0.3\cdot\textbf{1}_{2}^{\mathsf{T}},\textbf{0}_{3}^{\mathsf{T}})^{\mathsf{T}},&k=5,\ldots,10\\ \textbf{0}_{5},&k=11,\ldots,50,\\ \end{cases}$ (14)
-Under (12)-(14), we have $H_{j}^{(\text{fg})}$’s are false for $j=51,\ldots,250$ and there are 28 important features.
-Empirical FDR and power of the FVG filter in identifying important features under different sample sizes are visualized in Figure 1.
-For comparison, we also evaluate the performance of model-X knockoff filter (Candès et al.,, 2018), group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019) in FDR control and identifying important features.
+Unlike the FDR control, asymptotic power of our FVG filter can’t be rigorously proved.
+The main reason is that our FVG filter is a model-free approach without rigid parametric assumption on the conditional distribution $Y|\textbf{X}$ and its power relies on the consistency between $Y|\textbf{X}$ and the feature importance score.
+However, when a linear model is correctly used, our FVG filter can achieve asymptotic power $1$.
+For example, following the proof of Zou2006, the feature importance scores $W_{j}$’s will be positive with asymptotic probability $1$ for those features with nonzero contribution to $Y$, making their asymptotic probability to be included in the selection set $\mathcal{R}^{(\text{fg})}$ as 1 under mild conditions.
+3 Simulated Experiments with Real-world Genetic Data
+To evaluate the proposed FVG filter in controlling FDR and identifying important features with comparison to existing approaches, we conduct extensive simulated experiments.
+3.1 Data Generation
+We simulate datasets based on a real-world genetic data to mimic the dependency structure among features in real-world genetic analysis and investigate how the proposed method performs in real applications.
+Specifically, we simulate datasets based on $6095$ variants in the APOE/APOC region (chromosome 19, positions 44909011 $\sim$ 45912650) from the whole-genome sequencing data (NG00067.v5) of the Alzheimer’s Disease Sequencing Project (Leung2019).
+In our study, we restrict our focus to $6952$ individuals with estimated European ancestry rate [^2] Estimated by SNPWeights v2.1 (Chen2013) using reference populations from the 1000 Genomes Consortium (10002015).
+greater than 80% and $p=1157$ variants $X_{1},\ldots,X_{p}$ whose minor allele frequency (MAF) is larger than $1\%$ and pairwise correlations are in [0.95,0.95][-0.95,0.95] to avoid the existence of statistically indistinguishable variants with nearly-perfect correlations.
+By randomly sampling $n$ individuals without replacement, each simulated dataset is obtained by collecting the sampled individuals’ variants and generating responses $y_{1},\ldots,y_{n}$ from the linear model
+$Y=\sum_{j=1}^{p}X_{j}{\beta}_{j}+e,\quad e\sim\text{N}(0,4),$ (12)
+where only $k=15$ randomly selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.6^{2})$.
+In our experiments, we simulate 1000 datasets for each possible sample size $n=500,1000$ and $2000$.
+To circumvent the threshold phenomenon mentioned in Gimenez2019b, we implement the multiple knockoff counterpart described in Appendix E.1.
+3.2 Details of Implementation and Evaluation Metrics
+Among these $1157$ variants, we compute the correlation $\text{cor}(X_{i},X_{j})$ between each pair of variants $(X_{i},X_{j})$ over all 6952 individuals and construct $345$ variant groups ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(X_{i},X_{j})|)_{p\times p}$.
+We define important features as variants $X_{j}$’s whose corresponding coefficients $\beta_{j}$’s are nonzero and important groups as those containing at least one important feature.
+Specifically, most important groups only contain one important feature.
+In summary, we have
+  • $H_{j}^{(\text{ff})}$’s are false for 15 important features;
+  • $H_{k}^{(\text{gg})}$’s are false for groups containing important features;
+  • $H_{j}^{(\text{fg})}$’s are false for features in those groups with false $H_{k}^{(\text{gg})}$’s.
+Based on groups ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ and the correlation matrix $\boldsymbol{\Sigma}=(\text{cor}(X_{i},X_{j}))_{1157\times 1157}$, we generate second-order group knockoffs under the maximum entropy (ME) construction using Algorithm 2 of Chu2024 and compute feature statistics as $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ with the joint lasso importance scores in the implementation of our FVG filter.
+To evaluate the power and precision of inference results produced by our FVG filter, we derive catching sets of signals based on the selection set $\mathcal{R}_{\text{FVG}}$ as
+$\mathcal{CC}_{\text{FVG}}=\{\mathcal{S}_{k,\text{FVG}}={B}^{(2)}_{k}\cap\mathcal{R}_{\text{FVG}}|k=1,\ldots,345\}.$
+For comparison, we also implement several existing methods as follows.
+  • Model-X knockoff filter: Based on the correlation matrix $\boldsymbol{\Sigma}$, we generate second-order knockoffs at the variant level under the maximum entropy (ME) construction and use the lasso coefficient difference statistics (Candes2018) in the implementation of model-X knockoff filter.
+Based on the selection set $\mathcal{R}_{\text{model-X}}$, we derive catching sets as
+    $\mathcal{CC}_{\text{model-X}}=\{\mathcal{S}_{j,\text{model-X}}=\{j\}|j\in\mathcal{R}_{\text{model-X}}\}.$
+  • Group knockoff filter: We implement the group knockoff filter based on the same ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ and the same second-order group knockoffs as our FVG filter.
+Based on the selection set $\mathcal{R}_{\text{group}}$, we derive catching sets as
+    $\mathcal{CC}_{\text{group}}=\{\mathcal{S}_{k,\text{group}}={B}^{(2)}_{k}|k\in\mathcal{R}_{\text{group}}\}.$
+  • KnockoffZoom and KeLP: For these methods which identify signals at multiple layers of different resolutions, we define three resolution layers as (a) layer 0: the level of variants; (b) layer 1: the level of variant groups ${B}^{(1)}_{1},\ldots,{B}^{(1)}_{650}$ obtained by the hierarchical clustering (average linkage with cutoff value $0.25$); (c) layer 2: the level of variant groups ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ which are also used in the FVG filter.
+Here, layer 2 corresponds to the same variant group level as the group knockoff filter while layer 1 is an intermediate layers between layer 2 and the level of variants.
+Based on such a hierarchical structure of variant groups that each variant group at high-resolution levels is nested under one group at low-resolution levels, we implement both KnockoffZoom and KeLP at layers 0 and $2$ (2 layers) and layers 0, 11 and $2$ (3 layers).
+Thus, both KnockoffZoom and KeLP are tackling the same task as our FVG filter to refine those important variant groups in layer 2.
+For each layer, we generate second-order group knockoffs under the maximum entropy (ME) construction using Algorithm 2 of Chu2024.
+After implementing either KnockoffZoom or KeLP, we obtain catching sets as the most specific discoveries among all discoveries in selection sets (same as Figure 1 of Gablenz2024) as detailed in Appendix F.
+Specifically, to avoid obtaining empty catching sets, we implement KeLP in our article with carefully chosen tuning parameters as detailed in Appendix G.
+In the following section, we evaluate the empirical FDR of our FVG filter via (4), empirical power of all approaches as the proportion of important features (with nonzero coefficients) included in at least one catching sets.
+We also measure the informativeness of different approaches by measuring the average size and purity [^3] Purity refers to the minimum absolute correlation within the catching set.
+of catching sets.
+3.3 Performance Comparison of the FVG Filter with Existing Competitors
+Refer to captionRefer to caption
 (a) $n=500$.
+Refer to captionRefer to caption
 (b) $n=1000$.
+Refer to captionRefer to caption
 (c) $n=2000$.
-Figure 1: Empirical FDR and power of FVG filter with respect to the target FDR level ($\alpha$) over 1000 simulated datasets of different sample sizes in comparison to model-X knockoff filter (Candès et al.,, 2018), group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019).
-It is clear that the FVG filter manages to control $\text{FDR}^{(\text{fg})}$ under any target level $\alpha$.
-For a fixed target level $\alpha$, we find the empirical power increases to $1$ as $n$ grows, suggesting that the FVG filter is consistent in identifying important features.
-Although existing methods can also control FDR and have growing power, they suffer different types of performance deterioration in identifying important features.
-On one hand, group knockoff filter suffers power loss when the target FDR level is lower than $0.10$ due to the threshold issue mentioned in Gimenez and Zou, (2019).
-Even with power improvement as the target FDR level grows from $0.10$ and $0.20$, the informativeness of rejection sets obtained by group knockoff filter is poor as all important groups ($B_{1},\ldots,B_{10}$) are rejected as a whole without distinguishing which features are more important.
-As a comparison, the average size of nonempty catching sets $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ ($k=1,\ldots,50$) obtained by FVG filter is less than $2$ under the target level $\alpha=0.1$.
-On the other hand, although the multilayer knockoff filter has power improvement over model-X knockoff filter by focusing important feature selection on those identified feature groups, there remains a significant power gap between FVG filter and the multilayer knockoff filter when the target FDR is large.
-3.2 Simulated Experiments with Real-world Genetic Data
-To mimic the dependency structure among features in real-world genetic analysis and investigate how the proposed method performs in real applications, we simulate 500 datasets based on $6095$ variants in the APOE/APOC region (chromosome 19, positions 44909011 $\sim$ 45912650) from the whole-genome sequencing data (NG00067.v5) of the Alzheimer’s Disease Sequencing Project (Leung et al.,, 2019).
-Specifically, we restrict our study to $6952$ individuals with estimated European ancestry rate [^3] Estimated by SNPWeights v2.1 (Chen et al.,, 2013) using reference populations from the 1000 Genomes Consortium (The 1000 Genomes Project Consortium,, 2015).
-greater than 80% and $p=1000$ variants $X_{1},\ldots,X_{p}$ whose minor allele frequency (MAF) is larger than $1\%$ and pairwise correlations are in $[-0.95,0.95]$ to avoid the existence of statistically indistinguishable variants with nearly-perfect correlations.
-Among these variants, we compute the correlation $\text{cor}(X_{i},X_{j})$ between any pair of variants $(X_{i},X_{j})$ over all 6952 individuals and construct variant groups ${B}_{1},\ldots,{B}_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(X_{i},X_{j})|)_{p\times p}$.
-By randomly sampling $n=5000$ individuals without replacement, each simulated dataset is obtained by collecting the sampled individuals’ variants and generating responses $y_{1},\ldots,y_{n}$ from the linear model
-$Y=\sum_{j=1}^{p}X_{j}{\beta}_{j}+e,\quad e\sim\text{N}(0,4),$ (15)
-where only $k=10$ randomly selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.225)$.
-As a result, important features refer to variants $X_{j}$’s whose corresponding coefficients $\beta_{j}$’s are nonzero and important groups are those containing at least one important feature.
-Specifically, most important groups only contain one important feature.
-Figure 2: Empirical performance of the model-X knockoff filter (Candès et al.,, 2018), the group knockoff filter (Dai and Barber,, 2016), the multilayer knockoff filter (Katsevich and Sabatti,, 2019) and our FVG filter over 500 simulated data: (a) empirical FDR and power with respect to the target FDR level; (b) average size of catching sets; (c) average purity of catching sets.
-To illustrate the empirical benefit of the FVG filter, we apply it to find important features in these 500 simulated datasets with comparison to the model-X knockoff filter (Candès et al.,, 2018), the group knockoff filter (Dai and Barber,, 2016) and the multilayer knockoff filter (Katsevich and Sabatti,, 2019).
-Empirical FDR and power of all methods are presented in Figure 2 (a).
-For fair comparison, we use the joint lasso feature importance score when implementing both existing methods and our algorithm.
-It is found that our FVG filter can empirically control FDR.
-Compared to the model-X knockoff filter which can only distinguish around $50\%$ of important features when the target FDR level is relaxed to $20\%$, our FVG filter, the multilayer knockoff filter and the group knockoff filter are able to identify about $70\%\sim 80\%$ of important features (or feature groups) when the target FDR level is $10\%$.
-However, the inference results given by the multilayer knockoff filter and the group knockoff filter are far from being informative and precise.
-To illustrate this, we compute catching set $\mathcal{R}_{k}$ of each variant group from the rejection set $\mathcal{R}$, where
-  • $\mathcal{R}_{k}=B_{k}$ if $H^{\text{(gg)}}_{k}$ is rejected by the group knockoff filter;
-  • $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ if the rejection set $\mathcal{R}$ is obtained from the model-X knockoff filter or our FVG filter.
-  • For each group $B_{k}$, we let
-    • $\mathcal{R}_{k}=B_{k}$ if $H^{\text{(gg)}}_{k}$ but all $H^{\text{(ff)}}_{j}$’s ($j\in B_{k}$) is rejected by the multilayer knockoff filter;
-    • $\mathcal{R}_{k}=B_{k}\cap\mathcal{R}$ if $H^{\text{(gg)}}_{k}$ and some $H^{\text{(ff)}}_{j}$’s ($j\in B_{k}$) are rejected by the multilayer knockoff filter.
-We then calculate the average size and purity (minimum absolute correlation) of catching sets to measure the informativeness and precision of $\mathcal{R}$ given by different filters in Figure 2 (b)-(c).
-Here, it is desired that sizes of catching sets are smaller as more unimportant variants are excluded.
-Even when there remain multiple features in catching sets, high purity is desirable to exclude unimportant variants weakly correlated to the important ones (or poor-quality proxy variants).
-In contrast to the group knockoff filter whose average size and purity of catching sets are almost $6$ and smaller than $20\%$ respectively (which means that catching sets obtained by the group knockoff filter contain a lot of poor-quality proxy variants), the precision of our FVG filter in pinpointing true important variants is close to the model-X knockoff filter and is higher than the multilayer knockoff filter.
-With an average size close to $1$ and average purity greater than $80\%$, our FVG filter can directly obtain informative catching sets containing only one or two features in most cases.
-In summary, our FVG filter manage to simultaneously achieve high power and high informativeness in identifying important features.
+Figure 2: Empirical FDR and power of FVG filter with respect to the target FDR level ($\alpha$) over 1000 simulated genetic datasets of different sample sizes in comparison to model-X knockoff filter (Candes2018), group knockoff filter (Dai2016), the KnockoffZoom (Sesia2020) and the KeLP (Gablenz2024).
+Refer to captionRefer to caption
+(a) $n=500$.
+Refer to captionRefer to caption
+(b) $n=1000$.
+Refer to captionRefer to caption
+(c) $n=2000$.
+Figure 3: Average sizes and purity of catching sets obtained FVG filter over 1000 simulated genetic datasets of different sample sizes in comparison to model-X knockoff filter (Candes2018), group knockoff filter (Dai2016), the KnockoffZoom (Sesia2020) and the KeLP (Gablenz2024).
+Over 1000 simulated datasets of different sample sizes, the empirical FDR of our FVG filter and its power in comparison with all existing competitors under different target FDR levels are reported in Figure 2.
+We also present the average sizes and purity of catching sets produced by different methods in Figure 3.
+It is clear from Figure 2 that the FVG filter controls $\text{FDR}^{(\text{fg})}$ at any target level $\alpha$.
+Existing methods suffer different types of performance deterioration in identifying important features.
+On one hand, the model-X knockoff filter suffers great power loss in Figure 2 because high correlations of variants in the same group lower signal-to-noise ratio of false $H^{\text{(ff)}}_{j}$’s.
+On the other hand, although group knockoff filter manages to achieve higher power in Figure 2, such an advantage comes from the sacrifice of informativeness of rejection sets.
+This can be seen from catching sets with average size of around $6$ and average purity of around $0.6$ in Figure 3 because variant groups are rejected as a whole without distinguishing which variants are more important.
+Such a poor precision of group knockoff filter can be somehow relieved by both KnockoffZoom and KeLP, which however still suffer lack of precision.
+Specifically, KnockoffZoom (2 layers) manages to maintain similar power as the group knockoff filter as shown in Figure 2.
+By decreasing the average size from $6$ to $3$ and increasing average purity from $0.6$ to $0.85$ when $n=2000$, the improvement of precision provided by KnockoffZoom is still limited.
+Adding the intermediate layer (layer 1) in KnockoffZoom only slightly improves precision of catching sets (average size of around $2.5$ and average purity of around $0.89$) without any power gain.
+KeLP, on the other hand, improves the precision of catching sets much greater than KnockoffZoom.
+With the same 2 layers in analysis, KeLP (2 layers) returns catching sets of average sizes of around $2.3$ and average purity of around $0.9$ when $n=2000$ (except for target FDR level $0.05$ where all catching sets are of size 1).
+Adding the intermediate layer (layer 1) could further refine catching sets to average size of $1.7$ and purity of $0.95$.
+However, the power of KeLP is significantly lower especially under lower target FDR levels.
+In contrast, by simultaneously selecting important groups and refining the catching sets, the FVG filter return catching sets that are smaller and purer without significant power loss in identifying important genetic variants.
+As sample size $n$ grows, FVG filter maintains catching sets of small size (around $1.1$) and high purity (higher than $0.97$) for target FDR levels $0.05$, $0.1$, $0.15$, $0.20$.
+With minor power loss with respect to group knockoff filter, our FVG filter can locate the real important variants from important groups in informative catching sets most of the time.
+On the contrary, variant-level (layer 0) inferences of KnockoffZoom and KeLP still rely on the same variant-level knockoff of model-X knockoff filter.
+With low power at layer 0, a non-negligible fraction of catching sets are of size larger than $1$, leading to both large average size and lower average purity.
+Such an obstacle is elaborately circumvented by our FVG filter which uses group knockoffs to ease the identification of signals and lasso regression to refine the catching sets.
+3.4 Multiple Signals in the Same Group
+In the simulation setting of Section 3.1, there is only one signal within each nonnull group in most of time.
+As a result, heuristic identifying the contributing variants in nonnull groups are not that difficult.
+For example, one can always perform heuristic selection of contributing variants from variant groups identified by the group knockoff filter.
+However, such an identification task becomes nontrivial when there are multiple genetic signals in one variant group.
+To demonstrate the performance of the FVG filter in prioritizing important variants under such scenarios, we conduct more experiments with two settings of signal distributions, including:
+  • Setting A: $k=15$ selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.6^{2})$, where there is a randomly selected variant group containing 4 signals, two randomly selected groups containing 2 signals and seven randomly selected groups containing 1 signal;
+  • Setting B: $k=20$ selected coefficients ${\beta}_{j}$’s are nonzero and follow $N(0,0.6^{2})$, where there are ten randomly selected groups containing 2 signals.
+Under both settings, we simulate 1000 datasets of sample size $n=2000$. We compare our FVG filter with two heuristic approaches.
+  • Top-1 signal: Within each variant group selected by the group knockoff filter, we construct the catching set by heuristically including only the variant with the largest $|W_{j}|$’s.
+  • SuSiE: Within each variant group selected by the group knockoff filter, we implement the SuSiE method (Wang2020) and construct the catching set as the $95\%$ credible set.
+Performance of our FVG filter and two heuristic approaches are presented in Figure 4.
+We find that the FVG filter still achieves a balance between the informativeness of catching sets and power.
+On one hand, selecting only the top-1 signal within identified variant groups provides the most informative catching sets (with size 1 and purity 1).
+However, doing so may miss many signals, especially under setting B.
+On the other hand, using SuSiE within selected variant groups usually returns too large catching sets.
+In contrast, our FVG filter not only returns informative catching sets (average size <1.5<1.5 and average purity $\approx 0.9$) but also has similar power as SuSiE.
+That is to say, among all genetic variants selected by SuSiE that are not top signals in the groups, our FVG filter adaptively selects those contributing variants without proxy ones.
+Remark 2.
+Although all the simulations are conducted under the linear model, the FDR control of our FVG filter still stands for other conditional distribution $Y|\textbf{X}$.
+The reason is that our FVG filter is developed under the similar framework of model-X knockoff, where no assumption is imposed on $Y|\textbf{X}$.
+Thus, even $Y|\textbf{X}$ follows a logistic regression model, using the lasso feature importance scores would not violate our FVG filter’s FDR control.
+However, using a model that accurately depicts $Y|\textbf{X}$ to compute feature importance scores would further improve the power.
+Refer to captionRefer to captionRefer to caption
+(a) Setting A.
+Refer to captionRefer to captionRefer to caption
+(b) Setting B.
+Figure 4: Average power, sizes and purity of catching sets obtained by FVG filter over 1000 simulated genetic datasets of different settings in comparison to two heuristic approaches.
 4 AD Analysis of EADB–UKBB Dataset
-Given the promising performance of the proposed filter in simulated experiments, we apply our method to the EADB-UKBB dataset where model-X knockoff filter exhibits power loss and the group knockoff filter returns not informative enough results in Section 1.2.
-To infer which variants are associated with AD, we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ and construct variant groups $B_{1},\ldots,B_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$.
-By doing so, we obtain the same $321,535$ groups $B_{1},\ldots,B_{k}$ as the group knockoff filter in Section 1.2 where the average absolute correlation of variants from different groups are lower than 0.5.
-Based on the correlation matrix $(\text{cor}(G_{i},G_{j}))_{p\times p}$ and variant groups $B_{1},\ldots,B_{K}$, we generate group knockoffs $\widetilde{\mathbb{G}}$ under the maximum entropy (ME) construction of Chu et al., (2023).
+Given the promising performance of the proposed filter in simulated experiments, we apply it to the EADB-UKBB dataset where model-X knockoff filter and KeLP exhibits power loss and the group knockoff filter and KnockoffZoom returns not informative enough results in Section 1.2.
+To infer which variants are associated with AD, we first compute correlation $\text{cor}(G_{i},G_{j})$ between any pair of variants $G_{i}$ and $G_{j}$ using the UK Biobank directly genotyped data as the reference panel and construct variant groups $B_{1},\ldots,B_{K}$ by applying the hierarchical clustering (average linkage with cutoff value $0.5$) on the distance matrix $(1-|\text{cor}(G_{i},G_{j})|)_{p\times p}$.
+By doing so, we obtain the same $321,569$ groups $B_{1},\ldots,B_{k}$ as the group knockoff filter in Section 1.2 where the average absolute correlation of variants from different groups are lower than 0.5.
+Based on the correlation matrix $(\text{cor}(G_{i},G_{j}))_{p\times p}$ and variant groups $B_{1},\ldots,B_{K}$, we generate second-order group knockoffs $\widetilde{\mathbb{G}}$ under the maximum entropy (ME) construction using Algorithm 2 of Chu2024 by assuming that the distribution of X can be well approximated by a multivariate Gaussian distribution.
 Given the response of interest $\textbf{y}=(y_{1},\ldots,y_{n})^{\mathsf{T}}$ where $y_{i}=1$ if the $i$-th observation corresponds to a clinically diagnosed AD case and $y_{i}=0$ otherwise, we use absolute values of lasso estimators of the logistic regression model,
@@
 We then calculate feature statistics $W_{j}=\text{sign}\left(T_{j}-\widetilde{T}_{j}\right)\cdot\max\left(T_{j},\widetilde{T}_{j}\right)$ and implement the FVG filter with target FDR level $\alpha=0.10$.
-Variants in the rejection set $\mathcal{R}^{(\text{fg})}$ under target FDR level $\alpha=0.10$ are presented in Figure 3 with their post-selection importance scores $W_{j}\cdot\text{I}(W_{j}>0)$ and names of the closest genes.
 Figure 5: Manhattan plot of genetic variants under our FVG filter. Identified variants under target FDR level $\alpha=0.10$ are highlighted with names of their closest genes.
-Details of variants identified by the FVG filter under the target FDR level $0.10$ are provided in Table 3.
-In total, we identify 205 variants that contribute to AD variation substantially, which comes from 154 groups in 84 loci.
+Table 3: Details of variants identified by the FVG filter under the EADB–UKBB dataset and target FDR level $\alpha=0.10$.
+Here, the column “cS2Gene” denotes the gene that each genetic variant functionally annotates, and the column “Annotation” depicts the type of annotation.
+Identified Position cS2Gene Annotation Identified Position cS2Gene Annotation Identified Position cS2Gene Annotation
+Variant Variant Variant
+Chromosome 1 rs3740890 130385218 rs419010 44865063 PVRL2 GTeX
+rs2305463 20853688 EIF4G3 Exon Chromosome 12 rs365653 44858389
+rs782791 203029691 PPFIA4 GTeX,ABC rs7183 7070715 C1S Exon rs17561351 44869072 TOMM40 eQTLGen,ABC
+rs4844610 207629207 CR1 GTeX rs17860282 51346566 CELA1 Promoter rs73052307 44881148 APOE EpiMap
+Chromosome 2 Chromosome 13 rs11668327 44895376
+rs35933838 9544047 CPSF3 eQTLGen rs9516168 93085490 rs449647 44905307 APOE Promoter,ABC
+rs6726709 37270395 Chromosome 14 rs41290122 44880326
+rs77059113 43445369 rs17125924 52924962 rs10420036 44882772 APOE EpiMap
+rs113068192 76786226 rs12590654 92472511 rs283808 44883777 PVRL2 GTeX
+rs116038905 105805908 rs4904929 92470949 rs283813 44885917 PVRL2 Exon
+rs74851408 117201103 rs74093831 105733666 CRIP1 EpiMap rs7254892 44886339 APOE EpiMap=0.500
+rs6733839 127135234 ERCC3 Cicero rs873533 106667442 rs7412 44908822 APOE Exon
+rs744373 127137039 BIN1 GTeX Chromosome 15 rs445925 44912383 APOE GTeX
+rs2118506 127139927 BIN1 EpiMap rs12595082 50715532 rs390082 44913574 APOC1 Promoter,EpiMap
+rs4954187 134842848 rs2043085 58388755 LIPC GTeX rs190712692 44921921
+rs6725887 202881162 rs593742 58753575 rs141622900 44923535
+rs10933431 233117202 NGEF GTeX rs16946801 63312881 CA12 GTeX rs41290100 44867684
+rs7421448 233117495 INPP5D EpiMap rs117618017 63277703 APH1B Exon,GTeX,eQTLGen,ABC rs187183066 44872328 CEACAM22P GTeX
+Chromosome 3 rs56117790 79075361 rs41290108 44874585 PVRL2 Promoter
+rs4974180 56200107 Chromosome 16 rs73033507 44928146 APOC4 EpiMap
+rs7634492 136105288 SLC35G2 GTeX rs12325539 30022312 DOC2A Promoter rs79149284 45010596 DMWD GTeX
+rs16824536 155069722 rs11865499 31120929 KAT8 GTeX,eQTLGen rs77196615 44877078 ZNF296 EpiMap
+rs61762319 155084189 MME Exon rs78924645 31143037 PRSS36 Promoter,GTeX rs79701229 44881674 PVRL2 Promoter
+Chromosome 4 rs8058370 81738205 PLCG2 Promoter rs3745150 44882502 TOMM40 GTeX
+rs6448453 11024404 rs12446759 81739398 PLCG2 Promoter,Cicero rs157580 44892009 TOMM40 Promoter
+rs13116567 12239152 rs11548656 81883307 PLCG2 Exon rs2075649 44892073 TOMM40 Promoter,GTeX,eQTLGen
+rs502746 14019593 rs1071644 81937798 PLCG2 Exon rs8106922 44898409
+rs116791081 143428212 rs56407236 90103687 rs405697 44901434 TOMM40 Exon
+Chromosome 5 Chromosome 17 rs7259620 44904531 APOC1 GTeX
+rs25992 14707491 ANKH Exon,Cicero rs7225151 5233752 SCIMP GTeX,ABC,Cicero rs405509 44905579 APOE Promoter,ABC
+rs10068419 86923485 rs9901675 7581494 CD68 Exon rs440446 44905910 APOE Exon,ABC
+rs7268 140332965 HBEGF Exon rs79759095 44451394 rs142042446 44883210 APOE EpiMap
+rs335473 177559423 RAB24 EpiMap rs199451 46724418 KANSL1-AS1 GTeX rs12972970 44884339 TOMM40 ABC
+rs4700716 180138792 RASGEF1C Promoter rs616338 49219935 ABI3 Exon,EpiMap rs283815 44887076
+Chromosome 6 rs2632516 58331728 BZRAP1 ABC rs6857 44888997 PVRL2 Exon
+rs78799498 17319361 RBM24 ABC rs9896864 63458947 rs71352238 44891079 TOMM40 Promoter,GTeX,ABC
+rs2130357 27919052 PGBD1 GTeX rs4277405 63471557 ACE GTeX rs184017 44891712 TOMM40 Promoter,ABC
+rs3132451 31614248 AIF1 Promoter,EpiMap rs4309 63482562 ACE Exon,EpiMap rs2075650 44892362
+rs532965 32610196 HLA-DRB5 ABC rs3730025 63480412 ACE Exon rs157581 44892457 TOMM40 Exon
+rs9469112 32447376 Chromosome 19 rs34095326 44892587 BCAM GTeX
+rs9271162 32609938 HLA-DRB1 EpiMap rs4147918 1058177 ABCA7 Exon rs34404554 44892652
+rs9273349 32658092 19:1051137_CTG_C 1051138 ABCA7 Exon,GTeX,EpiMap,ABC rs157582 44892962 TOMM40 Promoter
+rs9275184 32686937 HLA-DQA2 EpiMap rs3752246 1056493 ABCA7 Exon,GTeX rs10119 44903416 TOMM40 Exon,GTeX
+rs3997854 32715138 TAP2 GTeX rs35917007 1849148 REXO1 Promoter,GTeX rs769449 44906745 APOE Promoter,ABC
+rs114812713 41066261 OARD1 Exon rs10404195 5908958 VMAC Exon rs75627662 44910319 SNRPD2 GTeX
+rs7748513 41160234 TREML2 GTeX rs60184524 43790063 rs7256200 44912678 APOC1 GTeX,EpiMap,ABC
+rs2093395 41187288 TREML2 EpiMap rs445752 43956053 rs438811 44913484 APOC1 Promoter,ABC
+rs143332484 41161469 TREM2 Exon,EpiMap rs2125579 44288548 ZNF235 Exon,GTeX,eQTLGen rs12721046 44917997
+rs9394778 41247320 TREML2 EpiMap rs8107530 44332247 rs12721051 44918903 APOC1 Exon
+rs3997700 41251889 rs62116778 44389907 rs111789331 44923868
+rs2171089 47548075 CD2AP GTeX rs151111866 44387243 ZNF285 Exon rs73936968 44892559
+rs117591349 78478451 rs2571174 44430314 ZNF229 Exon rs439401 44911194 APOE GTeX
+rs313161 85332752 rs17800789 44510962 rs59325138 44913034 APOC1P1 GTeX,Cicero
+rs976271 114361563 rs7254925 44513811 rs584007 44913221 APOE GTeX
+Chromosome 7 rs74602839 44481157 rs157595 44922203
+rs10952168 1554792 TMEM184A Promoter,GTeX,ABC rs62116891 44524436 rs1064725 44919304 APOC1 Exon
+rs12669393 7815748 rs930461 44539272 CEACAM22P Exon rs118060185 44928417 ZNF155 GTeX
+rs3173615 12229791 TMEM106B Exon rs16979269 44552920 ZNF180 GTeX rs144311893 44920687
+rs2189965 28132395 rs56757528 44561286 rs4803771 44924391 EML2 GTeX
+rs4917014 50266267 IKZF1 EpiMap,Cicero rs10405030 44565601 rs60049679 44926451 APOC1P1 Promoter
+rs1476679 100406823 LAMTOR4 GTeX rs2357100 44595918 rs114533385 44933496 ZNF226 GTeX
+rs41280968 100105324 AP4M1 Exon,ABC rs7259313 44600383 rs79429216 44942260 APOC4 Exon,GTeX,EpiMap
+rs221834 100745552 UFSP1 Cicero rs17658507 44617986 PVR GTeX rs5167 44945208 APOC4 Exon
+rs144867634 111940111 DOCK4 Exon rs80122548 44618415 rs3760627 44953923 CLPTM1 Promoter
+rs12703526 143410495 EPHA1-AS1 ABC rs62119261 44618732 APOC4 EpiMap rs16979595 44974124
+rs7805776 143427203 TAS2R60 eQTLGen rs203717 44636512 IGSF23 Exon rs73560271 45051621 ZNF296 EpiMap
+Chromosome 8 rs2965160 44692865 CEACAM19 eQTLGen rs78620885 45087826 PPP1R37 ABC
+rs4731 11808828 FDFT1 Exon,ABC rs2965164 44698782 CEACAM16 Promoter rs117106519 45047229 CLASRP ABC
+rs73223431 27362470 PTK2B GTeX,eQTLGen rs62120574 44711506 rs3178166 45090912 GEMIN7 Exon,GTeX,eQTLGen
+rs9297949 94957217 NDUFAF6 Promoter rs76419583 44714590 CEACAM22P GTeX rs10405086 45123977
+rs1693551 100663356 SNX31 Exon,GTeX rs2965109 44722081 BCL3 ABC rs2532037 45194545
+rs2436889 102564430 rs1004165 44728939 rs3848526 45195982 SYMPK GTeX
+rs79832570 144042819 PARP10 GTeX rs2965169 44747899 BCL3 Exon,ABC rs10411314 45224801 EXOC3L2 Exon
+rs61732533 144053248 OPLAH Exon rs116480598 44753996 BCL3 ABC rs17356664 45237513
+rs34674752 144099319 SHARPIN Exon,ABC=0.500 rs114294368 44789899 rs77740243 45248200
+Chromosome 9 rs75599500 44740439 BCL3 EpiMap rs11559024 45317925 CKM Exon
+rs1131773 93077974 SUSD3 Exon,GTeX,EpiMap rs2965174 44741758 CBLC GTeX rs17875640 45323890 CKM Promoter,ABC
+rs1883025 104902020 rs10408993 44802496 rs73047784 45458317
+Chromosome 10 rs117326714 44783846 rs3916883 45352060 ERCC2 ABC
+rs7920721 11678309 USP6NL EpiMap,ABC,Cicero rs28399654 44813331 BCAM Exon rs4802253 45401737 PPP1R13L ABC
+rs7068614 11679478 rs117142879 44802327 rs2336219 45409148 PPP1R13L GTeX
+rs117980033 29966853 rs28399637 44820881 BCAM Promoter,GTeX,EpiMap,ABC rs7256865 45515149 RTN2 eQTLGen
+rs1227759 69831131 COL13A1 GTeX,eQTLGen rs10405693 44823407 BCAM EpiMap rs73942917 45543797 DMWD Cicero
+rs6586028 80494228 TSPAN14 GTeX,ABC rs149529419 44868071 rs1264226 45559909 OPA3 GTeX
+Chromosome 11 rs140684051 44896199 rs73568973 45549792 OPA3 Exon
+rs1685404 47222114 rs73572039 44839373 PVRL2 GTeX rs73570937 45588621 GPR4 EpiMap
+rs1582763 60254475 rs76692773 44890954 TOMM40 Promoter,ABC rs75987583 45745299 DMWD GTeX
+rs11824773 60309467 MS4A6E Cicero rs61642202 44842026 ZNF224 GTeX rs875121 49948251
+rs536841 86076782 rs117310449 44890259 TOMM40 Promoter,ABC rs2075803 51125272 SIGLEC9 Exon,eQTLGen
+rs10792832 86156833 PICALM ABC rs77301115 44893716 TOMM40 Promoter,eQTLGen rs1710398 51223655 CD33 GTeX
+rs3844143 86139201 EED EpiMap rs79398853 44895528 rs3865444 51224706 CD33 Promoter
+rs74685827 121482368 SORL1 ABC,Cicero rs115881343 44899959 TOMM40 eQTLGen rs731170 54664811 LILRB4 Exon
+rs2298813 121522975 SORL1 Exon rs41289512 44848259 CLASRP Cicero Chromosome 20
+rs3781837 121578263 rs12691088 44915229 APOC1 Exon,ABC rs6064392 56409712 CASS4 GTeX,eQTLGen,EpiMap,ABC
+rs12272618 121589615 SORL1 Promoter rs3865427 44877704 PVRL2 GTeX rs6069749 56446519 CASS4 GTeX
+Variants in the rejection set $\mathcal{R}^{(\text{fg})}$ under target FDR level $\alpha=0.10$ are presented in Figure 5 with their positive part of importance scores $W_{j}\cdot\text{I}(W_{j}>0)$’s and names of the closest genes.
+Details of these variants are provided in Table 3.
+In total, we identify 266 variants that contribute to AD variation substantially, which comes from 198 groups in 89 loci.
 Similar to the literature, multiple variants are identified in the APOE/APOC region with the strongest association to AD (chromosome 19, positions 44909011 $\sim$ 45912650).
-In addition, we also manage to identify variants rs744373 and rs2118506 close to gene “BIN1”, variant rs1532278 close to gene “CLU”, variant rs10792832 close to gene “PICALM”, variants rs4277405, rs4309 and rs3730025 close to gene “ACE”, all of which are also reported in He et al., (2022) and Bellenguez et al., (2022).
-Compared with the model-X knockoff filter, the FVG filter only misses 2 loci identified by the model-X knockoff filter but discover 37 more loci, which suggests that our method would not suffer from the same power loss issue of the model-X knockoff filter.
-Table 3: Details of variants identified by the FVG filter under the EADB–UKBB dataset and target FDR level $\alpha=0.10$.
-Identified Position Closest Identified Position Closest Identified Position Closest Identified Position Closest
-Variant Gene Variant Gene Variant Gene Variant Gene
-Chromosome 1 rs3997700 41251889 TREML5P rs9516168 93085490 LINC00363 rs3852860 44879709 NECTIN2
-rs11579514 161173522 B4GALT3 rs9296558 47484147 CD2AP Chromosome 14 rs365653 44858389 NECTIN2
-rs782791 203029691 PPFIA4 rs10948363 47520026 CD2AP rs74093831 105733666 ELK2AP rs73052307 44881148 NECTIN2
-rs4562624 207512620 CR1 rs12525381 696327 PTPRK rs873533 106667442 LINC00221 rs11668327 44895376 TOMM40
-rs3818361 207611623 CR1 rs117591349 78478451 MIR10524 rs17125924 52924962 FERMT2 rs449647 44905307 APOE
-rs6701713 207612944 CR1 rs313161 85332752 LINC02535 rs4904929 92470949 SLC24A4 rs10420036 44882772 NECTIN2
-rs4844610 207629207 CR1 Chromosome 7 Chromosome 15 rs283813 44885917 NECTIN2
-Chromosome 2 rs221834 100745552 ZAN rs59685680 50709337 SPPL2A rs7254892 44886339 NECTIN2
-rs116038905 105805908 NCK2 rs144867634 111940111 DOCK4 rs28470808 50735410 SPPL2A rs7412 44908822 APOE
-rs74851408 117201103 DDX18 rs10952168 1554792 TMEM184A rs2043085 58388755 LIPC rs445925 44912383 APOC1
-rs6733839 127135234 BIN1 rs56402156 143406388 TCAF2 rs593742 58753575 ADAM10 rs390082 44913574 APOC1
-rs730482 127136908 BIN1 rs12703526 143410495 TCAF2 rs117618017 63277703 APH1B rs190712692 44921921 APOC1
-rs744373 127137039 BIN1 rs9640386 143435090 TCAF2 rs3816385 64126456 SNX1 rs41290108 44874585 NECTIN2
-rs745717 127136321 BIN1 rs2189965 28132395 JAZF1 rs56117790 79075361 RASGRF1 rs73033507 44928146 APOC1P1
-rs2118506 127139927 BIN1 rs2122623 37842715 NME8 Chromosome 16 rs79701229 44881674 NECTIN2
-rs10496731 134840058 ACMSD rs7809394 55077689 EGFR rs12325539 30022312 C16orf92 rs157580 44892009 TOMM40
-rs4954187 134842848 ACMSD rs12669393 7815748 UMAD1 rs11865499 31120929 KAT8 rs2075649 44892073 TOMM40
-rs3819121 134865290 ACMSD rs4725076 8204779 ICA1 rs8058370 81738205 CMIP rs8106922 44898409 TOMM40
-rs6725887 202881162 WDR12 rs1476679 100406823 ZCWPW1 rs12446759 81739398 CMIP rs405697 44901434 TOMM40
-rs10933431 233117202 INPP5D rs41280968 100105324 AP4M1 rs11548656 81883307 PLCG2 rs7259620 44904531 TOMM40
-rs7421448 233117495 INPP5D rs13228694 100342684 STAG3L5P-PVRIG2P-PILRB rs1071644 81937798 PLCG2 rs440446 44905910 APOE
-rs35349669 233159830 INPP5D Chromosome 8 rs56407236 90103687 FAM157C rs6857 44888997 NECTIN2
-rs3771371 71400409 ZNF638 rs2436889 102564430 ODF1 Chromosome 17 rs2075650 44892362 TOMM40
-rs113068192 76786226 LRRTM4 rs4731 11808828 FDFT1 rs79759095 44451394 GPATCH8 rs157581 44892457 TOMM40
-Chromosome 3 rs79832570 144042819 SPATC1 rs5910 44372421 ITGA2B rs34404554 44892652 TOMM40
-rs7634492 136105288 PPP2R3A rs61732533 144053248 OPLAH rs199439 46716137 NSF rs75627662 44910319 APOE
-rs16824536 155069722 MME rs34674752 144099319 SHARPIN rs199451 46724418 NSF rs438811 44913484 APOC1
-rs61762319 155084189 MME rs73223431 27362470 PTK2B rs2671659 49397167 PHB rs283814 44885967 NECTIN2
-rs4974180 56200107 ERC2 rs1532278 27608798 CLU rs616338 49219935 ABI3 rs439401 44911194 APOE
-Chromosome 4 rs9297949 94957217 NDUFAF6 rs2072153 49312652 ZNF652 rs59325138 44913034 APOC1
-rs6448453 11024404 CLNK Chromosome 9 rs2526378 58326988 TSPOAP1-AS1 rs584007 44913221 APOC1
-rs13116567 12239152 LINC02270 rs1131773 93077974 SUSD3 rs2632516 58331728 TSPOAP1-AS1 rs157595 44922203 APOC1
-rs502746 14019593 LINC01182 Chromosome 10 rs9896864 63458947 CYB561 rs4803770 44924096 APOC1P1
-rs116791081 143428212 GAB1 rs7920721 11678309 ECHDC3 rs4277405 63471557 ACE rs71352239 44926286 APOC1P1
-Chromosome 5 rs7068614 11679478 ECHDC3 rs4309 63482562 ACE rs118060185 44928417 APOC1P1
-rs25992 14707491 ANKH rs10764258 20634336 MIR4675 rs3730025 63480412 ACE rs144311893 44920687 APOC1
-rs871269 151052827 TNIP1 rs117980033 29966853 JCAD rs9901675 7581494 CD68 rs114533385 44933496 APOC1P1
-rs335473 177559423 FAM193B rs1227759 69831131 COL13A1 Chromosome 19 rs79429216 44942260 APOC4-APOC2
-rs4700716 180138792 RASGEF1C rs6586028 80494228 TSPAN14 rs445752 43956053 ZNF221 rs12721109 44943964 APOC4-APOC2
-rs10068419 86923485 LINC02059 Chromosome 11 rs2571174 44430314 ZNF229 rs16979595 44974124 CLPTM1
-Chromosome 6 rs74685827 121482368 SORL1 rs2293164 44475997 ZNF180 rs10411314 45224801 EXOC3L2
-rs976271 114361563 HDAC2-AS2 rs2298813 121522975 SORL1 rs1625502 44540845 CEACAM22P rs75987583 45745299 MEIOSIN
-rs2130357 27919052 OR2B2 rs3781837 121578263 SORL1 rs10405071 44732632 BCL3 rs875121 49948251 SIGLEC11
-rs13408 28244970 ZKSCAN4 rs3740890 130385218 ZBTB44-DT rs2965169 44747899 BCL3 rs1710398 51223655 CD33
-rs3130931 31167111 POU5F1 rs17439299 15599040 LINC02751 rs116480598 44753996 BCL3 rs3865444 51224706 CD33
-rs3132451 31614248 AIF1 rs1685404 47222114 DDB2 rs28399654 44813331 BCAM rs12974390 54264111 LILRB5
-rs532965 32610196 HLA-DRB1 rs2293579 47419207 PSMC3 rs10419669 44794812 CBLC rs10404195 5908958 VMAC
-rs9469112 32447376 HLA-DRA rs1582763 60254475 MS4A4E rs10405693 44823407 BCAM rs4147918 1058177 RDH13
-rs9271162 32609938 HLA-DRB1 rs11824773 60309467 MS4A4A rs73572039 44839373 NECTIN2 19:1051137_CTG_C 1051138 RDH13
-rs9273349 32658092 HLA-DQB1 rs17854357 65834089 SNX32 rs77301115 44893716 TOMM40 rs3752246 1056493 RDH13
-rs9275184 32686937 HLA-DQB1 rs536841 86076782 PICALM rs112849259 44894050 TOMM40 rs35917007 1849148 REXO1
-rs9275614 32716480 HLA-DQA2 rs10792832 86156833 PICALM rs79398853 44895528 TOMM40 Chromosome 20
-rs114812713 41066261 OARD1 rs3844143 86139201 PICALM rs75687619 44896087 TOMM40 rs6064392 56409712 CASS4
-rs7748513 41160234 TREM2 Chromosome 12 rs76366838 44896639 TOMM40 rs6069749 56446519 CASS4
-rs2093395 41187288 TREML2 rs17860282 51346566 CELA1 rs115881343 44899959 TOMM40 Chromosome 21
-rs143332484 41161469 TREM2 Chromosome 13 rs3865427 44877704 NECTIN2 rs62222965 26225443 APP
-rs9394778 41247320 TREML5P
-Figure 4: Comparison of our FVG filter and the group knockoff filter in analyzing the EADB–UKBB dataset: (a) Venn diagram of variant groups identified by different filters; (b) bar plot of size of catching sets identified by different filters (columns); and (c) histogram of purity of catching sets identified by different filters.
-Compared to the group knockoff filter which identifies 152 AD-associated variant groups from 84 loci under the same target FDR level $0.10$, our FVG filter does not suffer power loss.
-As shown in Figure 4 (a), among all variant groups identified by either the group knockoff filter or our FVG filter [^4] Here, we say a variant group is identified by our FVG filter if it contains at least one variant identified by our FVG filter ., 143 groups are identified by both methods while 11 groups and 9 groups are only identified by our filter and the group knockoff filter respectively.
-This suggests the consistency of inference results between our FVG filter and the group knockoff filter.
-However, our filter manages to detect signals at variant rs1532278 close to gene “CLU”, while none of variant groups in the “CLU” locus are identified by the group knockoff filter.
-Without significant power loss, our method is shown to succeed in pinpointing AD-associated variants with higher precision.
-As shown in Figure 4 (b)-(c), all catching sets obtained by our method are of size no larger than $6$ while most of catching sets (129/154) are of size 1 and purity greater than $90\%$.
+In addition, we also manage to identify variants rs6733839, rs744373 and rs2118506 close to gene “BIN1”, variant rs4844610 close to gene “CR1”, variants rs536841, rs10792832 and rs3844143 close to gene “PICALM”, variants rs4277405, rs4309 and rs3730025 close to gene “ACE”, all of which are also reported in He2022 and Bellenguez2022.
+Compared with the model-X knockoff filter, the FVG filter only misses 2 loci identified by the model-X knockoff filter but discover 53 more loci.
+The advantage of our FVG filter over KeLP is more dominating.
+No matter whether the intermediate layer is included in inference of KeLP, all KeLP-identified loci are discovered by our FVG filter while our FVG filter also returns at least another 44 new loci.
+This suggests that our method would not suffer from the same power loss issue of the model-X knockoff filter and KeLP.
+Compared to the group knockoff filter and KnockoffZoom [^4] KnockoffZoom’s result at the group-versus-group level is the same as the group knockoff filter, no matter whether the intermediate layer is included in inference.
+which identify 198 AD-associated variant groups from 91 loci under the same target FDR level $0.10$, our FVG filter does not suffer power loss.
+As shown in Figure 6 (a), among all variant groups identified by either the group knockoff filter (and KnockoffZoom) or our FVG filter [^5] Here, we say a variant group is identified by our FVG filter if it contains at least one variant identified by our FVG filter., 192 groups are identified by both methods while both our FVG filter and the group knockoff filter identify 6 groups not in common.
+This suggests the consistency of inference results between our FVG filter and the group knockoff filter (and KnockoffZoom).
+However, our FVG filter manages to detect signals at variant rs7225151 close to gene “LOC100130950”, while none of variant groups in the “LOC100130950” locus are identified by the group knockoff filter (and KnockoffZoom).
+(a) Venn diagram of variant groups identified by group knockoff filter and our FVG filter.
+(b) Manhattan plot of feature importance statistics of different genetic variant in a high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6, under the FVG filter.
+(c) Comparison of catching set size among different approaches.
+(d) Comparison of catching set purity among different approaches.
+Figure 6: Comparison of our FVG filter and other methods in analyzing the EADB–UKBB dataset.
+Without power loss, our FVG filter returns informative selection sets via simultaneous fine-mapping.
+This can be seen in Figure 6 (c)-(d) where empirical distributions of catching set sizes and purity are displayed.
+Specifically, all but one catching sets of our FVG filter are of size no larger than $8$ while most of catching sets are of size 1 and purity greater than $90\%$.
 In other words, most of catching sets only contain the important variants or strongly correlated proxy variants.
-In contrast, 37 catching sets obtained by the group knockoff filter (Dai and Barber,, 2016) contain more than $8$ variants and are far from being informative with strong impurity.
-Specifically, for the five largest catching sets obtained by the group knockoff filter shown in Table 4, our FVG filter manages to distinguish one or two variants from more than 100 proxy variants in each.
-This suggests the ability of the proposed method in pinpointing AD-associated variants.
-Comprehensive comparisons of our method with all existing methods are presented in Table 5.
+This results in the smallest average size (1.343) and highest average purity (0.951) of catching sets.
+In contrast, 32 catching sets obtained by the group knockoff filter contain more than $8$ variants with strong impurity and are far from being informative.
+Specifically, for the five largest catching sets obtained by the group knockoff filter shown in Table 4, our FVG filter manages to distinguish one or two variants from more than 50 proxy variants in each.
+For comparison, KnockoffZoom (2 layers) only manages to narrow the signal of group “6:31571218:32682663:64” down to the variant rs532965 while the rest four groups are included as a whole in the resolution-adaptive discoveries.
+Adding the intermediate layer only refines the catching set of group “17:43056905:45876021:36” (size: 314) to a catching set of subgroup with size 281 and the catching set of group “6:31571218:32682663:83” (size: 57) to a catching set of subgroup with size 19.
+Thus, the average size and purity of catching sets provided by KnockoffZoom is still suboptimal.
+In addition, catching sets provided by our FVG filter also dominate the KeLP in size.
+Specifically, among all catching sets provided by our FVG filter, more than 95% are if size 1 or 2 (less than 85% for KeLP (2 layers) and KeLP (3 layers)).
+This suggests the advantages of the proposed FVG filter in refining AD-associated genetic discoveries over KnockoffZoom and KeLP.
+Comprehensive comparisons of our method with all existing methods are summarized in Table 5.
 Table 4: List of variants identified by the FVG filter in the five largest catching sets obtained by the group knockoff filter.
-Identified Chromosome Position Catching sets obtained by Closest
+Identified Chromosome Position Catching sets obtained by Closest cS2Gene Annotation
 Variant the group knockoff filter (size) Gene
-rs13408 6 28244970 6:28017819:28917607:9 (121) ZKSCAN4
-rs3130931 6 31167111 6:30798168:31571217:20 (136) POU5F1
-rs3132451 6 31614248 6:31571218:32682663:7 (232) AIF1
+rs3132451 6 31614248 6:31571218:32682663:7 (232) AIF1 AIF1 Promoter,EpiMap
+rs532965 6 32610196 6:31571218:32682663:64 (84) HLA-DRB1 HLA-DRB5 ABC
 rs9469112 6 32447376 6:31571218:32682663:76 (117) HLA-DRA
-rs199439 17 46716137 17:43056905:45876021:36 (314) NSF
-rs199451 17 46724418
-Table 5: Summary of results by applying different methods on the EADB-UKBB dataset.
+rs9271162 6 32609938 6:31571218:32682663:83 (57) HLA-DRB1 HLA-DRB1 EpiMap
+rs9273349 6 32658092 HLA-DQB1
+rs199451 17 46724418 17:43056905:45876021:36 (314) NSF KANSL1-AS1 GTeX
+Table 5: Summary of results by applying our FVG filter to the EADB-UKBB dataset in comparison with existing methods, where bold values indicate the best performance under different evaluation metrics.
 Method Number of Average number of Average Average
@@
 loci identified locus catching sets catching sets
-Marginal association test 51 18.078 18.078 0.477
-Model-X knockoff filter 49 2.816 1.000 1.000
-Group knockoff filter 84 23.714 13.105 0.610
-FVG filter 84 2.440 1.257 0.958
+Marginal association test 54 17.093 17.093 0.489
+Model-X knockoff filter 38 4.263 1.000 1.000
+Group knockoff filter 91 19.923 9.157 0.651
+KnockoffZoom (2 layers) 91 15.152 5.621 0.851
+KnockoffZoom (3 layers) 91 14.681 5.409 0.858
+KeLP (2 layers) 45 10.422 2.535 0.918
+KeLP (3 layers) 38 10.605 2.488 0.932
+FVG filter 89 2.989 1.343 0.951
+: Size and purity of catching sets provided by model-X knockoff filter are trivially 1 and thus not included
+   in comparison.
+In addition to statistical performance, the biological informativeness of discoveries made by our FVG filter is also our concern.
+We apply a SNP-to-gene linking strategy (Gazal2022) to investigate whether the identified genetic variants are functionally enriched or not.
+Here, genetic variants in the same group can functionally annotate different genes or have different annotation types.
+There are in total 7 types of annotations as follows.
+  • Exon: The genetic variant is located less than 20 base pairs from an exon of the annotated gene.
+  • Promoter: The genetic variant lies in promoter regions of the annotated gene.
+  • GTEx: The genetic variant lies in an expression quantitative trait locus of the annotated gene across 54 cell types, according to the GTEx data.
+  • eQTLGen: The genetic variant lies in an expression quantitative trait locus of the annotated gene in blood, according to the eQTLGen data.
+  • Epimap: The genetic variant lies in enhancer regions of the annotated gene across 833 cell types under the EpiMap model.
+  • ABC: The genetic variant lies in Hi-C linked enhancer regions of the annotated gene across 167 cell types under the Activity-by-Contact model.
+  • Cicero: The genetic variant has co-accessibility links to the annotated gene’s promoter across 61,806 blood/basal cells.
+Among 266 genetic variants identified by our FVG filter in Table 3, 180 (67.67%) can be mapped with function evidence.
+This proportion of functional annotation is significantly higher than the average percentage of the background genome (30.52%) and variants in catching sets of the group knockoff filter (56.54%) as shown in Table 6.
+Such an advantage is more significant in annotation types “Exon”, “Promoter” and “ABC”.
+As a result, compared to the group-level inference, discoveries made by the FVG filter is more biological informative for further drug development or therapy designs.
+Table 6: Proportions of different annotations for variants in catching sets of group knockoff filter and the FVG filter.
+Annotation Background Genome Group knockoff filter FVG filter
+Exon 6.74% 11.64% 19.55%
+Promoter 2.68% 7.67% 12.03%
+GTeX 8.33% 22.78% 24.44%
+eQTLGen 1.57% 6.56% 6.02%
+Epimap 10.10% 12.24% 12.41%
+ABC 8.87% 9.76% 14.66%
+Cicero 0.98% 3.09% 4.51%
+Total 30.52% 56.54% 67.67%
 5 Discussions
 As more genetic variants are sequenced using the rapidly developing whole-genome sequencing technology, it is now possible to explain more diseases’ heritability via large-scale genetic studies.
-Noting that existing knockoff methods either lose power or lack informativeness because of strong correlations among variants in Alzheimer’s disease analysis of the EADB-UKBB dataset, we developed a new filter to select important features using group knockoffs.
+Noting that existing knockoff methods either lose power or lack informativeness because of strong correlations among variants in Alzheimer’s disease analysis of the EADB-UKBB dataset, we develop a new filter that uses group knockoffs to simultaneously select important groups and perform fine-mapping within each selected group.
 Specifically, we first define a new family of conditional independence hypotheses that allow inference at feature level under group knockoffs construction.
 Based on theoretical properties of test statistics when group knockoffs are used, we develop the FVG knockoff filter with FDR control.
-Extensive simulated experiments (with both synthetic data and real-world genetic data) empirically validate the FDR control of our proposed filter.
-Compared with the existing knockoff filters, our FVG filter is shown to be powerful in identifying important features with higher precision.
-When applying the proposed method to AD analysis of the real-world EADB–UKBB dataset, we identify as many signals as the group knockoff filter with as high precision as the model-$X$ knockoff filter at the same target FDR level $0.10$.
-Within several variant groups identified by the group knockoff filter, our method precisely distinguishes the variants that contribute more to AD variation.
-Simultaneous inference of conditional independence at different layers of groups is of great necessity as we need to balance the power and the resolution (informativeness) of groups (Katsevich and Sabatti,, 2019; Sesia et al.,, 2020; Gablenz and Sabatti,, 2023).
+Utilizing penalized regressions, our approach can efficiently learn the sparse feature importance and refine the catching set within each group by selecting a subset of most promising features.
+Extensive simulated experiments with real-world genetic data empirically validate the FDR control of our proposed filter.
+Compared with the existing knockoff filters, our FVG filter is shown to return small and pure catching sets without missing many important features.
+When applying the proposed method to AD analysis of the real-world EADB–UKBB dataset, our FVG filter achieves the balance between power (identifies as many signals as the most powerful group knockoff filter) and precision of catching sets (with small average size and high average purity).
+Simultaneous inference of conditional independence at different layers of groups is of great necessity as we need to balance the power and the resolution (informativeness) of groups (Katsevich2019; Sesia2020; Gablenz2024).
 For example, in genetic analysis, variant groups with higher resolution are typically smaller and have higher correlations, making it challenging to identify important signals, but easier to interpret if they are identified.
-Thus, it is interesting to incorporate the proposed filter into the multilayer testing framework (Barber and Ramdas,, 2017; Katsevich and Sabatti,, 2019; Sesia et al.,, 2020; Gablenz and Sabatti,, 2023) to perform simultaneous inference of $H^{(\text{fg})}$’s at multiple layers of group structure.
+Thus, it is interesting to incorporate the proposed filter into the multilayer testing framework (Barber2017; Katsevich2019; Sesia2020; Gablenz2024) to perform simultaneous inference of $H^{(\text{fg})}$’s at multiple layers of group structure.
 In addition, it is also of great interest to adopt the proposed filter in causal inference.
@@
 We leave these tasks as future works.
-Appendix A Proof of Theorem 1
+Appendix A Supplementary Figures for Section 1.2
+(a) Manhattan plot of feature importance statistics of different genetic variants under KnockoffZoom
+(2 layers) at the feature-versus-feature level.
+(b) Manhattan plot of the contribution of different variants to the importance statistics of the group under KnockoffZoom (3 layers) at the intermediate layer.
+(c) Manhattan plot of feature importance statistics of different genetic variants under KnockoffZoom
+(3 layers) at the feature-versus-feature level.
+Figure 7: Manhanttan plot under KnockoffZoom (2 layers) and KnockoffZoom (3 layers) in a high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6.
+Here we exclude plots at the group-versus-group level that is the same as Figure 1 (c).
+Appendix B Proof of Theorem 1
 Without loss of generality, we let $k=1$, $k^{\dagger}=2$, $j\in B_{1}$ and $j^{\dagger}\in B_{2}$.
@@
     $([\mathbb{X}_{j},\mathbb{X}_{B_{1}\setminus\{j\}},\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{1}\setminus\{j\}},\{\mathbb{X}_{-B_{1}},\widetilde{\mathbb{X}}_{-B_{1}}\}],\textbf{y})$ (15)
-[…]
+    ${\mathop{=\joinrel=}^{D}}$ $([\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{1}\setminus\{j\}},{\mathbb{X}}_{j},{\mathbb{X}}_{B_{1}\setminus\{j\}},\{\mathbb{X}_{-B_{1}},\widetilde{\mathbb{X}}_{-B_{1}}\}],\textbf{y}).$
     Note that swapping $\mathbb{X}_{j},\mathbb{X}_{B_{1}\setminus\{j\}}$ and $\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{1}\setminus\{j\}}$ does not change $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s, (6) and (15) implies that conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s,
     $(T_{j},\widetilde{T}_{j}){\displaystyle\mathop{=\joinrel=}^{D}}(\widetilde{T}_{j},T_{j}),$ (16)
-[…]
+    and thus $\text{\rm sign}(W_{j})$ uniformly distributes on $\{+,-\}$.
   • $\star$
@@
     $({\textbf{X}}_{B_{1}},\widetilde{\textbf{X}}_{B_{1}},{\textbf{X}}_{B_{2}},\widetilde{\textbf{X}}_{B_{2}})|(\textbf{X}_{-(B_{1}\cup B_{2})},\widetilde{\textbf{X}}_{-(B_{1}\cup B_{2})})$
-[…]
-[…]
-[…]
+    ${\mathop{=\joinrel=}^{D}}$ $({\textbf{X}}_{B_{1}},\widetilde{\textbf{X}}_{B_{1}},\widetilde{\textbf{X}}_{B_{2}},{\textbf{X}}_{B_{2}})|(\textbf{X}_{-(B_{1}\cup B_{2})},\widetilde{\textbf{X}}_{-(B_{1}\cup B_{2})})$
+    ${\mathop{=\joinrel=}^{D}}$ $(\widetilde{\textbf{X}}_{B_{1}},{\textbf{X}}_{B_{1}},{\textbf{X}}_{B_{2}},\widetilde{\textbf{X}}_{B_{2}})|(\textbf{X}_{-(B_{1}\cup B_{2})},\widetilde{\textbf{X}}_{-(B_{1}\cup B_{2})})$
+    ${\mathop{=\joinrel=}^{D}}$ $(\widetilde{\textbf{X}}_{B_{1}},{\textbf{X}}_{B_{1}},\widetilde{\textbf{X}}_{B_{2}},{\textbf{X}}_{B_{2}})|(\textbf{X}_{-(B_{1}\cup B_{2})},\widetilde{\textbf{X}}_{-(B_{1}\cup B_{2})})$
     If $j,j^{\dagger}\in\mathcal{H}_{0}^{(\text{fg})}$, by (6) and (16), we have conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s,
@@
     and thus $\text{\rm sign}(W_{j})$ and $\text{\rm sign}(W_{j^{\dagger}})$ are independent conditional on $|W_{1}|,\ldots,|W_{p}|$ and $\text{\rm sign}(W_{j})$ for those false $H^{(\text{fg})}_{j}$’s.
-Appendix B Proof of Theorem 2
-As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within each row of Table 2 satisfy the coin-flipping property introduced in Candès et al., (2018).
+Appendix C Proof of Theorem 2
+As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within each row of Table 2 satisfy the coin-flipping property introduced in Candes2018.
 In other words, for each $l$, $\text{sign}(W_{j})$’s for those $j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}$ are i.i.d.
 coin flips conditional on $\{W_{j}|j\in\mathcal{C}_{l}\}$ and $\{\text{\rm sign}(W_{j})|j\in\mathcal{C}_{l}\setminus\mathcal{H}_{0}^{(\text{fg})}\}$.
-Thus, by Lemma 3 of Katsevich and Sabatti, (2019), we have
-$\textbf{E}\left\{\sup_{t>0}\frac{V^{(l)}_{\mathcal{H}_{0},+}(t)}{1+V^{(l)}_{\mathcal{H}_{0},-}(t)}\right\}\leq 1.93,\quad\text{where }\begin{cases}V^{(l)}_{\mathcal{H}_{0},+}(t)=\#\{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}|W_{j}\geq t\},\\ V^{(l)}_{\mathcal{H}_{0},-}(t)=\#\{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}|W_{j}\leq-t\}.\\ \end{cases}$ (20)
+Thus, by Lemma 3 of Katsevich2019, we have
+[…]
 Thus, the FDR of the rejection set $\mathcal{R}^{(\text{fg})}$ obtained by solving (10) satisfies
 ${\text{FDR}}^{(\text{fg})}=$ $\mathbf{E}\Biggl\{\frac{\#(\mathcal{R}^{(\text{fg})}\cap\mathcal{H}_{0}^{(\text{fg})})}{1\vee\#\mathcal{R}^{(\text{fg})}}\Biggr\}$
-$=$ $\mathbf{E}\Biggl{\{}\frac{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\Biggr{\}}$
-$=$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot\frac{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
-$\leq$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
+$=$ $\mathbf{E}\Biggl\{\frac{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\cdot V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\Biggr\}$
+$=$ $\mathbf{E}\Biggl\{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\cdot\frac{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr\}$
+$\leq$ $\mathbf{E}\Biggl\{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\cdot\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr\}$
 $\leq$ $\mathbf{E}\Biggl\{\sum_{l=1}^{\infty}\frac{v_{l}\alpha}{1.93}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr\}$
@@
 by (17).
-Appendix C Extensions
-C.1 Multiple Knockoffs
-Although Algorithm 2 with group knockoffs can obtain the rejection set with FDR control, randomness in group knockoffs generation could produce greatly different $\mathcal{R}^{(\text{fg})}$ in different runs, especially in the case that the number of false $H_{j}^{(\text{fg})}$’s is small and Algorithm 2 could return an empty rejection set (Gimenez and Zou,, 2019).
+Appendix D Comparisons between the Naive Filter and the FVG Filter
+In this section, we empirically compare the naive filter (Algorithm 1), the FVG filter (Algorithm 2) with and without the $1.93$ factor using simulated datasets under the setting of Section 3.1.
+Over 1000 simulated datasets of different sample sizes, the empirical FDR and power under different target FDR levels are reported in Figure 8.
+Although both the naive filter and the FVG filter without the 1.93 factor does not possess theoretically rigorous FDR control, their empirical FDRs remain under control.
+In addition, they both have higher power than the FVG filter (Algorithm 2) with the $1.93$ factor.
+Refer to captionRefer to caption
+(a) $n=500$.
+Refer to captionRefer to caption
+(b) $n=1000$.
+Refer to captionRefer to caption
+(c) $n=2000$.
+Figure 8: Empirical FDR and power of the naive filter (Algorithm 1), the FVG filter (Algorithm 2) with and without the $1.93$ factor with respect to the target FDR level ($\alpha$) over 1000 simulated genetic datasets of different sample sizes.
+Appendix E Extensions
+E.1 Multiple Knockoffs
+Although Algorithm 2 with group knockoffs can obtain the rejection set with FDR control, randomness in group knockoffs generation could produce greatly different $\mathcal{R}^{(\text{fg})}$ in different runs, especially in the case that the number of false $H_{j}^{(\text{fg})}$’s is small and Algorithm 2 could return an empty rejection set (Gimenez2019b).
 Here we extend the proposed Algorithm 2 to multiple group knockoffs $\widetilde{\textbf{X}}^{(1)},\ldots,\widetilde{\textbf{X}}^{(M)}$ where original and knockoff features are simultaneously exchangeable at the level of feature groups.
@@
 $(\widetilde{\textbf{X}}^{(\sigma_{1}(0))}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{K}(0))}_{B_{K}},\widetilde{\textbf{X}}^{(\sigma_{1}(1))}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{K}(1))}_{B_{K}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{1}(M))}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(\sigma_{K}(M))}_{B_{K}})$ (19)
-[…]
+${\mathop{=\joinrel=}^{D}}$ $(\widetilde{\textbf{X}}^{(0)}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(0)}_{B_{K}},\widetilde{\textbf{X}}^{(1)}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(1)}_{B_{K}},\ldots,\widetilde{\textbf{X}}^{(M)}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(M)}_{B_{K}})$
 for any permutations $\sigma_{1},\ldots,\sigma_{K}$ of $\{0,1,\ldots,M\}$.
-With importance scores $\{T_{j}^{(m)}|j=1,\ldots,p;m=0,\ldots,M\}$ obtained in an analogous way of (5), we follow He et al., (2021) to compute feature statistics
+With importance scores $\{T_{j}^{(m)}|j=1,\ldots,p;m=0,\ldots,M\}$ obtained in an analogous way of (6), we follow He2021 to compute feature statistics
 $\kappa_{j}=\arg\max_{m}T_{j}^{(m)},\quad\tau_{j}=\max_{m}T_{j}^{(m)}-\text{\rm median}\{T_{j}^{(m)}|m\neq\kappa_{j}\},\quad j=1,\ldots,p.$ (20)
-Specifically, $\kappa_{j}$ and $\tau_{j}$ are the multiple group knockoffs counterparts of $\text{sign}(W_{j})$ and $|W_{j}|$ respectively (Gimenez and Zou,, 2019) that
+Specifically, $\kappa_{j}$ and $\tau_{j}$ are the multiple group knockoff counterparts of $\text{sign}(W_{j})$ and $|W_{j}|$ respectively (Gimenez2019b) that
   1. $\star$
@@
       (Between-Group Independence) for any $k\neq k^{\dagger}$, $\kappa_{j}$ and $\kappa_{j^{\dagger}}$ are independent for any $j\in B_{k}\cap\mathcal{H}_{0}^{(\text{fg})}$ and $j^{\dagger}\in B_{k^{\dagger}}\cap\mathcal{H}_{0}^{(\text{fg})}$.
-Based on $\kappa_{j}$’s and $\tau_{j}$’s, we provide Algorithm 3 as the multiple knockoffs counterpart of Algorithm 2.
+Based on $\kappa_{j}$’s and $\tau_{j}$’s, we provide Algorithm 3 as the multiple knockoff counterpart of Algorithm 2.
 Algorithm 3 Feature filter with multiple group knockoffs (naive version).
@@
 Here, $(\kappa_{(kl)},\tau_{(kl)})$ corresponds to the feature from the $k$-th group that is aligned in the $l$-th row.
-Row $B_{1}$ $B_{2}$ $B_{3}$cdots
-1 $(\kappa_{(11)},\tau_{(11)})$ $(\kappa_{(21)},\tau_{(21)})$ $(\kappa_{(31)},\tau_{(31)})$cdots
-[…]
-[…]
-[…]
-[…]
-C.2 Filters based on E-values
-Another variation relies on the connection between knockoff filter and $e$-values discussed in Ren and Barber, (2024).
-Following Ren and Barber, (2024)’s way of computing $e$-values of a set of hypotheses whose $W_{j}$’s possess the coin-flipping property, we can analogously compute $e$-values of hypotheses whose $W_{j}$’s are aligned in the same row of Table 2 as shown in steps 4-7 of Algorithm 4.
-By doing so, an e-value $e_{j}$ is calculated for each hypothesis $H_{j}^{(\text{fg})}$ and we obtain the rejection set $\mathcal{R}^{(\text{fg})}$ by applying the e-BH procedure (Wang and Ramdas,, 2022) to $e$-values $e_{1},\ldots,e_{p}$ (as shown in steps 8-10 of Algorithm 4).
+Row $B_{1}$ $B_{2}$ $B_{3}$ $\cdots$
+1 $(\kappa_{(11)},\tau_{(11)})$ $(\kappa_{(21)},\tau_{(21)})$ $(\kappa_{(31)},\tau_{(31)})$ $\cdots$
+2 $(\kappa_{(12)},\tau_{(12)})$ $(\kappa_{(22)},\tau_{(22)})$ $(\kappa_{(32)},\tau_{(32)})$ $\cdots$
+3 $(\kappa_{(13)},\tau_{(13)})$ $(\kappa_{(23)},\tau_{(23)})$ $(\kappa_{(33)},\tau_{(33)})$ $\cdots$
+4 $(\kappa_{(14)},\tau_{(14)})$ $(\kappa_{(24)},\tau_{(24)})$ $(\kappa_{(34)},\tau_{(34)})$ $\cdots$
+$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\ddots$
+E.2 Filter based on E-values
+Another variation relies on the connection between the knockoff filter and $e$-values discussed in Ren2024.
+Following Ren2024’s way of computing $e$-values of a set of hypotheses whose $W_{j}$’s possess the coin-flipping property, we can analogously compute $e$-values of hypotheses whose $W_{j}$’s are aligned in the same row of Table 2 as shown in steps 4-7 of Algorithm 4.
+By doing so, an e-value $e_{j}$ is calculated for each hypothesis $H_{j}^{(\text{fg})}$ and we obtain the rejection set $\mathcal{R}^{(\text{fg})}$ by applying the e-BH procedure (Wang2022) to $e$-values $e_{1},\ldots,e_{p}$ (as shown in steps 8-10 of Algorithm 4).
 Algorithm 4 Rigorous feature filter with group knockoffs using e-values.
@@
 Proof of Theorem 3.
-Because $W_{j}$’s that are in the same row of Table 2 possess the coin-flipping property, by (7) of Ren and Barber, (2024), we have
+Because $W_{j}$’s that are in the same row of Table 2 possess the coin-flipping property, by equation (7) of Ren2024, we have
 $\textbf{E}\left\{\frac{\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\geq t_{l,\alpha})}{1+\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}I(W_{j}\leq-t_{l,\alpha})}\right\}\leq 1.$
@@
 $\sum_{j\in\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}=\sum_{l=1}^{\infty}\sum_{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}\leq\sum_{l=1}^{\infty}p_{l}=p.$
-Finally, by Theorem 2 of of Ren and Barber, (2024), as $\sum_{j\in\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}\leq p$, the e-BH procedure in steps 8-10 of Algorithm 4 returns the rejection set $\mathcal{R}^{(\text{fg})}$ with FDR control.
+Finally, by Theorem 2 of of Ren2024, as $\sum_{j\in\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}\leq p$, the e-BH procedure in steps 8-10 of Algorithm 4 returns the rejection set $\mathcal{R}^{(\text{fg})}$ with FDR control.
 ∎
+Appendix F Detailed Definition of Catching Sets of Resolution-Adaptive Approaches
+Catching sets of resolution-adaptive approaches are formally defined as follows.
+  • $\star$
+    For KnockoffZoom (2 layers), based on selection sets $\mathcal{R}_{\text{KZ}}^{(0)},\mathcal{R}_{\text{KZ}}^{(2)}$ at layers 0 and 2, we derive catching sets as
+    $\mathcal{CC}_{\text{KZ}}=\mathcal{CC}_{\text{KZ}}^{(0)}\cup\mathcal{CC}_{\text{KZ}}^{(2)},$
+    where $\mathcal{CC}_{\text{KZ}}^{(0)}=\{\mathcal{S}_{j,\text{KZ}}^{(0)}=\{j\}|j\in\mathcal{R}_{\text{KZ}}^{(0)}\},$
+    $\mathcal{CC}_{\text{KZ}}^{(2)}=\{\mathcal{S}_{k,\text{KZ}}^{(2)}={B}^{(2)}_{k}|k\in\mathcal{R}_{\text{KZ}}^{(2)}\text{ and }j\notin\mathcal{R}_{\text{KZ}}^{(0)}\text{ for all }j\in{B}^{(2)}_{k}\}$.
+  • $\star$
+    For KnockoffZoom (3 layers), based on selection sets $\mathcal{R}_{\text{KZ}}^{(0)},\mathcal{R}_{\text{KZ}}^{(1)},\mathcal{R}_{\text{KZ}}^{(2)}$ at layers 0, 1 and 2, we derive catching sets as
+    $\mathcal{CC}_{\text{KZ}}=\mathcal{CC}_{\text{KZ}}^{(0)}\cup\mathcal{CC}_{\text{KZ}}^{(1)}\cup\mathcal{CC}_{\text{KZ}}^{(2)},$
+    where $\mathcal{CC}_{\text{KZ}}^{(0)}=\{\mathcal{S}_{j,\text{KZ}}^{(0)}=\{j\}|j\in\mathcal{R}_{\text{KZ}}^{(0)}\},$
+    $\mathcal{CC}_{\text{KZ}}^{(1)}=\{\mathcal{S}_{k,\text{KZ}}^{(1)}={B}^{(1)}_{k}|k\in\mathcal{R}_{\text{KZ}}^{(1)}\text{ and }j\notin\mathcal{R}_{\text{KZ}}^{(0)}\text{ for all }j\in{B}^{(1)}_{k}\}$,
+    $\mathcal{CC}_{\text{KZ}}^{(2)}=\{\mathcal{S}_{k,\text{KZ}}^{(2)}={B}^{(2)}_{k}|k\in\mathcal{R}_{\text{KZ}}^{(2)}\text{ and }k^{\prime}\notin\mathcal{R}_{\text{KZ}}^{(1)}\text{ for all }{B}^{(1)}_{k^{\prime}}\subset{B}^{(1)}_{k}\}$,
+  • $\star$
+    For KeLP (2 layers), based on selection sets $\mathcal{R}_{\text{KeLP}}$, we derive catching sets as
+    $\mathcal{CC}_{\text{KeLP}}=\mathcal{CC}_{\text{KeLP}}^{(0)}\cup\mathcal{CC}_{\text{KeLP}}^{(2)},$
+    where $\mathcal{CC}_{\text{KeLP}}^{(0)}=\{\mathcal{S}_{j,\text{KeLP}}^{(0)}=\{j\}|\{j\}\in\mathcal{R}_{\text{KeLP}}\},$
+    $\mathcal{CC}_{\text{KeLP}}^{(2)}=\{\mathcal{S}_{k,\text{KeLP}}^{(2)}={B}^{(2)}_{k}|{B}^{(2)}_{k}\in\mathcal{R}_{\text{KeLP}}\}.$
+  • $\star$
+    For KeLP (3 layers), based on selection sets $\mathcal{R}_{\text{KeLP}}$, we derive catching sets as
+    $\mathcal{CC}_{\text{KeLP}}=\mathcal{CC}_{\text{KeLP}}^{(0)}\cup\mathcal{CC}_{\text{KeLP}}^{(1)}\cup\mathcal{CC}_{\text{KeLP}}^{(2)},$
+    where $\mathcal{CC}_{\text{KeLP}}^{(0)}=\{\mathcal{S}_{j,\text{KeLP}}^{(0)}=\{j\}|\{j\}\in\mathcal{R}_{\text{KeLP}}\},$
+    $\mathcal{CC}_{\text{KeLP}}^{(1)}=\{\mathcal{S}_{k,\text{KeLP}}^{(1)}={B}^{(1)}_{k}|{B}^{(1)}_{k}\in\mathcal{R}_{\text{KeLP}}\}.$
+    $\mathcal{CC}_{\text{KeLP}}^{(2)}=\{\mathcal{S}_{k,\text{KeLP}}^{(2)}={B}^{(2)}_{k}|{B}^{(2)}_{k}\in\mathcal{R}_{\text{KeLP}}\}.$
+Appendix G Our Implementation of KeLP (Gablenz2024)
+We describe our detailed implementation of KeLP (2 layers) under the simulated experiments in Section 3, while the implementation of KeLP (3 layers) is analogous.
+As a procedure based on $e$-value, KeLP first compute $e$-values for all $H_{j}^{\text{(ff)}}$’s at the feature-versus-feature level and all $H_{k}^{\text{(gg)}}$’s at the group-versus-group level as follows.
+  • For $j=1,\ldots,p$ where $p=1157$, the $e$-value $e_{j}^{\text{(ff)}}$ of $H_{j}^{\text{(ff)}}$ is computed as
+    $e_{j}^{\text{(ff)}}=\frac{c^{(\text{ff})}\times I(W_{j}^{\text{(ff)}}\geq t^{\text{(ff)}}_{\gamma^{\text{(ff)}}})}{1+\#\{j|W_{j}^{\text{(ff)}}\leq-t^{\text{(ff)}}_{\gamma^{\text{(ff)}}}\}},$
+    where $W_{j}^{\text{(ff)}}$ is the feature importance score of $H_{j}^{\text{(ff)}}$ computed using model-X knockoff and
+    $t^{\text{(ff)}}_{\gamma^{\text{(ff)}}}=\min\Biggl\{t>0\Bigg|\frac{1+\#\{j|W_{j}^{\text{(ff)}}\leq-t\}}{1\vee\#\{j|W_{j}^{\text{(ff)}}\geq t\}}\leq\gamma^{\text{(ff)}}\Biggr\}.$
+    In other words, $e_{j}^{\text{(ff)}}$’s are $e$-values computed by model-X knockoff filter under the target FDR level $\gamma^{\text{(ff)}}$.
+  • The $e$-value $e_{k}^{\text{(gg)}}$ of $H_{k}^{\text{(gg)}}$ with respect to ${B}^{(2)}_{k}$ is computed as
+    $e_{k}^{\text{(gg)}}=\frac{c^{(\text{gg})}\times I(W_{k}^{\text{(gg)}}\geq t^{\text{(gg)}}_{\gamma^{\text{(gg)}}})}{1+\#\{j|W_{k}^{\text{(gg)}}\leq-t^{\text{(gg)}}_{\gamma^{\text{(gg)}}}\}},$
+    where $W_{k}^{\text{(gg)}}$ is the feature importance score of $H_{k}^{\text{(gg)}}$ computed using group knockoff with respect to ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ and
+    $t^{\text{(gg)}}_{\gamma^{\text{(gg)}}}=\min\Biggl\{t>0\Bigg|\frac{1+\#\{j|W_{k}^{\text{(gg)}}\leq-t\}}{1\vee\#\{j|W_{k}^{\text{(gg)}}\geq t\}}\leq\gamma^{\text{(gg)}}\Biggr\}.$
+    In other words, $e_{k}^{\text{(gg)}}$’s are $e$-values computed by group knockoff filter with respect to ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ under the target FDR level $\gamma^{\text{(gg)}}$.
+Based on $e_{j}^{\text{(ff)}}$’s and $e_{k}^{\text{(gg)}}$’s, KeLP returns the resolution-adaptive selection set under the FDR level $\alpha$ via solving the constrained optimization problem
+$\max\left(\sum_{j=1}^{1157}x_{j}^{(\text{ff})}+\sum_{k=1}^{345}\frac{x_{k}^{(\text{gg})}}{|{B}^{(2)}_{k}|}\right)$, (22)
+s.t. $x_{1}^{(\text{ff})},\ldots,x_{1157}^{(\text{ff})},x_{1}^{(\text{gg})},\ldots,x_{345}^{(\text{gg})}\in\{0,1\},$
+[…]
+[…]
+[…]
+Here, $D_{\text{total}}=1157+345$ equals the total number of $e$-values computed at both levels.
+Finally, catching sets can be obtained as $\{j|\hat{x}_{j}^{(\text{ff})}=1\}\cup\{{B}^{(2)}_{k}|\hat{x}_{k}^{(\text{gg})}=1\}$ based on the solution $(\hat{x}_{1}^{(\text{ff})},\ldots,\hat{x}_{1157}^{(\text{ff})},\hat{x}_{1}^{(\text{gg})},\ldots,\hat{x}_{345}^{(\text{gg})})$ of (22).
+In Gablenz2024, it is said that choice of $c^{(\text{ff})}$, $c^{(\text{gg})}$ and $\gamma^{(\text{ff})}$, $\gamma^{(\text{gg})}$ are left to users.
+Thus, in our simulation (Section 3), we use $c^{(\text{ff})}=1157$ and $c^{(\text{gg})}=345$ to equal the number of hypotheses in both levels.
+Suppose $\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}$ is the selection set computed by model-X knockoff filter under the target FDR level $\gamma^{\text{(ff)}}$ and $\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}$ is the selection set computed by the group knockoff filter under the target FDR level $\gamma^{\text{(gg)}}$, we have that
+  • $e_{j}^{\text{(ff)}}\geq\frac{1157}{\gamma^{\text{(ff)}}\times|\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}|}$ for all $j\in\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}$;
+  • $e_{k}^{\text{(gg)}}\geq\frac{345}{\gamma^{\text{(gg)}}\times|\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}|}$ for all $k\in\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}$.
+To avoid the case that no catching set is obtained with nonempty $\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}$ and $\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}$, we elaborately tune $\gamma^{(\text{ff})}$, $\gamma^{(\text{gg})}$ as follows.
+Specifically, our target is to include two trivial points in the feasible region of (22), including
+  • the point where $x_{j}^{(\text{ff})}=\text{I}(j\in\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}})$ for $j=1,\ldots,1157$ and $x_{k}^{(\text{gg})}=0$ for $k=1,\ldots,345$;
+  • the point where $x_{j}^{(\text{ff})}=0$ for $j=1,\ldots,1157$ and $x_{k}^{(\text{gg})}=\text{I}(k\in\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}})$ for $k=1,\ldots,345$.
+By doing so, we have nonempty catching sets as long as either $\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}$ or $\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}$ is nonempty.
+To achieve this, we require that
+$\frac{1157}{\gamma^{\text{(ff)}}\times|\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}|}$ $\geq\frac{D_{\text{total}}}{\alpha\times|\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}|},$
+$\frac{345}{\gamma^{\text{(gg)}}\times|\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}|}$ $\geq\frac{D_{\text{total}}}{\alpha\times|\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}|},$
+from the third and forth line of (22).
+This leads to our choice that $\gamma^{\text{(ff)}}=1157/D_{\text{total}}\times\alpha$ and $\gamma^{\text{(gg)}}=345/D_{\text{total}}\times\alpha$.
+When more than $2$ layers are included in the inference, we analogously implement KeLP with $\gamma$ of each layers as $\alpha$ times ratio between the number of hypotheses in this layer and the total number of hypotheses in all layers and the constant $c$ of each layer as the number of hypotheses in this layer.